A Bibliography of Publications by, and about, Max Born

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

18 October 2022 Version 1.126

Title word cross-reference

1/r [DL08]. 137 [Bor35b]. $2.50 [Mor54]. $3.80 [Kle70c]. $4.95 [Kle70b]. _{2} [Bor18g, Pac10]. n log n [ADO11].

1 [Lor10]. 14.99/$25.00 [Ber04]. 1801 [vS21]. 1911 [Meh75]. 1916-1955 [Mar73]. 1926 [Bor54a]. 1928 [HKP^{+}32a, HKP^{+}32b]. 1933 [Krö98]. 1945 [Bor45b]. 1947 [Bor63-28]. 1951 [Bor63b]. 1955 [For70]. 1957 [BBF^{+}57a]. 1970 [KS71, Lin98]. 1995 [Krö98].

20 [Goe88]. 24.80 [Kle70a]. 240pp [Ber04].

3.85 [Bro72]. 37th [Bor53c].

4 [RSS^{+}07]. 480pp [Ber04]. 4th [Bor50b].

50th [Bor59b, Bor63v]. 55 [BB69a].

60th [Ber15, Sho20].

7 [Sho20]. 7th [Sho20].

80th [Ano63, CS79b, Ros79].

978 [Sho20]. 978-1-108-47743-7 [Sho20].

în [Bor69f].

Abbildungsfehler [Bor32c]. aberrations [Bor32c]. Abhandlungen [BB22a, BB63g, Bor63c, Bor63d, Seg64]. Ableitung [Bor10b]. Ablenkung [vS21]. Abraham [BvL23, BvL63]. Absolute [BS35, BL18b, Bor18d, BL63b]. absoluten [Bor18d]. absorbance [BL11]. absorbierenden [BL11]. absorbing [BL11]. absorption [Bor32d]. Absorptionsbanden [Bor32d]. Absorptionsvermögen [BL11]. Achievement [Bor68a]. Achtzehn [BBF^{+}57b, ABH^{+}55, ABH^{+}87]. achtzigsten [Hei62]. Activity [Bor35c, Bor15d, Bor15e, Bor22e, Bor36a, Bor63-51, Bor63-52]. Adiabatenprinzip [Bor26b, Bor63j]. Adiabatensatzes [BF28, BF63b]. adiabatic [Bor26b, BF63b, BF28, Bor63j]. Adsorption [BW63a, BF30, BW31a, BW31b, BF63c]. Adsorptionskatalyse [BF30, BW31a, BW31b, BF63c, BW63a]. Aesthetics [Sur99]. Aether [Bor54a]. Affairs [Bor50d, GR63]. affinité [Bor24c]. affinity [Bor19a, BB20a, BG21a, Bor24c, Bor63l, BB63d]. Affording [Ure67]. after [AS96, Kae48, Krö98]. Again [Ano67, Jon05]. against [ABH^{+}55, ABH^{+}87]. Age [GR63, Bor55e, Bor62a, BB71c, BB84, BB90, Cha05, And81]. ago [Bor01a]. aide [dB53]. air [Hag08]. Aktivität [Bor15d, Bor15e, Bor22e, Bor63-51, Bor63-52]. al [Mor54]. Albert [Bor56a, Bor63u, Bro72, For70, Kle70a, Mar73, Bor55a, Bor56a, Bor63u, BB69a, BB71a, BB71b, BB05, Sch49a, Sch49b, Wei12]. alchemy [Hag08]. Alfons [Krö98]. Alkali [Bor63o, Bor32d, Bor32a]. Alkalihalogenide [Bor32d]. Alkalihalogenidkristalle [Bor32a, Bor63o]. allgemeinen [Bor16a]. Alloys [Bor37a]. Alte [Lor10]. Althoff [Som05, Som05]. Am [Ber06]. Ambition [Jon08]. American [Ber04]. Amplitude [BB21b, BB22a, BB63g, BB63h]. Analysis [KED^{+}60, Man18, Man19]. Analysis/Mehlin [KED^{+}60]. Anatas [BB25a]. angeregten [BH23a, BH63a]. anisotrope [Bor16b, BS16]. anisotropen [Bor17]. anisotroper [Bor18b]. anisotropic [Bor16b, BS16, Bor17, Bor18b]. Anniversaries [Ano52]. anniversary [Bor59b, Bor63v, Sho20]. Ansätze [BR31]. Anthology [BB63a]. Anton [Fis10a, Fis12a, Fis10a, Fis12a]. Antoon [Bor28a, Bor63a]. Antwort [BL19]. Anwendung [Bor19c, Bor63s]. ’Apeiron [BP44c]. aperiodic [BJ25b, BW26b, BW26a, BJ63b]. aperiodischer [BJ25b, BJ63b]. Appleton [Mor54, MPK^{+}54]. Application [Bor38a, Bor19c, Bor31a, Bor63s, Car89]. applied [SG82]. Approaches [BR31]. Approximation [ADO11, Dem03, Hov18, SG52, BS85, GDR04, Hol07, Koz91, PST07, SG82]. April [BBF^{+}57a]. Arbeit [Bor09a, Bor10a, Sch05]. Arbeiten [BL19, Ber15]. Archiv [Lin98]. Arms [vW57]. army [Pal00]. Arnold [Bor42a, Bor42d, Bor63p, Bor51a, Bor51b, Bor52b, Bor63b]. aro [BB90]. article [Bor13d]. Articles [BR95, GR63]. Artikel [Bor13d]. Ascent [Bro73a]. Ash [Krö98]. Ash/Alfons [Krö98]. Aspects [Bor27b, Bor40b, Man18, Bor36c, Bor37b, Meh75]. assistant [Enz02a]. Association [Bor28b, Bor31d]. Assoziation [Bor28b]. astounding [Haw11]. Astrology [Ano01]. astronauts [Bor58c]. Astronomical [Bor53a]. Astronomy [KED^{+}60, FF91]. Astronomy/Born [KED^{+}60]. Äther [Bor19f, Lor10]. Atom [BH27, Bor57b, BL63b, Bor63-27, Bor64a, DM69, Rod19, Bor20a, BB20a, BH23a, BH25, Bor26a, Bor26c, BJ30, Bor60c, BH63a, BB63d, Bor67a, Bor20a, Bor64a, Fow27]. Atombau [Bor40c]. Atomdynamik [Bor26c]. Atome [Bor20e, Bor63r]. Atomen [BH23b, BJN25]. Atomforschern [BBF^{+}57b]. Atomgefahren [BKCK58]. Atomic [And81, BBD^{+}46, Bor35a, Bor46a, Bor47c, Bor51c, BBSR69, Bor70c, Com56, FR13, GR63, Pei97, vW57, BBF^{+}57b, Bor09d, BL18b, Bor18h, BL18a, BL19, Bor20b, Bor24d, Bor26a, Bor26c, BL63a, BB71c, BB84, BB90, Bor23a, Bor27a, Bor65c, Bor71, BRB71, CZPS37]. Atómica [Bor65c, BRB71, Bor71, BB71c, BB84, Alb71]. atomikoan [BB90]. Atomistik [Bor20b]. Atomkonstanten [Bor27a]. Atommechanik [BJ30, BH25, Giu13]. Atommodell [Bor09d, BL18a, BL19, BL63a]. Atommodelle [BL18b, Bor18h, BL63b]. Atoms [Bor26c, Sal19, Bor19a, Bor20e, BH23b, BJN25, Bor63r, Bor63l]. Atomtheorie [Bor23a, Bor24d]. Atomwaffen [ABH^{+}55, ABH^{+}87]. Atomzeitalters [Hei72, Bor55e]. Attempt [Bor65i, BF40e]. attraction [vS21]. Attraktion [vS21]. Aufbau [Bor20b, BB27]. Auffassungen [LBBH69]. Aufsätze [Bor20b]. Ausgewählte [Bor63c, Bor63d, Seg64]. Aussprache [BKCK58]. Authority [Sur99]. Axioms [Man19]. Axis [Bor35c, Bor36a].

B [BDR84]. Bahnbrecher [Hei72]. Baked [BB63a]. Balmer [BFL46]. Band [BJ30, Giu13]. bands [Bor32d]. Based [Khr09]. basic [Bor10b]. Basis [Bor19b, Bor22a, BG48]. Basque [BB90]. Basta [CZPS37]. Baumeister [Gre06, GE08]. Bd [Lin98]. ,,von [Som05]. Be [Ano10, Enz02b, Nik08]. bearbeitet [Bor10b]. Became [Bor65f]. Bedeutung [Bor28e]. before [Nik08]. Began [Kae48]. begriffliche [Bor54b]. Begründer [Bor28c, Bor63-41]. Begründung [Bor14b, BHJ62]. Behandlung [Bor39c]. Beitrag [BF30, BF63c]. Beiträge [KED^{+}60]. belief [Bor57d]. Bellman [KED^{+}60]. Bemerkung [Bor32b, Bor63q]. Bemerkungen [BL19, Bor53g, Bor15a, BB22a, BF25a, Bor61a, Bor63e, BB63g].

Berechnung [BL18b, Bor18d, BL18c, BL63b, BL63c]. Berichtigung [Bor09a, Bor10a, Bor63f]. Berliner [Bor42a, Bor42d, Bor63p]. Bernard [Hei72]. Bertrand [Bro72, Mar73]. Besprechung [Som15, Som34, Som50]. Besprechungen [BvLM23, HKP^{+}32a, HKP^{+}32b]. Best [AG20]. Bestimmung [Bor18e, BB21a, Bor63-50]. Besuch [Bor58c]. Betrachtungen [Bor20g, Bor20h, Bor20i, Bor57a, Bor63z, Bor63g, Hab19]. between [BL11, Bor18e, Bor20c, BK23, Bor40g, Bor43c, Bor63-50, Bor63-46, BB71a, BB71b, BB05, Bro72, CS79b, Ros79]. Beugung [Bor59c]. Beweglichkeit [Bor20j, Bor20k, Bor63-49]. bewegten [Bor10b, Bor12a]. bewegter [Bor10a, Bor10d, Bor14b]. Bewegung [vS21]. Beweis [BF28, BF63b]. Beyond [Bow08, De 16, Khr09, Rod19, Bor66d]. Beziehung [Bor18e, Bor63-50]. Beziehungen [Hei25]. Big [Hei72]. Binary [Bor35c, Bor36a]. binding [Bor31b]. Bindung [Bor24b, Bor31b, Bor30b]. Biographie [GE08, Sto98, SJ98]. biography [Enz02b, GE08, Sto98]. biological [HP02]. Biradicals [BS50]. Birthday [Ano52, Ano63, CS79b, Ros79, Ber15, Hei62, Paw08]. bis [Fis10a, Fis12a, Hei72]. bleibt [Bor63-58]. blessing [Bor66g]. Blessings [Bor66b]. Bodie [Bor10a]. Bodies [Bor10d, Bor10b, Bor10e, BL11, Bor14b, Bor26a, BB63g]. body [Bor10c, vS21]. Bohm [Nik08]. Bohr [BL18b, BL18a, BL19, BH23b, BL63a, CS79b, Nik08, Ros79]. Bohrian [BL63b]. Bohrschen [BH23b]. Bohrscher [BL18b, BL63b]. Bohrsches [BL18a, BL19, BL63a]. bond [Bor24b]. bonding [Bor30b]. Book [A.53, Alb71, And81, Ano46, Ber50, Bey06, Bla58, Bor37a, Bor40c, Bor44a, Bor44b, Bor46b, Bor46c, Bor47d, Bor51d, Bor53b, Bor54a, Bor68a, Bre73, Bro72, Cas05, CZPS37, Cop61, Din50, Din23, Far79, For68, For70, Fow27, Han80, Hec00, Hei72, Hei79a, Hei79b, Hol88, Kle70b, Kle70a, Kle70c, Kör59, Lax55, Lin98, Lin58, Mar50, Mar73, McC49, Mil60, Mil49, Mor54, MPK^{+}54, Nor71, Oes68, Pag63, Pry60, Rom65, Sch79, Seg64, Sho20, Str64, Str60, Süs57, Wol66, da 33, HKP^{+}32a, HKP^{+}32b]. Books [Ano56, BDR84, BB63a, CZPS37, Bor40b]. border [Bor31c, Sch05]. Born [Alb71, Ano46, Ano63, Ano02, Ano10, Anoxx, ABdB^{+}53, Ber06, Bey06, Bla58, BR95, Bre73, Bro72, Cas05, CS83, Cop61, Fis10b, Fis12b, For70, Fow27, Gre06, GE08, Hec00, Hei72, Hei62, K^{+}83, KS71, Kle70b, Kle70a, Kle70c, KED^{+}60, Lax55, Lem82, Lin98, Mar73, Mil60, Sch05, Sch07, Sho20, Som05, Som33, Str64, Str60, Wei82, Wol66, Nik08, A.53, Alb71, Ali05, ADO11, And81, Ano52, Ano56, Ano63, Ano70, Ano82, Anoxx, AG20, Bak83, Bel90, Ber50, Ber05, BB99, BS85, Bla08, Bor10b, Bor20d, BI34c, BBB50, Bor63-57, Bor64c, Bor65f, Bor65g, Bor65h, Bor68b, BB69a, Bor69b, BB71a, BB71b, Bor79, Bor02, BB05, BC80, Bow69, Bow08, Bre70, Bro72, Bro73b, Car89, Cas12, Che70, CS79b, CZPS37, Coo02, De 16, DL08, Dem03]. Born [DM69, Din50, Din23, Dre17, Ein91, Enz02a, Far79, FP09, For68, For70, Fre01, FL10, Fre34, Gal01, GDR04, Giu13, Gow79, Gre05a, Gre05b, Gre06, GE08, GS47, Hai12, Han80, Hei79a, Hei79b, Hei62, Hol07, Hov18, HMLM88, Ily16, Jon05, Khr09, Kle70c, Kon78, Kör59, Koz91, Lan83, Lap69, Lem82, Lem83, Lin58, Man18, Man19, Mar50, Mat78, McC49, Mil49, MS69, Nik08, Nor71, Oes68, Pac10, Pag63, Pai82, PST07, Paw08, Pei71, Pry60, Rom65, Ros79, RS51, SK66, Sch05, Sch07, Sch79, SF05, SG52, Seg64, Sig96, SGM^{+}12, Som05, Som15, Som34, Som50, SG82, Sur99, Süs57, Wei12, Wei82, Wil67, Wol83, Wol95, da 33, vLP52, Kle70a, Mor54, MPK^{+}54]. Borns [Lem83]. Breakdown [BS85]. Breaking [Bor40g]. Brechungsindex [Bor18e, Bor63-50]. Breslau [Anoxx]. bridge [Bor20c]. brief [Enz02b]. Briefe [Som05]. Briefwechsel [BB69a, EB82, Bor69b, For70, Kle70a]. Brillouin [Bor46b]. British [Bor31d, Bon05, Bor31d, Pal00]. Broglie [Dre17]. Brücke [Bor20c]. Bücher [BDR84, KED^{+}60]. Buchwesens [Lin98]. buckling [Bor09g]. builder [Gre06, GE08]. Bulletin [GR63]. Bundesrepublik [Sch05, Sch07]. Busch [Bor62f]. Byers [MPK^{+}54].

C [Ber04, Bor53b, HKP^{+}32a, HKP^{+}32b]. calculate [BF40e]. calculating [BL63c]. Calculation [BL63b, Bor63-47, BS85, Bor13b, BL18b, Bor18d, BL18c, Bor23c, Bor23d, Bor43d, BB43c]. calculus [BC80, GXB21]. Cambridge [Krö98, Sho20]. can [Bor59a]. Cantilever [BP63b, BP42]. Case [Joh13, Bel90]. Cassidy [Ber04]. Catalog [Lem83]. Catalysis [BW63a, BF30, BW31a, BW31b, BF63c]. Cause [Ber50, Bor39b, Bor39a, Bor49f, Din50, Mar50, McC49, Mil49, Som50]. Caused [BP63b, BP42]. Centenary [CS83, Ano82, Bor31d, Lan83]. Centre [BF40a, BF40b]. Century [Bro72, Bor50e, Bra09, Ber04]. Certain [Bey06, Cas05, BB43c, Gre05a]. Chair [ABdB^{+}53]. champ [Bor38e]. Chance [Bor49f, Som50, Ber50, Din50, Mar50, McC49, Mil49]. Change [Krö98, AS96]. character [Bor21b, Bor63-54]. charge [Bor13d, Bor18e, Bor63-50]. Charged [Bor36d, Bor36e]. Charles [Bor15a]. Chemical [Bor30a, Bro68, Bro06, FW08, Hab22, Tho22, Bor24b, Bor24e, BH24a, Bor24a, Bor24c, Bor29c, Bor29a, Bor30b, Bor31f, BH63b, Bor63-60, Cha05, Hab86, HP02, Hol88, Pal00, Bor31b]. Chemie [Ber15, Bor20c, KED^{+}60, Bor31c]. Chemiker [Sto98]. chemische [Bor24b, BH24a, Bor24a, BH63b, Bor31b]. chemischen [Bor24e, Bor29c, Bor30b, Bor31f, Bor40c, Bor63-60]. Chemist [Sto98]. Chemistry [Joh13, Ber15, Bor20c, Bor31c]. Chemists [Joh13]. Chimie [Bor40c]. chimique [Bor24c]. Churchman [KED^{+}60]. Ciencia [BB71c, BB84, Alb71]. Circuits [MPK^{+}54]. Classical [Khr09, RSS^{+}07, Bor55f, Bor58h, Bor59a, Bor59e, Lan83, Bor36d, Bor36e]. classique [Bor59a]. clock [Bor58c, BB58, BB63b]. closely [vS21]. cloud [Hab86, Hol88]. cohesive [Bor19d, Bor19e]. Coined [Gal01]. Cold [Sal19]. Collected [BDR84]. Collision [Bor77, BJN25, Bor26e, Bor26d, Bor26f, Bor28e, Bor63-59, Bor63-36, Bor63-64]. Colloquium [Bor45b]. color [Bor63g]. Comment [Bor58f]. commentaries [BB05, Bro72]. Commentary [Got83, Kle70a]. Comments [Ano67, Bla67, Bor68b, Ure67, Wil67, BB22a, Bor63e, BB63g]. Commission [BBD^{+}46]. communication [Bor15a]. Comparison [BR95]. Composite [Bla08]. compressibility [BL18c, BL63c]. Concept [Bor58a, Bor58b]. conceptia [Bor69f]. Conception [CZPS37]. Conceptions [LBBH69]. Concepts [Man19, Bel90]. conceptual [Bor53c, Bor54b]. conciencia [Alb71, BB71c, BB84]. condensée [Bor52d]. Condensing [BF38, Bor63-44, Bor37c]. conditions [Bor06]. Conference [CS83, BB55, BV09]. conferences [Meh75]. Conflict [CS79b, Ros79]. Congress [Bor28b]. Conscience [BB69b, Lem82, BB71c, BB84, BB90]. Considerations [Bor63z, Bor20g, Bor20h, Bor20i, BF40e, Bor63g]. Constant [BS35, Cas12, Bor18e, Bor63-50]. Constants [Bor46d, Bor46e, Bor63t, BH24a, Bor24a, Bor27a, BH63b]. constitution [Bor23b]. constrained [BP22, BP63a]. contingency [Sig96]. Continuity [Bor55b, Bor63h]. Continuum [SK66]. contribution [Bon05, BF30, BF63c]. Control [Sur99]. Controversy [Sur99]. Cooperation [Sal19]. Correction [Bor09a, Bor10a, Bor63f]. Correlation [MS69]. Correspondance [Mar73]. Correspondence [BB69a, Bro72, BB71a, BB71b, BB05, Pei71, EB82]. Corridors [Bor52a]. Corrigenda [Bor44c, Bor63i]. Cosmic [Bor34a]. Cosmological [Cas12]. Cottrell [MPK^{+}54]. Coulomb [Hol07]. Coupled [Bor35c, Bor36a]. couples [Bor37d]. Coupling [Bor51g, Bor63y]. Courant [Som05, Som05]. Covariant [Bor42e]. creators [Str11]. Criterion [Hov18]. Critical [Bor20g, Bor20h, Bor20i, Bor63z]. cross [Koz91]. crossers [Sch05]. crossroads [BV09]. Crystal [BH54, BL63b, Bor63i, Bor63-31, Bor63-32, Bor63-47, Bor63-46, BB63j, Lax55, Bor15b, BL18b, Bor18d, BL19, BS19, Bor21d, BB22b, Bor23a, BGM33b, BT34, Bor40e, BF40e, BM40, Bor42e, Bor43d, BB43c, Bor43c, Bor44c, BS63b, Bor63-63, Bor64d, BH88, Mis40, Som15, BL18a, BL63a]. Crystals [BB43a, BB42a, Bor39e, Bor63-48, BB63f, Bor63o, GS47, BB47a, Bor15c, Bor15e, BL18c, Bor18a, Bor18e, Bor18f, BS19, Bor22e, BM32, Bor32a, BGM33a, BB42b, BL42, BB43b, Bor46f, BB47b, Bor49h, Bor51e, Bor51g, Bor59c, BL63c, BM63, Bor63y, Bor63-50, Bor63-52, BS63b, Bor63-53, HKP^{+}32a, HKP^{+}32b, MPK^{+}54]. cubic [Bor18h, BF40e]. Curies [Kae48]. Current [Bor68b]. curse [Bor66g]. Curved [Cas12]. Czukor [Bor15a, Bor15a].

D [Fow27, Lem82]. Dachboden [Som05]. Damped [BP63b, BP42]. dangers [BKCK58]. Dans [Bor59a, CZPS37]. dargestellt [Bor20d, Bor22a]. Darstellung [Bor20g, Bor20h, Bor20i, Bor63z]. David [Ber04, Som05, Som05]. day [HHW99]. death [Bor59b, Bor63v]. debate [Car89, BKCK58]. Debye [Som05, Som05]. Declaration [BBF^{+}57a, BBF^{+}57b]. Dedicated [CS79b, Ros79]. Déduction [BI34d]. déduit [Bor49e]. Dee [Ber04]. Definition [Bor10c]. Definitions [KED^{+}60]. deformability [BH24a, Bor24a, BH63b]. deformations [Bor42e]. Deformierbarkeit [BH24a, BH63b]. Delta [Lap69]. Delta-Function [Lap69]. Demonstrability [BS50]. Density [BL43, BY50, Bor36b]. dependence [Bor22f]. dependent [PST07]. Derivation [Dem03, SF05, Ali05, Bor10b, BI34d]. derived [Bor49e]. Deriving [AG20]. Detection [Khr09]. determination [Bor18e, BB21a]. determining [Bor63-50]. Determinism [Bor63h, Bor55b, Bor55h]. Determinismus [Bor55h]. deterministic [Bor55f]. deterministisch [Bor55f]. Deutscher [Sto98]. Deutschland [Sch07, Som05]. Deutung [Bor24e, Bor32d, Bor55d, Bor61a, Bor63e, Bor63m]. development [Bor53c, Lor10, Meh75, Bor55e]. deviation [vS21]. Diagram [Bow69, Che70]. Dialectical [Fre01, FL10]. Dialog [LBB68]. dialogue [Bel99]. Diamanten [Bor14d, Bor63-61]. Diamond [Bor46d, Bor63t, Bor14d, Bor49h, Bor63-61]. diatomic [Bor18f, BF33, BF63a, Bor63-53]. Dicke [BB21a]. Did [Bor65g]. dielectric [Bor18e, Bor63-50]. dielectrics [Bor15a]. Dielektrika [Bor15a]. Dielektrizitätskonstante [Bor18e, Bor63-50]. different [Bor06]. differential [Koz91]. Diffraction [BW59, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, Hec00, SG52, Wol66, Bor59c, Str64, Str60]. Diffuse [Bor41a, Bor42f, BLS42, Bor42g, BL42]. Dimensional [Lap69]. dimensions [Bor18d]. Dipolcharakters [Bor21b, Bor63-54]. dipole [Bor21b, BK23, Bor63-54]. Dipolmoments [BK23]. Dirac [BI34d, Bor52c]. direct [Bor20e, Bor21b, Bor63r, Bor63-54]. direkte [Bor20e, Bor63r]. direkten [Bor21b, Bor63-54]. discoveries [Bra09]. discovery [Bor59c, BC80, Hag08, Kae48]. Discussion [Bor43d, Bor63-47]. Discussione [HSBA60]. Discussions [HSBA60]. disintegration [Bor29d, Bor29e, Bor63-62]. Dislocations [MPK^{+}54]. Dispersion [Bor15c]. Dissipation [BF25a]. Distribution [BG63, Lin98, BvK13a, BG46, BvK63a]. Disturbance [BP63b, BP42]. DM [Kle70a]. d’ondes [BI34d, dB53]. doomed [Hag08]. Dr [Bor42a]. Dr. [Bor42d, Bor63p]. dreams [Haw11]. Drehung [Bor15c]. Drehungsvermögens [Bor18b]. Drei [Bor20b]. Dreier [Sch05]. Dualism [BB68, LBB68]. Dublin [Bor45b]. Duckworth [Ber04]. Duckworth/Ivan [Ber04]. due [vS21]. dünner [BB21a]. Duration [BFL46]. durch [Bor59c, vS21]. dynamic [Bor24b]. Dynamical [BH54, BH88, BB47a, BGM33b, BGM33a, BG47a, Lax55]. Dynamics [BL43, Bor63-46, Bor70c, Bow69, Sur99, Bor09e, Bor09f, Bor23a, Bor26a, Bor26c, Bor42c, Bor43c, BH55, BH56, Bor64d, Bor15b, Som15]. Dynamik [Bor09e, Bor09f, Bor23a, BH55, Bor15b, Som15]. Dynamische [BGM33b, BGM33a]. dynamisches [Bor24b].

Early [Jam74a, Sch05, Sch07, Sal19]. Ebene [Bor06, Bor63-43]. Economy [Bor39b, Bor39a]. ECOSA [CS83]. Edinburgh [Bor36c, CS83, ABdB^{+}53]. edited [Bor10b, BB63a]. edition [HKP^{+}32a, HKP^{+}32b, Kle70c, Sho20, Bor50b]. Editor [Bor47a, Ano67, BR57, Hab22, Tho22]. Edmund [Bor54a]. Edward [Mor54, MPK^{+}54]. Effect [Bor42b, Bor46i, BB63f, Bor28f, BS41, BB45, BB47b, Bor49h]. Egypt [BB46]. Ehrenfest [Som05, Som05]. Eigenfrequenzen [Bor32a, Bor63o].

Eigenschwingungen [BvK13a, Bor14c, Bor18f, Bor63-53, BvK63a]. Eight [Bor53b]. eighteen [BBF^{+}57b, ABH^{+}55, ABH^{+}87, Got83]. eightieth [Hei62]. einem [Som05]. Einfluß [BS19, BH24a, Bor24a, BH63b, BS63b]. Einstein [Ano06, Bor55a, BB69a, Bor69b, Bre73, Bro72, For70, Kle70a, Mar73, Som05, Bor16a, Bor19b, Bor20d, BHvL21, Bor22a, Bor22c, BB24, Bor49b, Bor50a, Bor50b, Bor50d, Bor55a, Bor56a, Bor62b, Bor63u, Bor64b, Bor65b, Bor65a, BI67, BB69a, Bor69b, BI69, Bor69c, BB71a, BB71b, Bor01b, BEP03, BB05, Car89, CS79b, Pei71, Ros79, Sch49a, Sch49b, Som05, Wei12]. Einstein” [Som05]. Einsteins [Bor19b, Bor20d, Bor22a, Bor64b, Bor69c, Bor01b, BEP03, Din23, Rom65]. Einwande’ [Bor13d]. Einzeldarstellungen [BF31]. Elastic [Bor46d, Bor46e, Bor63t, Bor06, Bor63-43]. Elasticity [Bor11]. elastischen [Bor06, Bor63-43]. Elastizitätstheorie [Bor11]. Electomagnetic [Wol66]. Electric [BFP45, BFP63, BS85, Bor36e]. electrical [Bor19d, Bor19e, Bor19f, Bor24e]. Electricity [Bor54a, MPK^{+}54]. Electro [Str64]. Electro-optical [Str64]. Electrodynamics [BB99, Bor10a, Bor10d, BCG49, Bor63k, Bor07, Bor14b, Bor14a, BR31, BI34d, BI34e, Bor34c, Bor35e]. électrodynamique [BI34d, BI34e]. electrolytic [Bor20j, Bor20k, Bor63-49]. Electromagnetic [BI33a, Bor34b, BW59, Bor63-30, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, Hec00, Str64, Str60, Bor10b, Bor18a, Bor33c, Bor36b, Bor38e, BF39a, BF39b, Bor43b, Bor53e, Bor65d, Bor72, BP46]. électromagnetique [Bor38e]. Electron [Bor52c, CZPS37, Fre34, SG52, Bor09a, Bor09e, Bor09f, Bor09b, Bor19a, Bor20b, BB20a, BH23a, Bor32b, Bor33b, BI34c, BP46, Bor51g, Bor63q, Bor63f, Bor63l, BB63d, Bor63y, Bor63n, Bor18b, BG21a, BH63a]. electrons [Bor10b]. electrostatic [Bor21c, Bor63-55]. elektrische [Bor19d, Bor19e, Bor24e]. elektrischen [Bor19f]. Elektrodynamik [Bor10a, Bor10d, Bor14b, Bor14a, Bor63k]. elektrolytischen [Bor20j, Bor20k, Bor63-49]. elektromagnetische [Bor18a]. elektromagnetischen [Bor10b, Bor33c, Bor43b, Bor53e, Bor65d, Bor72]. Elektronen [Bor51g, Bor63y]. Elektronen- [Bor51g, Bor63y]. Elektronenaffinität [BB20a, Bor63l, BB63d, Bor19a, BG21a]. Elektronenbahnen [BH23a, BH63a]. Elektronenradius [Bor32b, Bor63q]. elektronentheoretische [Bor14b]. Elektronentheorie [Bor10b, Bor18b, Bor20b]. Elektrons [Bor09a, Bor09e, Bor09f, Bor09b, Bor63f, Bor63n]. elektrostatische [Bor21c, Bor63-55]. elemental [Bor22c]. Elementar [Bor22a]. Elementare [BJ30]. Elementary [Bor22a, Bor46j, Bor49c, Bor49g, CZPS37, Bor22c, Bor65e, BJ30, Bor46c]. Elements [Hei72]. elle [Bor59a]. Elsasser [And81]. Elster [Ber15, Ber15]. Embattled [Sal19]. Embryologist [BBB50]. Emigration [Krö98]. Emigré [Krö98, AS96]. Emil [Cop61, Hec00, Mil60, Str64, Str60, Wol66]. emission [BL11]. Emissions [BL11]. Emissions- [BL11]. Encyclopedia [Bor20f].

End [Bey06, Cas05, Gre05a]. endlicher [BB21b, BB22a, BB63g, BB63h]. enduring [Bon05]. Energie [Bor14a, Bor63k]. Energy [BBD^{+}46, Bor63k, Bor07, Bor14a, BS19, Bor36b, Bor46f, Bor47c, BS63b]. enigma [Ber04, Ber04]. Entanglement [Bla08]. Entdeckung [Bor59c]. Entgegnung [Bor13d]. entitled [Fre01]. entscheidende [Bor55c]. Entwicklung [Lor10, Bor55e]. envariance [Bla08]. Enzyklopädie [Bor20f]. Eotvös [BC13b, BC13a]. Eotvösschen [BC13b, BC13a]. episodes [Bra09]. Epistemological [CS79b, Ros79]. equation [BI34d, BI34d]. Equations [Bor61c, BI63b, Ily16, Bor10b, Bor33b, BI34b]. era [Alb71, BB71c, BB84]. Erforschung [Fis10a, Fis12a]. erhobenen [Bor13d]. Erinnerungen [Bor69a, Bor75, Bor56a, Bor59b, Bor59c, Bor63u, Bor63v, Bor65b, BI67, Bor68c, BI69]. Erklärung [BBF^{+}57b]. Ernst [Anoxx, Bor48a, Bor63-28, Anoxx]. Errata [Ano02]. Erratum [Ber06, BW31a, Bor36a, BW31a]. Erster [Giu13]. Erwiderung [Bor55c]. Erwin [BG61, Bor63w]. escaped [Cou13]. Essay [Gow79]. essays [Bor20b]. Estate [Lem83, Bor10b]. eta [BB90]. ether [Bor19f, Lor10]. Euler [Bor09g]. Eulerschen [Bor09g]. Europaische [Bor57a]. Europe [Bor58d, Rod19]. European [Bor57a]. Events [SGM^{+}12, Bla08]. Everettian [Man18, Man19]. Evils [Bor66b]. Ewald [HKP^{+}32a, HKP^{+}32b]. excitation [BJN25]. excited [BH23a, BH63a]. Exclusive [SGM^{+}12]. exile [Sig96]. Expelled [MP01]. Experiment [Bor43a, Bor56b, Bor69e, Ano56, Kör59]. Experimental [BBB50]. Explanation [AG20, Car89]. Exposición [Bor22c]. extent [Bor59a]. Extraordinary [Bor68a].

F [Bor37a, Bor53b, CZPS37, Hol88, MPK^{+}54, Som34]. F.R.S [Bor40d]. F.R.S. [BR95, Bor58g, Bor63-39, Bor63-40]. F.R.S.E. [BS63a]. Fajans [MS69]. fall [Cha05]. Fame [BHvL21]. Family [Ano02, Bor02]. Faraday [Bor31d]. Faraday-Jahrhundertfeier [Bor31d]. Farbenlehre [Bor63g]. fed [Hag08]. Federal [Sch05, Sch07]. Feinstruktur [BB33, BB63c]. Ferdinand [Lin98]. festen [Bor23a, Bor26c, BB27]. fester [Bor19d, Bor19e, BB21c, BB22a, BB63g, BB63i]. Festschrift [Ber15, Ber15]. Field [BI33b, Bor34a, Bor34b, BS35, Bor35d, BI63a, BI63b, Bor63-30, KED^{+}60, Ali05, Bor33b, BI34a, BI34b, BI35, Bor36b, Bor36d, Bor36e, Bor38e, BP46]. Fields [BP44c, Bor50c, Khr09, Ber15, BP44a, BP44b, BP46]. fifties [Fre01]. Fifty [Bor51f, Bor51h, Kae48]. Film [Bor68a]. fine [BB33, BB63c]. Finite [BB63g, BB21b, BB22a, Bor33b, BB63h]. First [Hol88, Hab86]. Fisher [Fow27]. Fisica [Bor71, Bor61b, HSBA60, Bor65c, BRB71]. fisicos [Bor22c]. five [Bor40b]. Fizica [Bor69f]. Flight [Bor58f]. Flow [MPK^{+}54, BBK13]. fluctuations [BF39a, BF39b]. Fluid [Bow69, Bor21b]. fluids [Bor16b, BS16, BG47b]. Flüssigkeiten [Bor15d, BS16, Bor63-51, Bor16b, Bor18b]. Flüssigkeitsmolekeln [Bor21b, Bor63-54]. Flux [Bor57e, Bor58e, Bor59d, Bor66e, Bor69f, Bor83, Bel90]. For.Mem.R.S.

[Bor47b, Bor63-35]. Force [CZPS37]. Forced [AS96, Krö98]. Forces [Bor47e, Bor19d, Bor19e, Bor24e, Bor29c, Bor31f, Bor63-60, Bor47d]. Foreword [Bro72]. form [HP82, HP02]. Formal [Man18]. Formalism [Wei12]. Formation [BS50, BF25b, BF63d]. Formeln [Bor53g]. former [Bor57d]. Formula [Dem03, Bor09g, Bor53g, Bor54c]. Formulation [BW63b, BW26b, BW26a, GXB21]. Formulierung [BW26a, BW63b]. Forum [Dys63]. found [Som05]. foundation [Bor14b]. Foundations [BI33b, BI34a, BHJ62, BI63a, BB47a, Bor20d, BB27, Bor49d]. founder [Bor28c, Bor63-41]. Fourier [BFP45, BFP63]. Fourier-Transformer [BFP63]. Frage [Bor55h]. Fragen [Lor10]. fragment [BGĬ90]. Francis [Bor58g, Bor63-39]. Franck [Ano52, K^{+}83, Lem82, vLP52, BW64a, HMLM88, Lem82, Wei82]. Frank [Lin98]. free [Bor20e, Bor30a, Bor46f, Bor63r]. freien [Bor20e, Bor63r]. French [Bor24c, BH28, Bor31e, BI34d, BI34e, Bor38e, BC48a, Bor49e, BB55, Bor59a, BH84, dB53]. Frenkel [BI34c, Bor62c]. Frequencies [BL43, Bor63o, Bor18f, Bor32a, BT34, Bor63-53]. Freund [Anoxx]. Freundlich [Bor53g, Bor54c]. Freundlichs [Bor53g]. Friedrich [Hei72]. friend [Anoxx]. friendship [BB05]. Frisch [KED^{+}60]. Fritz [Cha05, HH70, MVM10, Sto98, SJ98]. front [Pal00]. frontier [Sch07]. frühen [Sch05, Sch07]. Fuchs [Bon05]. fueled [Hag08]. Function [Lap69]. Functions [BG63, BF40d, BF40c, BG46]. fundamental [Bor22c]. fundamentos [Bor22c]. Furth [Ano46]. future [Bor53c, Bor69d]. futuro [Bor69d].

G [CZPS37, HKP^{+}32a, HKP^{+}32b, Krö98]. Gallery [Pai00]. Gänge [Bor52a]. gas [HP82]. Gasen [Bor15d, BG21b, Bor63-51]. Gases [RS51, Bor15d, BG21b, Bor63-51]. Gebieten [Ber15]. Geburtstag [Ano63, Ber15, Hei62]. Gedanken [Bor68c]. Gedanken” [Bor69a]. gefunden [Som05]. gegen [Bor13d]. Gehrcke [Bor13d, Bor13d]. Geitel [Ber15, Ber15]. gemeinsam [Anoxx]. gemeinverständlich [Bor20d]. General [Bor36a, BG49, BGK50, BG63, Che70, BG46, BG47a, BG47b, Bor16a, Bor35c, RSS^{+}07]. generalization [Bor09g]. Generalized [ADO11]. generatiei [Bor69f]. Generation [Bor56c, Bor69g, Bor70a, Bor70b, Hei72, Kle70c, Nor71, Pry60, Süs57]. Genesis [RSS^{+}07]. Genius [Pai00, Hag08]. Geometrical [Bor32c]. Geometrische [Bor32c]. geradlinigen [vS21]. Gerald [Ber04]. germ [HP82]. German [Krö98, ABH^{+}55, ABH^{+}87, Ano63, Anoxx, AS96, Ber15, BBF^{+}57a, BBF^{+}57b, BO07, Bor09a, Bor09e, Bor09f, Bor09d, Bor09c, Bor09b, Bor09g, Bor10b, Bor10a, Bor10c, Bor10d, Bor10e, Bor11, BL11, Bor12a, Bor12b, BK12, Bor13b, Bor13e, BBK13, Bor13d, Bor13c, BC13b, BC13a, BvK13b, Bor13a, BvK13a, Bor14b, Bor14a, Bor14c, Bor14d, Bor15a, Bor15c, Bor15b, Bor15d, Bor15e, Bor16b, BS16, Bor16a, Bor17, BL18b, Bor18d, BL18c, Bor18a, Bor18b, Bor18h, BL18a, Bor18e, Bor18f, Bor18g, BL19, Bor19a, Bor19d, Bor19e, BB19, Bor19f, BS19, Bor19b, Bor19c, Bor20a, Bor20b, Bor20j, Bor20k, Bor20c, Bor20e, Bor20f, BB20a, BB20b, BB20c, Bor20g, Bor20h, Bor20i, Bor20d, Bor20l, Bor21b, Bor21c, BG21a, BB21a, Bor21a, BHvL21, BB21b, BB21c, Bor21d, BG21b, BB22a, Bor22b, Bor22d, Bor22e, BP22, Bor22a, BB22b, Bor22f, Bor23a, BvLM23, BH23a, BH23b, BH23c, Bor23c, Bor23d, BK23]. German [Bor24d, Bor24b, Bor24e, BH24a, Bor24a, Bor24f, BH24b, BF25a, BB25a, BB25b, BJ25b, BJ25a, BF25b, BF31, BJN25, BH25, Bor26b, BW26a, Bor26c, BHJ26, Bor26e, Bor26d, Bor26f, BB27, Bor27a, BO27, Bor27c, BF28, Bor28e, Bor28b, BEvL^{+}28, Bor28c, Bor28d, Bor28f, Bor29c, Bor29b, Bor29d, Bor29e, BF30, BJ30, Bor30b, BR31, Bor31b, BW31a, Bor31d, Bor31c, BW31b, Bor31f, Bor32b, Bor32d, BM32, Bor32c, Bor32e, Bor32a, BGM33b, BGM33a, BB33, Bor33a, Bor33c, BF33, Bor39c, Bor43b, Bor48b, BS48, Bor51e, Bor51g, Bor52a, Bor53e, Bor53g, Bor54b, Bor55a, Bor55c, Bor55e, Bor55h, Bor55f, BH55, Bor55d, Bor56a, Bor56d, Bor57a, Bor57e, BKCK58, Bor58c, Bor58e, BL58, Bor58b, BB58, Bor58h, Bor59b, Bor59c, Bor59d, Bor59e, Bor60a, Bor60e, Bor61a, BHJ62, Bor62a, Bor63c, Bor63d, BL63b, Bor63j, BF63b, Bor63q, Bor63-49, Bor63g, BL63c, BF63c, Bor63e, Bor63f, BB63g, Bor63r]. German [Bor63-54, Bor63u, BH63b, Bor63-55, Bor63l, Bor63v, BH63a, BB63d, BB63c, BM63, BB63e, Bor63x, Bor63k, BL63a, Bor63z, Bor63y, Bor63-50, BW63b, Bor63-51, Bor63-52, BS63b, BW63a, BJ63b, Bor63-60, BP63a, Bor63-56, BJ63a, BHJ63, BL63d, BF63d, BO63, Bor63-37, Bor63-59, Bor63-36, BF63a, Bor63-61, Bor63-41, Bor63m, Bor63-43, BB63h, BvK63b, BB63i, Bor63s, Bor63-63, Bor63-62, BB63j, Bor63n, BvK63c, Bor63o, Bor63-53, BB63b, BvK63a, Bor63-57, Bor63-58, Bor63-64, Bor64a, Bor64b, Bor64e, Bor64f, Bor65b, Bor65d, Bor65j, Bor65i, Bor65k, Bor66c, Bor66f, Bor66e, Bor66g, BI67, Bor68c, BB69a, Bor69b, BI69, Bor69a, Bor69e, BB69b, Bor69c, Bor72, Bor75, Bor83, Bor01b, BEP03, Bus65, EB82, Fis10a, Fis12a, Gre06, GE08, Hab19, HH70, Hei25, Hei62, HKP^{+}32a, HKP^{+}32b, Inf55, Joh13, LBBH69, Lem82, Lem83, Lor10, Sch05, Sch07, vS21, Som05, Som15, Som33]. German [Som34, Sto98, Wei82]. German-Speaking [Krö98, AS96]. Germany [Joh13, Sch07, Som05, vW57]. Gesammelte [BDR84]. Geschichte [Inf55, Lin98]. Gesetzes [BC13b, BC13a]. gestörter [BP22, BP63a]. Gewissens [BB69b, Lem82]. Gift [MP01]. Gitterpotentiale [Bor21c, Bor63-55]. Gittertheoretische [Bor39c]. Gittertheorie [BB19, BB20b, BB20c, BB25a, BB25b, Bor26c, BM32, BGM33b, BGM33a, BM63, BB63e, BL18c, Bor19c, BG21a, BL63c, Bor63s]. Gleason [De 16]. Gleason-Type [De 16]. Goeppert [Goe88]. Goettingen [Som05]. Goldschmidt [Bor47b, Bor47f, Bor63-35]. Good [BB63a]. Göttingen [Bor09d, HMLM88, Enz02a, Got83, K^{+}83, Wei82, Bor09d]. Göttinger [Som05, Wei82]. Gravitation [Bor16a, Bor16a, RSS^{+}07]. Gravity [Cas12]. great [Kae48]. Green [RS51]. Greenspan [Bey06, Cas05]. Grenzbedingungen [Bor06]. Grenzgänger [Sch05, Sch07]. Grenzgebiet [Bor31c]. grids [BvK13a, BvK63a]. Grosse [Hei72]. Ground [BP63b, BP42].

Grundgleichungen [Bor10b]. Grundlagen [KED^{+}60, Bor20d, Bor22a, Bor19b, BB27]. Guide [Rod19]. Gültigkeitsgrenze [Bor51e]. Gustav [Bor14a, Bor63k, Bor14a, BBB50, Bor63k, Bor07]. Guthrie [Bor53c].

H [Bor18g, Bor37a, MPK^{+}54, Bor18g, Pac10]. H. [Bor18c]. Haber [Sto98, SJ98, Cha05, HH70, Hol88, MVM10, MS69, Sto98, Tho22]. Habilitation [Bor09d]. Habilitationsvortrag [Bor09d]. Hafner [Mor54]. half [Bor50e]. half-century [Bor50e]. Halide [Bor63o, Bor32a]. halides [BK23, Bor32d]. Hall [Ber04]. halogen [Bor19a, Bor63l]. Halogenatome [Bor19a, Bor63l]. Halogenwasserstoffe [BK23]. Hand [Bor40c]. Hand- [Bor40c]. Hans [Ber15, Ber15]. Happy [Paw08]. hardback [Sho20]. hardcover [Ber04]. harmonic [BP42]. Harmonie [BP63b]. Hartree [Fow27]. harvest [Bra09]. Hauptsatz [Bor48b]. Hayes [Bor53b]. Health [KED^{+}60]. Heat [BB42a, BvK13b, Bor13a, Bor14c, Bor20l, BB21c, BB22a, BK23, BF25a, BB42b, BB63g, BB63i, BvK63c, Hab19]. heavenly [vS21]. Hedwig [Bro72, For70, Kle70a, BB69a, BB05]. Hedwing [Alb71]. Heidelberg [Anoxx, Kle70c, Anoxx]. Heinrich [Bor46g, Bor46h, Bor63-34]. Heisenberg [Bor53b, Bro72, KED^{+}60, Mar73, Sch05, Sch07, BDR84, Bor49i, Hei72, Sch05, Sch07]. Heitler [Bor46c]. helium [BH23a, BH63a]. Heliumatom [BH23a, BH63a]. Hellinger [Anoxx, Anoxx]. Herkner [Bor18c]. Hermann [Bor10b, Bor59b, Bor63v, Bor10b, Bor59b, Bor63v, Wei12]. Herneck [Hei72]. Herrn [Bor13d, Bor15a, BL19]. hexagonal [Bor42e]. high [BB21c, BB22a, Bor47g, BB63g, BB63i]. higher [HP82, HP02]. Hilbert [Som05, Bor22b, Bor63x, Som05]. Hilfe [BL18b, BL63b]. Hintertreppe [Fis10a, Fis12a]. Historical [Jam74b, Kon78]. Histories [Pei97]. History [Ano02, BN13, Bor54a, Bor02, WP85, HP02, Inf55]. Hitler [Hag08, MP01]. hoffen [Bor63-58]. Hoffnung [Bor66c]. hohen [BB21c, BB22a, BB63g, BB63i]. Holl [Lin98]. homöopolaren [Bor32e]. homopolar [Bor31a, Bor32e]. Hon [BS63a]. Hope [Bor64g, Bor63-58, Bor66c]. Horace [MPK^{+}54]. Hrsgg [Krö98]. Huang [Lax55]. Hull [CZPS37]. Hydratationswärme [Bor20l]. hydration [Bor20l]. Hydrogen [BFL46, DM69, Bor22d, BK23]. hyperpolarizabilities [BS85]. hypothesis [Ali05, Bor13a].

Ice [Bor46e]. ideal [Bor51e]. idealen [Bor51e]. ideas [Lor10, BB63a]. if [Nik08]. Ignited [Bey06, Cas05, Gre05a]. ihre [Bor19b, Bor20d, BB21b, BB22a, Bor22a, Bor51e, BB63g, BB63h]. ihren [BS19, BS63b]. ihrer [Lem82]. II [Bor26c, BJ30, Bor36a, BS16, BB20c, BB22b, Bor26a, Bor26c, BHJ26, BJ30, BI35, Bor35c, BN36b, Bor36e, BF40c, BB43c, BP44b, BC48d, BHJ63, BB63j, Bor65g, Bor65j, BHJ67, BGĬ90, Man19, Mis40]. III [BF40d, BF40e, Bor42b, BP46, BG47a, Bor65h]. Illus [Mor54, Kle70c]. im [Bor10e, BH23a, Bor31d, Bor57e, Bor58e, Bor59d, BH63a, Bor66e, Bor83, Lin58]. Imperfect [RS51]. importance [Bor28e]. Impression [Bor79]. Impuls [Bor14a, Bor63k]. Impuls-Energie-Satz [Bor14a, Bor63k]. Inaugural [Bor36c]. Including [De 16]. index [Bor18e, Bor63-50]. Indian [Sur99]. inertial [Bor09c, BL58, BL63d]. Infeld [BB99]. Inference [AG20]. Infinitesimal [Bor13b]. Infinitesimalrechnung [Bor13b]. influence [BS19, BH24a, Bor24a, BH63b, BS63b]. infra [Bor18f, Bor32a, Bor63o, Bor63-53]. infra-red [Bor18f, Bor32a, Bor63o, Bor63-53]. inmediato [Bor69d]. Innovator [Bak83]. Instrumente [Bor32c]. instruments [Bor32c]. interaction [BP46]. interactions [Pei71]. Interference [BW59, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, Hec00, Wol66, Str64, Str60]. interferometric [BB21a]. interferometrische [BB21a]. Interpretation [Bor53d, Bor55g, Kon78, Pai82, Bel90, Bor24e, Bor32d, Bor54c, BB55, Bor55d, Bor61a, Bor63e, Bor63m, dB53]. Interpretations [Jam74a, Jam74b]. intimate [Far79]. Introduction [Bro72, KED^{+}60, Bor22c]. Investigations [Bor63-46, Bor43c]. ion [Bor20j, Bor20k]. Ionen [BH24a, BH63b, Bor20j, Bor20k, Bor20l, Bor63-49]. Ionendeformation [Bor24a]. Ionenkristalle [BM63]. Ionenkrystalle [BM32]. Ionenladung [Bor18e, Bor63-50]. ionic [Bor18e, BM32, BM63, Bor63-50]. ions [Bor20l, BH24a, Bor24a, Bor63-49, BH63b]. Irene [Bro72]. ISBN [Sho20]. Ismen [Bor60a]. isms [Bor60a]. Isotopes [MPK^{+}54]. isotroper [Bor18b]. Isotropic [Bor35c, Bor18b, Bor36a]. Italian [Bor60b, Bor61b, Bor71, HSBA60]. IV [BM40, BG47b]. Ivan [Ber04]. IX [Bor44c, Bor42e, Bor63i, Bor63-32].

J [Ano52, Ber06, Bla58, Bor40d, Bor40h, Bor41b, BB63a, Bor63-40, Fow27, MPK^{+}54, vS21, Ber04]. J.-P [Bla58]. Jacob [Bor62c]. Jaffe [Hei72]. Jahrbuch [Bor40c]. Jahre [BG61]. Jahrhundertfeier [Bor31d]. Jahrhunderts [KED^{+}60]. Jahrhunderts/Bellman [KED^{+}60]. James [Lem82, BW64a, K^{+}83, Lem82, vLP52]. Jan [CZPS37]. Jeremy [Ber04]. Jew [Sto98]. Jewish [Hag08]. Johannes [Bor52b, Bor63b]. Jones [Bor37a]. Jordan [Sch05, Sch07, FP09, Sch05, Sch07]. Joule [Bor50f, Bor50g]. Journalistic [Lem83]. Journey [FR13]. Jude [Sto98]. Julius [Ber15, Ber15]. jump [BBK13]. Jungk [Hei72]. Justified [Bla67]. justify [Dre17].

Karl [Bor15a, Bor48a, Bor63-28]. Katalog [Lem83]. Keil [BG61]. Kelvin [Bor38f]. Kernbewegung [Bor51g, Bor63y]. Kernzerfalls [Bor29d, Bor29e, Bor63-62]. killing [HP82, HP02]. kinematic [Hei25]. kinematics [Bor09a, Bor09e, Bor09f, Bor09b, Bor10c, Bor63f, Bor63n, Bor10e]. Kinematik [Bor09a, Bor09e, Bor09f, Bor09b, Bor10c, Bor10e, Bor63f, Bor63n]. kinematischer [Hei25]. Kinetic [BG47c, BG49, BGK50, BG63, Bor13e, BG46, BG47a, BG47b, BG48]. kinetischen [Bor13e]. klassische [Bor55f]. klassischen [Bor58h, Bor59e]. Klaus [Bon05]. Klecksel [Bus65]. kleinsten [Fis10a, Fis12a]. Knickformel [Bor09g]. Kohäsionskräfte [Bor19d, Bor19e]. Koller [CZPS37]. Kompressibilität [BL18c, BL63c]. Kongress [Bor28b]. Konstanten [BH24a, BH63b, Bor24a]. kontzientzia [BB90]. Kopplung [Bor51g, Bor63y]. Körper [BB21c, BB22a, BB63g, BB63i, Bor10a, Bor10d, Bor14b, Bor19d, Bor19e]. Körpereigenschaften [Bor27a]. Körpern [Bor10b, BL11]. Körpers [Bor10c, Bor10e]. Kräfte [Bor31f, Bor24e, Bor29c, Bor63-60]. kräftefreien [BL58, BL63d]. Kristalldimensionen [Bor18d]. Kristalle [BL18c, BS19, Bor51e, BL63c, BS63b, Bor15e, Bor18a, Bor18f, Bor22e, BGM33b, BGM33a, Bor59c, Bor63-52, Bor63-53]. Kristalleigenschaften [BL18b, BL63b]. Kristallen [Bor15c, Bor18e, Bor51g, Bor63y, Bor63-50]. Kristallgestalt [BS19, BS63b]. Kristallgitter [BL19, BB22b, Bor23a, BB63j, Bor15b, BL18a, Bor21d, BL63a, Bor63-63, Som15]. Kritische [Bor20g, Bor20h, Bor20i, Bor63z]. kubische [Bor18h]. Kun [Lax55].

L [BL19, Bor48c, CZPS37, Hol88, MPK^{+}54]. Laboratory [Sur99]. Last [Bor51h]. later [Bor50d]. Lattice [Bor42c, BL43, Sur99, Bor14d, BL18c, BL18a, BL19, BB19, Bor19c, BB20b, BB20c, Bor21c, BG21a, BB22b, BB25a, BB25b, Bor26a, Bor26c, BM32, BGM33b, BGM33a, BT34, BF40e, Bor42e, BB43c, BL63c, Bor63-55, BM63, BB63e, BL63a, Bor63-61, Bor63s, Bor39c]. Lattice-theoretical [Bor39c]. Lattices [BH54, Bor63i, Bor63-31, Bor63-32, Bor63-47, BB63j, BK12, Bor15b, Bor21d, Bor23a, Bor40e, BF40e, BM40, Bor42e, Bor43d, BB43c, Bor44c, BvK63b, Bor63-63, Bor64d, BH88, Mis40, Som15, Lax55]. Laue [Bor59c]. Laues [Bor59c]. launched [Cha05]. laureate [Bor78, Cha05, Bor75, Sto98, And81, Far79, Han80, Hei79a, Sch79, Gow79, Hei79b]. laureates [ABH^{+}55, ABH^{+}87]. Law [Bor48b, Bor49e, BC13b, BC13a, Bor07]. Laws [Bor38f, Bor39b, BW63b, BW26b, BW26a, Bor39a]. layers [BB21a]. leap [Fis10a, Fis12a]. Leben [Bor75, HH70]. Leccia [Mar73]. Lecture [Bor36c, Bor38f, Bor47c, Bor53c, Bor09d, Bor50f, Bor50g, Lor10, BB55]. Lectures [Bor33a, Bor40c, Bor49a, Som34, Bor26a, BJ30, Bor49i, BH25, Bor53b]. Left [Bor64g, Bor63-58]. Legendre [Bow69, Che70]. Lehrbuch [Bor33c, Bor40c, Bor43b, Bor53e, Bor65d, Bor72]. Léon [CS79a]. Letter [BI33a, BI33b, BR57, Hab22, Tho22]. Letters [BR95, Bor47a, Bor69b, Bro72, BB71a, BB71b, BB05, Ein91, Som05, Bre73]. Levinson [SG82]. Levinson-modified [SG82]. L’expérience [Bla58]. Library [Kle70c]. Lichtes [Bor17, Bor18g, BG21b]. Lichtfortpflanzung [Bor12a]. Lichtquantum [Bor55a]. Lichtstrahls [vS21]. Lichttheorie [Bor33c, Bor43b, Bor53e, Bor65d, Bor72]. Life [And81, Bey06, Bor68a, Bor75, Cas05, Far79, For68, Han80, Hei79a, Hei79b, Kle70b, Oes68, Sch79, Bor68d, Bor78, Gow79, Gre05a, HH70, Sig96]. Light [BW59, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, GS47, Bor17, Bor18g, BG21b, Bor33c, BN36a, BN36b, Bor43b, Bor53e, Bor55a, Bor65d, Bor72, vS21, Bor12a, Hec00, Wol66, Str64, Str60]. limiting [Bor06]. limits [Bor51e]. line [Bor06, Bor63-43, vS21]. linéaire [Bor38e]. Linear [Ano10]. linearization [Bor36b]. Linie [Bor06, Bor63-43]. link [BK23]. L’interprétation [dB53, BB55]. liquid [Bor63-54]. Liquids [BG47c, BG47d, BG49, BGK50, BG63, Bor15d, Bor18b, BG46, BG47a, BG47b, Bor63-51]. List [Bor63-57]. Literary [BR95]. Literatur [Lin98]. Localizable [Bor50c]. loft [Som05]. London [Bro72, Bor31d]. Long [BFL46]. Looks [Bor50d]. Lorentz [Bor28a, Bor61c, Bor63a]. lost [GXB21]. Ludwig [Bor48a, Bor63-28]. L’Unité [CZPS37, CZPS37]. L’universo [Bor60b]. Luxury [BB69b, Lem82]. Luxus [BB69b, Lem82].

M [And81, Ano52, Bor47b, Bor47f, Bor63-35, Sho20, Som15, Som34, Som50]. Machine [Hei72]. Macmillan [Bro72]. made [Haw11]. Magnetic [BS50, Bor36d, Bor36e]. main [Bor55c]. making [Bel99]. Maler [Bus65]. Man [Bor57b, Bor63-27, Bro73a, Bor64a]. Mandl [KED^{+}60]. Manhattan [HHW99]. Manifesto [Ano06]. Manner [Bor65i, Sch05]. Maria [Goe88]. Marley [KED^{+}60]. Mass [Bor35d, BF40a, BF40b, Bor09c, Bor18a, BI33a]. Masse [Bor09c, Bor18a]. Masses [BR49, BGCR49]. Master [Cha05]. materialism [Fre01, FL10]. Materie [Bor20b, BF31, BB27, Bor33a, Som34, da 33, Bor13e, Bor19f]. Mathematics [Bor20f, RSS^{+}07, Ber15, FF91]. Mathematik [Ber15, Bor20f]. Mathématique [Bor40c]. Mathieu [Bla58]. Matière [CZPS37, Bor52d]. Matrix [BHJ62, KED^{+}60, BC80]. Matrizenmechanik [BHJ62]. Matter [BF31, Bor33a, Som34, Bor13e, Bor19f, Bor20b, Bor23b, BB27, Bor36d, Bor36e]. Max [A.53, Alb71, And81, Ano46, Ano56, Ano70, Ano02, ABdB^{+}53, Ber50, Ber06, Bey06, Bla58, Bro72, Bro73b, Cas05, CS83, CZPS37, Cop61, Din50, Din23, Ein91, Far79, Fis10a, Fis12a, For68, For70, Fow27, Gow79, Han80, Hec00, Hei72, Hei79a, K^{+}83, Kle70b, Kle70a, Kle70c, Kör59, Lax55, Lem83, Lin98, Lin58, Mar50, Mar73, McC49, Mil60, Mil49, Mor54, MPK^{+}54, Nor71, Oes68, Pag63, Pry60, Rom65, Sch05, Sch07, Sch79, Seg64, Str64, Str60, Süs57, Wol66, da 33, Ali05, Ano63, Ano82, Anoxx, Ber05, BR95, Bor10b, BvL23, Bor48a, Bor48c, Bor57c, Bor59c, BvL63, Bor63-28, Bor63-57, Bor64c, Bor65f, Bor65g, Bor65h, Bor68b, BB71a, BB71b, Bor79, Bor02, BB05, CS79b, Coo02, Fis10a, Fis10b, Fis12a, Fis12b, Fre01, FL10, Giu13, Gre05a, Gre05b, Gre06]. Max [GE08, Hai12, Hei79b, Hei62, KS71, Kle70c, Lem82, Lem83, Mat78, Pai82, Paw08, Ros79, Sch05, Sch07, Wei12, Wil67, Wol83, Wol95]. Maxwell [Hei72, Bor18e, Bow69]. Maxwellian [Bor63-50]. Maxwellsche [Bor18e, Bor63-50]. Meaning [Bor50b, Bor28d, Bor29b]. Measurement [KED^{+}60, Bor20e, Bor63r]. Measurements [De 16]. Mécanique [BH84, BH28, Bor31e, BB55, Bor59a, dB53]. Mechanical [BB63g, Bor19f, Bor21b, BB21b, BB22a, BP22, Bor63-54, BP63a, BB63h, BC80, Hei25]. Mechanics [Ano10, BH27, Bor27b, BF38, Bor40c, BP44c, Bor46j, Bor53d, Bor55g, Bor60c, BHJ62, Bor63j, BW63a, Bor63-37, Bor63-44, Bor67c, BHJ67, BH75, BJ77, Bor77, Fow27, GGK02, Jam74b, Man18, Man19, Pai82, SGM^{+}12, vdW67, Bor24f, BJ25a, BH25, Bor26b, BHJ26, Bor26e, Bor26d, Bor26f, Bor27c, Bor28e, BH28, BJ30, Bor31b, BW31a, BW31b, Bor31e, BF33, Bor37c, Bor37d, BP44a, BP44b, BP46, BG47b, Bor48b, Bor48d, Bor55f, BB55, Bor55d, BL58, Bor58h, Bor59a, Bor59e, Bor61a, Bor63e, Bor63-56, BJ63a, BHJ63, BL63d, Bor63-59, Bor63-36, BF63a, Bor63m, Bor63-64, Bor67a, BJ67, BH84, BGĬ90, FP09, Gal01, GXB21, dB53, vdW07, Bor46c]. Mechanik [Bor55f, Bor58h, Bor59e]. mechanischen [Bor19f, Bor21b, BB21b, BB22a, BB63g, Bor63-54, BB63h]. mechanischer [BP22, BP63a, Hei25]. Medaille [BEvL^{+}28]. Medal [BEvL^{+}28]. media [Bor12a]. Medien [Bor12a]. Mehlin [KED^{+}60]. mehratomigen [Bor32e]. mehratomiger [BH23c]. mehrfach [BH55]. Mein [Bor75, HH70]. Meine [Wei82]. meiner [Bor10a, Bor57e, Bor58e, Bor59d, Bor66e, Bor83, Lin58]. mele [Bor69f]. Melting [Bor39e, Bor40g, Bor63-48, Bor39c]. Mem [Bor47f]. Memo [BBD^{+}46]. Memoirs [And81]. Memorial [Bor50f, Bor50g]. memoriam [BBB50]. Memories [Bor56a, Bor59b, Bor59c, Bor63u, Bor63v, Bor65b, BI67, Bor68c, BI69, Bor69a]. men [Sch05]. Mensch [Bor64a]. Meson [Bor44f, Bor47e, BR49, Bor47d]. Message [Bor64c]. Messung [Bor20e, Bor63r]. mesure [Bor59a]. Metals [Bor37a]. metaphysics [Bor50f, Bor50g]. method [Bor14c, Bor18e, BB21a, Bor63-50]. Methode [Bor14c, Bor18e, BB21a, Bor63-50]. Methods [BS50, Bor63-47, Bor43d]. Mie [Bor14a, Bor63k, Bor07, SG82]. migration [AS96]. mind [Cha05]. Minimum [Bor39b, Bor39a]. Minkowski [Bor10b, Bor59b, Bor63v, Bor10b, Bor59b, Bor63v, Wei12]. Missing [Joh13]. Mitchell [Krö98]. Mitteilung [Bor15a]. mobility [Bor20j, Bor20k, Bor63-49]. Mobilization [Joh13]. Model [Khr09, Bor09d, BL18a, BL19, Bor22d, BH23b, BL63a]. Modell [Bor22d]. Modellen [BH23b]. Models [BL63b, Joh13, BL18b, Bor18h]. Modern [Bor54a, CZPS37, Bor20b, Bor36c, Bor37b, Bra09, Fre01, FL10, HSBA60, HBSA61, Bor33a, Som33]. moderna [HSBA60]. moderne [Bor20b, Bor33a, Som33, da 33]. Modernizing [Joh13]. Modified [Bor33b, SG82]. Molecular [BG63, Joh13, BF25b, BG46, BF63d, BBK13]. molecule [Bor22d]. Molecules [Bor36a, Bor17, Bor21b, BH23b, BH23c, BH24b, BJN25, BO27, Bor31a, Bor32e, Bor51g, Bor63-54, Bor63y, BO63, BO98, Bor35c]. Molekelbildung [BF25b, BF63d]. Molekeln [Bor51g, Bor63y, Bor17, BH23b, BH23c, BH24b, BO27, BO63]. Molekularströmung [BBK13]. Molekülen [BJN25, Bor32e]. moment [BK23, Bor36d, Bor36e]. Momentum [Bor07]. Monde [CZPS37]. Monographs [BF31]. Moon [Bor66d]. Morgan [KED^{+}60]. Moseley [Hei72]. most [Haw11]. Motion [Ily16, Bor51g, Bor63y, vS21]. Motiv [Bor55c]. Mott [Bor37a]. Movement [Gre05b]. movimento [Bor60b]. Moving [Bor10a, Bor10d, Wei72, Wei85, Bor10b, Bor12a, Bor13d, Bor14b]. Mr. [Bor13d, Bor15a, BL19]. MSS [Goe88]. multi [BH55]. multi-periodic [BH55]. multiply [BH56]. multiply-periodic [BH56]. Munich [Kle70a]. Mutually [SGM^{+}12]. My [Bor56c, Bor57e, Bor58e, Bor59d, Bor66e, Bor68a, Bor68d, Bor69f, Bor69g, Bor70a, Bor70b, Bor78, Bor83, For68, Hei72, Kle70b, Kle70c, Nor71, Oes68, Pry60, Bor10a, Bor50d, Bor64d, Süs57, Bor57d, Bor75, HH70, Wei82, And81, Far79, For68, Gow79, Han80, Hei79a, Hei79b, Kle70b, Oes68, Sch79]. mysterious [Bor35b]. Mystery [FR13].

N [Bor18g, Bor37a]. Nachlass [Bor10b, Lem83]. Nachweis [Bor21b, Bor63-54]. nähe [vS21]. Nancy [Cas05, Bey06]. narrative [Com56]. National [GR63, RW94]. Natur [Bor19d, Bor19e, Bor21a]. Natural [ABdB^{+}53, Bor36c, Bor39b, Bor49f, Bor63o, BvK13a, Bor14c, Bor15d, Bor18b, Bor18f, Bor22e, Bor32a, Bor39a, Bor63-51, Bor63-53, BvK63a, Som50, Ber50, Din50, Mar50, McC49, Mil49]. naturalist [Bor57a]. Nature [Bor38f, Bor19d, Bor19e, Bor21a, Bor55e]. Naturforscher [Hei72]. Naturforschers [Bor57a]. naturliche [Bor22e, Bor15d, Bor63-51]. natürlichen [Bor18b]. Naturwissenschaft [KED^{+}60]. Naturwissenschaft/Heisenberg [KED^{+}60]. Naturwissenschaften [Bor65i]. Naturwissenschaftlers [Bor65k]. naturwissenschaftliche [Bor65i]. Nazi [Joh13, MP01]. Nazis [Cou13]. near [Bor69d]. necessity [BKCK58]. Needs [Bla67]. Nernst [Som05, Som05]. neue [BW26a, BW63b, Lor10]. neutral [Bor20e, Bor63r]. neutraler [Bor20e, Bor63r]. neutrino [BN36a, BN36b, Bor37d]. neutron [Bor37d]. nicht [Bor65i]. nichtperiodische [BW26a, BW63b]. Ninth [Bor38f]. No [Enz02b]. Nobel [And81, Bey06, BB55, Cas05, Far79, Gow79, Han80, Hei79a, Hei79b, Sch79, ABH^{+}55, ABH^{+}87, BB55, Bor75, Bor78, Cha05, Gre05a, Sto98]. Nobelpreisträger [ABH^{+}55, ABH^{+}87, Sto98]. Nobelpreisträgers [Bor75]. noch [Bor63-58]. Non [Bor50c, BW63b, Bor38e]. non-linéaire [Bor38e]. Non-Localizable [Bor50c]. Non-Periodic [BW63b]. Nonlinear [BB99, Bor38e]. nostro [Bor61b]. Note [Ano67, Dem03, Bor32b, BT34, Bor63q]. Notwendigkeit [BKCK58]. nouveau [Bor52d]. nouvelle [BI34e]. Nuclear [BBF^{+}57a, Bor47e, BY50, Bor53b, ABH^{+}55, ABH^{+}87, Bor29d, Bor29e, Bor47d, Bor51g, Bor55e, BKCK58, Bor63y, Bor63-62, Kae48, Koz91]. nuclei [Bor38a]. Nucleus [FR13]. Number [Bor63-38, Bor35b, Bor39d]. Numbering [Hei72]. Nymphenburger [Kle70a].

O [Bor18g]. O.M [Bor40d, Bor63-40]. Oberflächenenergie [BS19, BS63b]. Obituaries [Bor42d]. Obituary [Ano70, Bor40a, Bor40d, Bor46g, Bor47b]. objectivity [Bor66f]. Objektivität [Bor66f]. occurring [BB43c]. Old [Lor10]. Once [Ano67]. ondulatoire [dB53]. One [Lap69]. One-Dimensional [Lap69]. Operations [BW63b]. Operator [GXB21]. Oppenheimer [Ber04, Ber04, BS85, Bow08, DM69, GDR04, Pac10, PST07, Ber04]. Optical [Bor35c, Bor15d, Bor15e, Bor18b, Bor22e, BH24a, Bor24a, Bor32c, Bor36a, BH63b, Bor63-51, Bor63-52, Str64]. OPTICS [CS83, BW59, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, Hec00, KED^{+}60, Str64, Str60, Wol66, Bor33c, Bor43b, Bor53e, Bor65d, Bor72, Cop61, Mil60, Sho20]. Optics/Mandl [KED^{+}60]. Optik [Bor33c, Bor43b, Bor53e, Bor65d, Bor72]. optische [Bor15d, Bor15e, Bor22e, BH24a, BH63b, Bor63-51, Bor63-52]. optischen [Bor18b, Bor32c]. orbits [BH63a]. order [Bor49h]. oscillations [Bor14c]. Oscillators [Bor35c, Bor36a]. Otto [Bor40a, Bor63-33, Bor81, Bor40f]. Our [Bor61b, BL19, BB22a, Bor62a, BB63g, Lor10].

P [Bla58, HKP^{+}32a, HKP^{+}32b]. Pacifist [Sal19]. painter [Bus65]. Paper [Kle70c, BI34c, Bor44c, FP09, Fre01]. Papers [BDR84, CS79a, Goe88, Mor54, MPK^{+}54, ABdB^{+}53, BB22a, BB63g, Haw11]. paradox [Bor58c, BB58, BB63b]. Paradoxes [Joh13]. Part [Bor63-38]. particle [Bor36d, Bor36e, BL58, BL63d]. Particles [Bor49g, Khr09, Bor49c, Bor65e, Fis10a, Fis12a]. Particular [BB63f, BB47b, Bor49h]. partir [BI34d]. Partisans [Sal19]. Partly [BB63a]. Partly-Baked [BB63a]. Pascual [Sch05, Sch07, Sch05]. passes [vS21]. Passion [Jon08]. past [HH70]. path [Bor20e, Bor46f, Bor63r]. Pauli [Bor47d, Enz02b]. Peace [Gre05b, Sal19, Bor47c]. Peaceful [Sal19]. Pedagogy [BN13]. period [Bon05]. Periodic [Bor46k, BW63b, BW26b, BW26a, Bor46b, BH55, BH56]. periodische [BW26a, BW63b]. periodischer [BH55]. personal [Com56]. Perspective [Jam74b]. perturbation [Bor23c, Bor23d]. peut [Bor59a]. peut-elle [Bor59a]. Phase [Cas12, BH23b]. Phase-Spaces [Cas12]. Phasenbeziehungen [BH23b]. phenomena [BW26b]. Philosopher [Sch49a, Sch49b]. Philosopher-Scientist [Sch49a, Sch49b]. Philosophic [Bor53b]. philosophical [Bor36c, Bor37b]. Philosophie [Bor65i, KED^{+}60]. Philosophie/Frisch [KED^{+}60]. philosophieren [Bor65i]. Philosophize [Bor65i]. Philosophy [ABdB^{+}53, Ber50, Bor36c, Bor49f, Bor65i, Din50, Jam74b, Mar50, McC49, Mil49, Som50, Fre01]. Photo [BFP45, BFP63]. Photo-Electric [BFP45]. Photochemical [BS50]. Photoelectric [Ano46]. Photon [Bor44f]. Phys [Ber06]. Physical [Bor19b, Bor27b, Bor53c, Bor53f, Bor20d, Bor21a, Bor22a, Bor27a, Bor28d, Bor29b]. Physicist [And81, Bey06, Bor50d, Bor65f, Bor65g, Cas05, Bor68c, Gre05a]. Physicists [BBF^{+}57a, Sal19, Lem82, Bor28b]. Physics [BN13, Bor12b, Bor33a, Bor39b, Bor45a, Bor50e, Bor50f, Bor50g, Bor51h, Bor56c, Bor58a, Bor60d, Bor62d, Bor62e, Bor65c, BBSR69, Bor69e, Bor69g, Bor70a, Bor70b, BRB71, Bre70, CZPS37, De 16, HMLM88, KED^{+}60, MPK^{+}54, RSS^{+}07, Sig96, Ber15, Bon05, Bor20c, Bor22b, Bor22c, Bor31c, Bor35a, Bor36c, Bor37b, Bor39a, Bor40b, Bor42h, Bor43a, Bor44d, Bor45d, Bor46a, Bor49i, Bor51c, Bor51f, Bor53c, Bor54b, Bor56b, Bor58b, Bor62a, Bor63x, Bor71, BB05, Bra09, Con62, FF91, Haw11, HSBA60, HBSA61, Kae48, Koz91, Lor10, Meh75, Str11, WP85, Wei72, Wei85, Bor56d, Bor57e, Bor58e, Bor59d, Bor60a, Bor60e, Bor61b, Bor66e, Bor69f, Bor83, Sch07, Ano56, CZPS37, Kör59, Som33, Hei72, Kle70c, Nor71, Pag63, Pry60, Süs57, Fre01, FL10]. Physics/Churchman [KED^{+}60]. Physik [Ber15, Bor31c, Bor33a, Bor40c, Bor60a, Bor62a, KED^{+}60, Lor10, da 33, Bor12b, Bor20c, Bor22b, Bor40c, Bor54b, Bor56d, Bor57e, Bor58e, Bor58b, Bor59d, Bor60e, Bor63x, Bor66e, Bor69e, Bor83, Sch07, Som33, Lin58]. physikalische [Bor21a, Bor24a]. physikalischen [Bor19b, Bor20d, Bor22a, Bor28d, Bor29b]. Physiker [Lem82, Lin98, Bor28b]. Physikers [Bor68c]. Physique [CZPS37, Bla58]. Pi [Ber04]. Pierre [Mar73]. Pieter [BS45, BS63a]. pilot [Dre17]. Planck [Bor63-28, Fis10a, Fis12a, Som05, BEvL^{+}28, Bor48a, Bor48c, Bor57c, Fis10a, Fis12a, Som05]. Planck-Medaille [BEvL^{+}28]. plane [Bor06, Bor15c, Bor63-43]. Plastic [MPK^{+}54]. poem [Bor62f]. Pohl [HMLM88]. point [BvK13a, Bor52d, BvK63a]. poisonous [Hab86, Hol88]. polar [BT34]. Polarisationsebene [Bor15c]. polarization [Bor15c]. political [Sch05, Sch07]. Politics [Bor60d, Bor60e, Bor62d, Bor62e, BB05, Fre01, Sch07, Pag63]. Politik [Sch07, Bor60e]. politische [Sch05, Sch07]. polyatomic [BH23c, Bor31a, Bor32e]. Portrait [Ber04, Pai00]. Portuguese [Bor65c, BRB71]. Potential [DL08, Lap69, Pac10]. potentials [Bor21c, Bor63-55]. power [Bor18b]. Pp [Bro72, Kle70b, Kle70a, Kle70c, Mor54, Sho20]. predict [Bor59a]. Predictability [Bor58h, Bor59e]. prédire [Bor59a]. Prentice [Ber04]. Presented [Mor54, MPK^{+}54, ABdB^{+}53]. Press [Ber04, Krö98, Sho20]. Press/Prentice [Ber04]. principe [Bor49e]. Principes [BI34e]. Principle [Bor63j, Bor26b, Bor48d, Bor49c, Bor09a, Bor09e, Bor09f, Bor09c, Bor09b, Bor10c, Bor10e, Bor13d, BR49, Bor63f, Bor63n]. Principles [Bor39b, BW59, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, KED^{+}60, BO07, Bor39a, Bor12b, BI34e, Cop61, Hec00, Mil60, MPK^{+}54, Sho20, Str64, Str60, Wol66]. Pringle [Ano46]. Prinzipien [Bor12b]. prinzips [Bor10c]. Probabilistic [Kon78, Bel90]. Problem [BF30, BF63c, Bor24b, Bor62a, Bor24b]. Problematik [Bor62a]. Probleme [Bor26c]. problèmes [Bor31e]. Problems [Bor26a, Bor53b, Bor70c, Bor31e, Bor26c]. Processes [Bor77, Bor10b, BJ25b, BW26a, Bor26e, Bor26d, Bor26f, Bor28e, Bor39c, BJ63b, Bor63-59, Bor63-36, Bor63-64]. Produktion [Lin98]. Prof. [Bor40a, Bor46g, Bor47b, Bor63-33]. Professor [Bor36c, Ano70, Bor46h, Bor47f, Bor63-34, Bor63-35, Bor81]. Profs. [Ano52].

Program [Ure67]. progress [Bor44d]. Project [HHW99]. Projective [De 16]. Prolog [Bor67b]. Promise [RSS^{+}07]. Proof [Man18, Man19, Bor63-54, BF28, BF63b]. Propagation [Bor46k, BW59, BW64b, BW65, BW69, BW75, BW^{+}80, BW97, BW99, BWB19, Hec00, Wol66, Bor12a, Bor46b, Str64, Str60]. Properties [Bor37a, BL63b, BL18b, Bor27a, BG47a]. prospects [Bor53c]. Proton [Bor35d]. publizistische [Lem83]. Pulse [Bor63k, Bor14a]. Pulse-Energy [Bor63k, Bor14a]. Punktgittern [BvK13a, BvK63a]. pure [BP44b, BP44a]. purely [BF40e]. Purpose [Bor39b, Bor39a]. Putnam [MPK^{+}54]. Pyroelectricity [Bor45c, Bor63-29, Bor22f]. Pyroelektrizität [Bor22f].

Quanta [BH84, BH28, Bor49e]. Quantelung [BB22a, BP22, BP63a, BB21b, BB63h]. Quantelung” [BB63g]. Quantenelektrodynamik [BR31]. Quantengesetze [BW26a, BW63b]. Quantenhypothese [Bor13a]. Quantenmechanik [BHJ26, Bor26e, BJ30, BF33, Bor48b, BL58, BHJ63, BL63d, Bor63-59, BF63a, Bor24f, BJ25a, Bor26b, Bor26d, Bor27c, Bor28e, Bor31b, BW31a, BW31b, Bor55d, Bor61a, Bor63j, Bor63e, BW63a, Bor63-56, BJ63a, Bor63-37, Bor63-36, Bor63m]. Quantensprung [Fis10a, Fis12a]. quantentheoretische [Hei25]. Quantentheorie [BH23c, BH24b, BJ25b, BO27, Bor29c, Bor30b, BJ63b, Bor63-60, BO63, Bor23c, Bor23d, BF25b, BF63d, LBBH69]. Quantenwelt [GE08, Gre06]. quantique [Bor31e, BI34d, BI34e, BB55]. Quantisation [BI63b, BI34b]. Quantised [Bor35d]. Quantization [BB63g, BB21b, BB22a, BP22, BW26b, BI35, BP63a, BB63h]. Quantum [Ano10, BN13, Bel99, Bey06, Bor27b, Bor30a, Bor34c, Bor34b, Bor35e, Bor38b, Bor40c, BLS42, Bor42g, BP44a, BP44b, Bor45c, BP46, BG47d, BGCR49, Bor53d, Bor55g, Bor63j, BW63b, Bor63-36, Bor63-30, Bor63-29, Bor63-45, Bor67c, BHJ67, BH75, BJ77, Bor77, BGĬ90, Cas05, De 16, GGK02, Ily16, Jam74b, Joh13, Jon08, Khr09, KED^{+}60, Man18, Man19, Pai82, SGM^{+}12, vdW67, Ber05, Ber06, Bor13a, BH23c, Bor24f, BH24b, BJ25b, BJ25a, Bor26b, BW26a, BHJ26, Bor26e, BO27, Bor28e, Bor29c, Bor29a, BJ30, Bor30b, BR31, Bor31b, Bor31e, BF33, BI34d, Bor38d, Bor48d, Bor49d, Bor49e, Bor55a, BB55, Bor55d, BL58, Bor61a, Bor63e, BJ63b, Bor63-60, Bor63-56, BJ63a, BHJ63, BL63d, BO63, Bor63-59, BF63a, Bor63m, BJ67, BB68, BO98, BC80, Con62, FP09, Fis10a, Fis12a, Gal01, GXB21, Gre05a, Gre06, GE08, Haw11, Hei25, LBBH69]. quantum [Str11, vdW07, BV09, Bor23c, Bor23d, BF25b, Bor26d, Bor27c, BW31a, BW31b, BG47b, Bor48b, BW63a, BF63d, Bor63-37]. quantum-mechanical [BC80]. quantume [BI34e]. Qubits [De 16]. quelle [Bor59a]. Quelques [Bor31e]. quest [Com56]. question [Bor55h]. questions [Lor10].

R [Ano46, Ber04, Bor47f, CZPS37, Fow27, HKP^{+}32a, HKP^{+}32b, MPK^{+}54]. race [Bor66d]. Radiation [Bor33a, Bor13a, BB33, BF39a, BF39b, BB63c, Som34]. Radioactive [MPK^{+}54]. Radium [Kae48]. radius [Bor32b, Bor33b, Bor63q]. Raman [BB63f, BB42a, Bor28f, BB42b, BB45, Bor46i, Bor47g, BB47b, Bor49h, Sur99]. Raman-Effect [BB63f]. Ramaneffektes [Bor28f]. Random [Khr09]. Ranging [Ano02, Bor02]. Raum [Bor06, Bor13c, Bor63-43]. Raumfahrern [Bor58c]. Raumgittern [BK12, BvK63b]. Raumgittertheorie [Bor14d, Bor63-61]. Rausch [Bor46g, Bor46h, Bor63-34]. Ray [Bor63-46, Bor42c, BLS42, Bor42g, Bor43c, vS21]. Rays [BB43a, Bor34a, Bor41a, BB47a, Bor21a, BS41, Bor42b, BL42, BB43b, Bor59c]. reaction [BF25a]. Read [MPK^{+}54]. Reader [Bor58f]. Reaktionswärme [BF25a]. Realitätsbegriff [Bor58b]. Reality [Bor53f, Bor58a, Bor63h, Bor64e, Bor64f, Bor65j, Bor65i, Bor55b, Bor58b, Bor66a, Bro73b]. really [Bor55f]. realm [Kae48]. reason [Bor55c, Bor66c]. récentes [Bor24c]. Recherches [Bor24c]. Reciprocal [Cas12, BF40d]. Reciprocity [Bor39d, BF40d, BF40c, BCG49, Bor49g, Bor50c, Bor63-38, Bor65e, DL08, Bor38a, Bor38c, Bor48d, Bor49c, BR49]. Recollections [And81, Bor65f, Bor65g, Bor65h, Far79, Gow79, Han80, Hei79a, Hei79b, Sch79, Wol83, Wol95, Bor53a, Bor78, Bor75]. reconsidering [BV09]. red [Bor18f, Bor32a, Bor53g, Bor54c, Bor63o, Bor63-53]. Reflections [Wil67, Bor57a, HH70, Bor65h, Hab19]. Reflexion [Bor41a]. reflexions [BLS42, Bor42g]. refractive [Bor18e, Bor63-50]. Regime [MP01]. region [Bor31c]. regions [dB53, dB53]. regular [BL18c, BL63c]. regulärer [BL18c, BL63c]. reinterpretation [Hei25]. Relation [Bor40g, Bor63-46, Bor43c]. Relations [Bow69, BH23b, Hei25]. relationship [BL11, Bor18e, Bor63-50]. relatividad [Bor22c]. Relativistic [Bor48d]. Relativitaätstheorie [Bor64b, Bor69c]. Relativität [Bor16a, Bor56d]. Relativitäts [Bor10c]. Relativitäts-prinzips [Bor10c]. Relativitätsprinzip [Bor13d, Bor09c]. Relativitätsprinzips [Bor09a, Bor09e, Bor09f, Bor09b, Bor10e, Bor63f, Bor63n]. Relativitätstheorie [Bor13d, Bor19b, Bor20d, Bor22a, Bor01b, BEP03, Din23, Rom65, Bor11, Inf55]. Relativity [Bor19b, BB24, Bor63-45, Wei12, Bor38d, Bor65a, Bor09a, Bor09e, Bor09f, Bor09c, Bor09b, Bor10c, Bor10e, Bor11, Bor13d, Bor16a, Bor20d, Bor22a, Bor22c, Bor38b, BF40a, BF40b, Bor50a, Bor50b, Bor56d, Bor62b, Bor63f, Bor63-42, Bor63n, Bor64b, Bor69c, Bor01b, BEP03, Inf55, RSS^{+}07]. Remarks [BI34c, Lan83, BL19, Bor38c, Bor53g, BF25a, Bor61a]. Reminiscences [Bor64d, Far79]. reply [Bor13d, Bor55c]. Reponse [Bor15a]. Representation [Bor63z, Khr09, Bor20g, Bor20h, Bor20i]. Republic [Sch07, Sch05]. Research [BN13, Bla67, Bor24c]. residual [BB33, BB63c]. resolution [Bor47g]. Response [BL19]. responsibility [Bor65k]. rest [Bor36d, BGCR49]. rest-moment [Bor36d]. Restless [A.53, Bor35f, Bor51i, Bor60b]. restraints [Bro68, Bro06]. Reststrahlen [BB33, BB63c]. rethinking [BKCK58]. retirement [ABdB^{+}53]. Reversal [Ily16]. Review [A.53, Alb71, Ano46, Ber50, Ber04, Bey06, Bla58, BDR84, Bor37a, Bor40b, Bor44a, Bor44b, Bor46b, Bor47d, Bor49i, Bor50b, Bor50d, Bor51d, Bor53b, Bor54a, Bre73, Bro72, Cas05, Cop61, Din50, Din23, Far79, For68, For70, Fow27, Gow79, Han80, Hec00, Hei72, Hei79a, Hei79b, Hol88, Kle70b, Kle70a, Kle70c, Kör59, Lax55, Lin98, Lin58, Mar50, Mar73, McC49, Mil60, Mil49, Mor54, MPK^{+}54, Nor71, Oes68, Pag63, Pry60, Rom65, Sch79, Seg64, Sho20, Str64, Str60, Süs57, Wol66, da 33, Bor50e, BvLM23, Som15, Som33, Som34]. Reviews [And81, Bor40c, Bor46c, Bor68a, CZPS37, HKP^{+}32a, HKP^{+}32b]. Revolution [Kae48, Bel99, Gre05a, Bey06, Cas05]. Rezension [Som33]. Rezensionen [Krö98]. Richard [MPK^{+}54]. rigid [Bor09a, Bor09b, Bor10c, Bor10e, Bor26a, Bor63f, Bor63n]. rise [Cha05, Hag08]. Robert [Ber04, Hei72]. Rock [Bor46i, BB63f, BB45, Bor47g, BB47b]. Rock-Salt [BB63f, Bor46i, Bor47g]. Romanian [Bor69f]. Röntgenstrahlen [Bor59c, Bor21a]. Roots [Kon78]. Rosenfeld [CS79a]. rotation [Bor15c]. rotatory [Bor18b]. Rotverschiebung [Bor53g]. Ruhm [BHvL21]. Rule [AG20, De 16, Ily16, Khr09, Man18, Man19, SF05, SGM^{+}12, Bla08, Dre17]. Russell [Bro72, Mar73, Ano06]. Russian [Bor28b]. russischen [Bor28b]. Rutherford [Dem03]. rutile [BB25b]. Rutils [BB25b].

s [Sal19]. S. [Bor47f]. Salt [BB63f, BB45, Bor46i, Bor47g, BB47b]. Satz [Bor14a, Bor63k]. Sauter [Som33, Som34]. Scalar [BF40c]. Scattering [BB43a, Bor42f, Bor63-46, GS47, Lap69, BB47a, Bor17, Bor18g, BG21b, BS41, Bor42b, Bor42c, BL42, Bor43c, BB43b, Hol07]. Schichten [BB21a]. Schmelzprozesses [Bor39c]. Scholars [Krö98, AS96]. school [Bor28c, Bor63-41]. Schriften [Bor63-57]. Schrödinger [BG61, Bor44a, Bor44b, Bor46c, Bor51d, Bor63w, Sig96]. Schrödinger/Wilhelm [BG61]. Schule [Bor28c, Bor63-41]. Schwefelatoms [BB20a, BB63d]. Schwingungen [BK12, BB21b, BB22a, BB63g, BB63h, BvK63b]. Science [Ano01, Bey06, Bla67, BB46, Bor53b, Bor58d, Bor65i, Cas05, Fre01, GGK02, Jon08, Kle70c, Pai00, RW94, Bor57d, Gre05a, BB71c, BB84, BB90]. Sciences [Str64]. Scientific [BDR84, BR95, Bor65i, Krö98, Lem83, Rod19, AS96, Bor63-57, Enz02b, Hag08, Haw11, ABdB^{+}53, Mor54, MPK^{+}54]. Scientist [BB63a, Sch49a, Sch49b]. Scientists [GR63, Krö98, MP01, AS96, BBF^{+}57b, Bor65k, Cou13, Str11, GR63]. Scotland [CS83]. Scribner [Kle70b]. Second [Hol07, Kle70c, Bor49h, Bor49e, HKP^{+}32a, HKP^{+}32b, Bor48b, Bor49e]. second-order [Bor49h]. secret [HP82, HP02]. sections [Koz91]. Seeking [Pal00]. Segen [Bor66g]. sein [Lin98]. seinem [Anoxx]. seiner [vS21]. seines [Bor59b, Bor63v]. Selected [CS79a, Bor63c, Bor63d]. Semester [Anoxx]. Semiclassical [Jam74a]. separate [Car89]. September [CS83]. Series [Lap69, BDR84]. Seven [Bor33a, Som34]. Seventieth [Ano52]. shape [BS19, BS63b]. Shea [MPK^{+}54]. Shell [BY50]. shift [Bor53g, Bor54c]. shook [Haw11]. Should [vW57]. Sieben [Bor33a, da 33, Som34]. Silent [Bor52a]. Simon [Bor58g, Bor63-39]. since [Meh75]. singular [dB53]. singulières [dB53]. Sinn [Bor28d, Bor29b]. Sir [Mor54, Bor40d, Bor40h, Bor41b, Bor58g, Bor63-39, Bor63-40]. Situation [Bor54b, Bor53c]. Sixty [FR13]. smallest [Fis10a, Fis12a]. Socialism [RW94]. Society [GGK02, Bor53c]. Soldner [BHvL21]. solid [Bor23a, Bor26c, BB27, Bor44d, BB63g]. Solids [SK66, Bor19d, Bor19e, BB21c, BB22a, BB63i]. Söllner [Krö98]. Solvay [BV09, Meh75]. Some [Bor36c, Bor37b, Bor38c, Bor42e, Bor31e]. Sommerfeld [Bor63b, Som05, Bor28c, Bor51a, Bor51b, Bor52b, Bor63-41, Som05]. Sommersemester [Anoxx]. Sonderdruck [Lin98]. Sorry [BR57]. Sources [vdW67, vdW07]. Space [Ano67, Bla67, Bor51j, Bor58f, Bor66b, Ure67, Wei12, Bor06, BK12, Bor14d, Bor63-61, Bor63-43, BvK63b, Bor66g, Bor13c, Bor51d]. Space-time [Bor51d]. Spaces [Cas12]. Spanish [Bor22c, Bor69d, BB71c, BB84]. Speaking [Krö98, AS96]. Special [Wei12, Bor63-42]. Specific [BB42a, BvK13b, Bor14c, BB21c, BB22a, BB42b, BB63g, BB63i, BvK63c]. specifische [BB21c, BB22a]. specifischen [BvK13b, Bor14c]. Spectrum [BFL46, GS47, BT34, Bor47g]. Speculation [Dys63]. Speculator [BB63a]. Spektrallinien [Bor40c]. spezifische [BB63g, BB63i]. spezifischen [BvK63c]. sphalerite [BB19, BB20b, BB20c, BB63e]. spherical [SG82]. Spiegelungen [HH70]. Springer [Kle70c, Lin98]. Springer-Verlag [Kle70c]. square [SG82]. Stabilität [Bor06, Bor63-43]. Stability [Bor63i, Bor63-31, Bor63-32, Bor06, Bor40e, BF40e, BM40, Bor42e, Bor44c, Mis40, Bor63-43]. Stahlung [da 33]. staircase [Fis10a, Fis12a]. standpoint [Bor10b]. Standpunkte [Bor10b]. stark [BL11]. starren [Bor09a, Bor09b, Bor10c, Bor10e, Bor63f, Bor63n]. state [Bor23a, Bor26c, Bor44d]. static [BF40e]. Statistical [BF38, Bor38f, BP44c, Bor44e, Bor46j, Bor49b, Bor55g, BH56, Bor63-44, Pai82, Bor37c, BB55, Bor55d, Bor61a, Bor63e, Bor63m, Bor44a, Bor44b, Bor46c, BH55]. Statistics [Bor63-37, Bor27c, Bor49d, BP44b]. Statistik [Bor27c, Bor63-37]. statistique [BB55]. statistische [Bor55d, Bor63m, BH55]. statistischen [Bor61a, Bor63e]. stellar [Bor53g]. stellare [Bor53g]. still [Kae48]. Stille [Bor52a]. Störungsrechnung [Bor23c, Bor23d]. Story [Jon08, MP01, HP82]. Stossanregung [BJN25]. Stoßvorgänge [Bor28e, Bor26e, Bor26d, Bor26f, Bor63-59, Bor63-36, Bor63-64]. Strahlung [Bor33a, Som34]. straight [vS21]. straight-line [vS21]. strength [BF40e]. Strong [Pei71]. strongly [BL11]. Structure [BY50, Bor51j, Bor20b, Bor26a, Bor26c, BB27, BB33, BB63c, Bro73b, HKP^{+}32a, HKP^{+}32b, BF31, Bor51d]. Structures [Bor46k, Bor46b]. Struktur [Bor26c, BF31]. Strukturbericht [HKP^{+}32a, HKP^{+}32b]. Studienjahre [Wei82]. Studies [Anoxx, BL19, Bor06]. Studium [Anoxx]. study [Bel90, Bro68, Bro06, Fis10a, Fis12a, Wei82]. stuff [Haw11]. Styles [BR95]. sub [Bro73b]. sub-structure [Bro73b]. subjectivity [Bor66f].

Subjektivität [Bor66f]. sublimation [BK23]. Sublimationswärme [BK23]. substances [Bor17]. Substanzen [Bor17]. Suggestion [Bor63-45, Bor38d]. sulfur [BB20a, BB63d]. sulla [HSBA60]. Summer [Anoxx]. sums [BB43c]. sun [HHW99]. superconductivity [BC48a, BS48, BC48c, BC48d, BC48b]. superiority [Bor57d]. supraconductivité [BC48a]. Supraleitfähigkeit [BS48]. surface [BS19, BS63b]. surmounting [Bor51e]. Survey [CZPS37]. sus [Bor22c]. Symbol [Bor66a, Bor64e, Bor64f, Bor65j, Bor65i]. Symmetry [Bor35c, Ily16, Bor36a]. System [Bor10e, Bor35c, BB63g, BB21b, BB22a, Bor36a, BB63h]. Systeme [BP22, BH55, BP63a]. Systems [BB21b, BB22a, BF38, BB63g, Bor63-44, BB63h, Bla08, BP22, BF33, Bor37c, BH55, BH56, BP63a, BF63a].

T [MPK^{+}54]. Tait [ABdB^{+}53, Bor36c]. Talks [BDR84]. tatsächlich [Bor55f]. teaching [Bor42h]. Technology [RW94]. Teilchen [Fis10a, Fis12a]. Teilchens [BL58, BL63d]. Temperaturabhängigkeit [Bor22f]. Temperature [Bor42f, BL42, BBK13, Bor22f]. Temperaturen [BB22a, BB21c, BB63i]. ‘Temperaturen”’ [BB63g]. temperatures [BB21c, BB22a, BB63g, BB63i]. Temperatursprung [BBK13]. tempo [Bor61b]. Ten [Jon08]. tensile [BF40e]. teoría [Bor22c]. Term [Gal01]. Testing [SGM^{+}12]. text [FL10]. textbook [Bor33c, Bor43b, Bor53e, Bor65d, Bor72]. Textbooks [BN13]. Their [HHW99, BS19, BS63b, Lem82]. Theorem [Bor63k, De 16, Bor14a, BF63b, BF28]. theoretic [Ali05]. Theoretical [Bor43c, Bor45a, Bor45d, Bor63-46, Bor14b, BB27, Bor39c, Bor40b, Bor42h, Bor49i, Hei25, Bor53g]. Theoretische [BB27, Bor53g]. theoretischen [Bor40c]. Theorie [Bor09a, Bor09b, Bor13e, BC13b, BC13a, BvK13b, Bor13a, Bor14c, Bor15a, Bor16a, BJN25, Bor28f, Bor29d, Bor29e, Bor32e, BS48, Bor51e, Bor63f, Bor63-62, Bor63n, BvK63c, Bor69e, Hab19, Bor38e, Bla58, Bor24c, BC48a, Bor49e, Bor52d]. Theorien [Bor28d, Bor29b]. Theories [Bor44f, Bor49b, Bor54a, Bor28d, Bor29b, KED^{+}60]. Theory [Ano56, BB42a, Bor11, BC13b, BC13a, Bor19b, Bor20d, Bor22a, Bor22c, BB24, Bor30a, BI33b, Bor34a, Bor34b, BS35, Bor35d, Bor35c, Bor37a, Bor38b, Bor42f, Bor45c, BG47c, Bor47e, BG47d, BC48c, BC48d, BG49, BCG49, Bor49g, BGK50, Bor52c, BH54, BW59, Bor62b, BI63a, BG63, Bor63-30, Bor63-29, Bor63-42, Bor63-45, BB63f, Bor63n, BW64b, Bor64b, BW65, Bor69e, BW69, Bor69c, BW75, BW^{+}80, BW97, BW99, Bor01b, BEP03, BWB19, Fre34, GS47, Hec00, KED^{+}60, Kör59, Lax55, RS51, SK66, Wol66, BV09, Ber05, Ber06, Bor09a, Bor09b, Bor10b, Bor13e, BvK13b, Bor13a, Bor14c, Bor14d, Bor15a, BL18c, Bor18b, BB19, Bor19c, Bor20b, BB20b, BB20c, BG21a, Bor23a, BH23c, Bor23c, Bor23d, Bor24d, BH24b, Bor24c, BB25a, BB25b, BJ25b, BF25b, BJN25, Bor26a, Bor26c, BO27, Bor28f, Bor29c, Bor29a, Bor29d, Bor29e, Bor30b, Bor31a, BM32, Bor32e, BGM33b]. theory [BGM33a, Bor33c, BI34a, BI34c, BI35, Bor36a, BN36a, BN36b, Bor36d, Bor36e, Bor38d, Bor38e, BLS42, Bor42g, BB42b, Bor42e, Bor43a, Bor43b, BG46, Bor47d, BG47a, BG47b, BB47b, BC48a, BS48, BC48b, BGCR49, Bor49e, Bor51e, Bor53e, Bor56b, Bor63g, BL63c, BM63, BB63e, BJ63b, Bor63-60, BF63d, BO63, Bor63-61, Bor63s, Bor63-62, BvK63c, Bor65a, Bor65d, Bor65e, BB68, Bor72, BH88, BO98, Dre17, Hab19, LBBH69, SG82, Str64, Str60, Bor11, Bor13d, Bor16a, Bor63f, Inf55]. Theory/Marley [KED^{+}60]. Thermal [BB43a, BB47a, BB43b, BS41, Bor42b]. thermochemical [Bor19c, Bor63s]. thermochemische [Bor19c, Bor63s]. Thermodynamics [Bor39e, Bor44e, Bor46j, Bor63z, Bor63-47, Bor63-48, BB63j, BO07, Bor20g, Bor20h, Bor20i, Bor21d, BB22b, Bor43d, BB43c, BG48, Bor63-63, Lan83, Bor48b, Bor49e, Bor46c, Bor44a, Bor44b]. Thermodynamik [Bor21d, BB22b, Bor63-63, BB63j, Bor20g, Bor20h, Bor20i, Bor48b, Bor63z]. thermodynamique [Bor49e]. thickness [BB21a]. Thin [BP63b, BB21a, BP42]. Thomson [Bor40d, Bor63-40, Bor09d, Bor40h, Bor41b]. Thomsonsche [Bor09d]. Thorndike [Bey06, Cas05]. Thoughts [Bor69a, Bor68c]. Three [SGM^{+}12, Bor20b, Sch05]. Thunderstorm [MPK^{+}54]. Time [Bor51j, Bor57e, Bor58e, Bor59d, Bor61b, Bor66e, Bor69f, Bor83, Ily16, Wei12, Enz02b, Lem82, PST07, Bor51d]. time-dependent [PST07]. Time-Reversal [Ily16]. times [BB05]. tint [Hab19]. Todestages [Bor59b, Bor63v]. Toeplitz [Bor40a, Bor40f, Bor63-33, Bor81]. together [Anoxx]. topics [Bor49i]. toy [GDR04]. Traditional [Bor63z, Bor20g, Bor20h, Bor20i]. traditionellen [Bor20g, Bor20h, Bor20i, Bor63z]. träge [Bor09c]. Tragedy [Jon08]. trajectoires [Bor59a]. trajectories [BH23a, Bor59a]. Trans [Bro72]. transfer [Bor46f]. Transformation [Bor61c, Che70]. Transformer [BFP45, BFP63, Ano46]. Transforms [Bow69]. Transistor [MPK^{+}54]. Translation [Bor62f]. Traubenberg [Bor46g, Bor46h, Bor63-34]. Travel [Bor66b, Bor66g]. Travelling [Rod19]. treasury [FF91]. Treatment [DM69, Bor22a, Bor36d, Bor36e, Bor39c]. Trouble [Cas05]. True [MP01]. Tubes [CZPS37]. Twentieth [Bro72]. Twenty [Bor38f]. Twenty-Ninth [Bor38f]. Twilight [RSS^{+}07]. Two [Bor49i]. tycoon [Hag08]. Type [De 16].

U.S.S.R. [Bor45d]. Übersichtsartikel [BDR84]. Überwindung [Bor51e]. Uhrenparadoxon [Bor58c, BB58, BB63b]. ultraroten [Bor18f, Bor32a, Bor63o, Bor63-53]. ultraviolet [Bor32d]. ultravioletten [Bor32d]. Umdenkens [BKCK58]. Umdeutung [Hei25]. unbounded [Kae48]. uncertain [BB05]. understanding [Bor28e]. undulatorische [Bor32c]. undulatory [Bor32c]. Unification [Bor44f]. Uniforme [CZPS37]. Unifying [Bor63-45, Bor38d]. Unitary [Bor36d, Bor36e]. Universe [Bor35f, Bor51i, Bor60b, A.53]. universities [Bor42h]. University [ABdB^{+}53, Bor36c, Krö98, Sho20]. unpublished [Fre01, FL10]. Unravelling [FR13]. Unsegen [Bor66g]. unseren [BL19, BB22a, BB63g]. unserer [Lor10]. unseres [Bor62a]. Untersuchungen [Bor06]. unto [Cas05]. use [Bor47c]. Using [BL63b, BL18b, Dre17, dB53].

V [Bor47b, Bor47f, Bor63-35, Bor46g, Bor59c]. Valence [Bor30a, Bor29a, Bor32e]. valency [Bor31a]. Valenz [Bor32e]. Validity [Hov18, Bor51e]. variation [BL42]. Variational [BO07]. Variationsprinzipe [BO07]. Vegard [BL19, BL19]. Verallgemeinerung [Bor09g]. Verantwortung [Bor65k]. Vergangenheit [HH70]. Verhältnis [BL11]. Verlag [Kle70c]. Verleger [Lin98]. Vernunft [Bor66c]. verschiedenen [Bor06]. version [GDR04]. Verständnis [Bor28e]. Versuch [Bor65i]. Verteilung [BvK13a, BvK63a]. Vertical [BP63b, BP42]. Verzeichnis [Bor63-57]. vi [Kle70b]. Vibration [BP63b]. Vibrations [BB22a, BP42, BB63g, BK12, BvK13a, BB21b, BS41, Bor42b, BB63h, BvK63b, BvK63a]. victory [Pal00]. View [Bre70]. Views [Bor68a, Bor68d, For68, Kle70b, Oes68]. viii [Kle70c]. visit [Bor58c]. Vocation [Kle70c]. Volume [Bor20l, BJ30, RSS^{+}07]. Volumen [Bor20l]. vom [Bor10b, Lor10, Bor19f, Bor66g]. vor [ABH^{+}55, ABH^{+}87]. Voraussagbarkeit [Bor59e]. vorbeigeht [vS21]. Vorgänge [Bor10b, BJ25b, BW26a, BW63b, BJ63b]. Vorhersagbarkeit [Bor58h]. Vorlesungen [BJ30, BH25, Giu13]. Vorstellungen [Lor10]. Vortrag [Lor10]. Vorträge [BDR84, Bor33a, Som34, da 33]. vue [Bor52d].

W [Ano46, Bor53b, Fow27, HKP^{+}32a, HKP^{+}32b, MPK^{+}54]. W. [Bor49i]. Wallace [Man18, Man19]. Walter [And81]. Wandel [Bor57e, Bor58e, Bor59d, Bor66e, Bor83, Lin58]. Wandlungen [KED^{+}60]. War [Pal00, Bor47c, Hab86, Sal19, Hol88]. Warfare [FW08, Hab22, Tho22, Bro68, Bro06, Cha05, Hab86, HP82, HP02, Hol88, Pal00]. Wärme [BB21c, BB22a, BB63g, BB63i, BvK13b, Bor14c, BvK63c]. Wärmestrahlung [Bor13a]. Wärmetheorie [BO07]. Wärmetönung [Hab19]. warn [ABH^{+}55, ABH^{+}87]. warnen [ABH^{+}55, ABH^{+}87]. Was [Bor31f, Bor63-58]. Wasserstoffmolekel [Bor22d]. Wave [Bor37d, Bor46j, Bor46k, Bor26f, BI34d, BF40d, BF40c, Bor46c, Bor63-64, Dre17, dB53, Bor46b]. waves [dB53]. weapons [ABH^{+}55, ABH^{+}87]. Weglänge [Bor20e, Bor63r]. Weird [Str11]. Weise [Bor65i]. welchem [vS21]. well [SG82]. Wellenmechanik [Bor26f, Bor63-64]. Weltkörpers [vS21]. Weltraumfahrt [Bor66g]. We’re [BR57, Nik08]. Werke [BDR84]. Werner [Bro72, Mar73, Sch05, Sch07, BDR84, Sch05, Sch07]. Wesen [Bor55e]. western [Pal00]. Weyl [Som05, Sig96, Som05]. Where [Car89]. Whitehouse [MPK^{+}54]. Whittaker [Bor54a]. Who [Bey06, Cas05, Cha05, Cou13, Gre05a]. Wide [Ano02, Bor02]. Wide-Ranging [Ano02, Bor02]. Wiederkehr [Bor59b, Bor63v]. Wilhelm [Bor52b, BG61, Bor63b, Bor62f]. Wirklichkeit [Bor65j, Bor65i, Bor64e, Bor64f]. wissenschaftliche [Lem83, BDR84]. wissenschaftlichen [Bor63-57]. wissenschaftlicher [Lin98]. within [Dem03]. Wolf [KED^{+}60, Sho20, Cop61, Hec00, Mil60, Str64, Str60, Wol66].

Wolfgang [Enz02b]. women [HHW99]. work [Bor09a, Bor10a, Bor64d, Sch05]. Works [BDR84, Bor63c, Bor63d, Ber15]. World [Bey06, Bor50d, Cas05, GR63, Hab86, Hol88, FF91, Gre05a, Gre06, GE08, Hag08, Haw11, Pal00]. Would [Nik08]. writings [Bor63-57]. Wrocław [Anoxx]. Wyckoff [HKP^{+}32a, HKP^{+}32b].

X [BB43a, BB47a, Bor21a, Bor41a, BS41, Bor42b, Bor42c, BLS42, Bor42g, BL42, Bor43c, BB43b, Bor59c, Bor63-46]. X-Ray [Bor63-46, Bor42c, BLS42, Bor42g, Bor43c]. X-Rays [BB43a, Bor41a, BB47a, Bor21a, BS41, Bor42b, BL42, BB43b, Bor59c]. xi [Bro72].

Year [FR13]. Years [Bor51h, Bor51f, Bor01a, Con62, Kae48, Wei82, Bor50d]. York [Kle70b, Kle70c, Mor54].

Zeeman [BS45]. Zeemann [BS63a]. Zeilinger [Fis10a, Fis12a]. Zeit [Lem82, Bor57e, Bor58e, Bor59d, Bor66e, Bor83, Lin58]. Zeitalters [Bor62a]. Zerstreuung [Bor17, Bor18g, BG21b]. Zientzia [BB90]. Zinkblende [BB20c, BB19, BB20b, BB63e]. zum [Ano63, Ber15, BF30, BF63c, Fis10a, Fis12a, Hei62, Bor13d, BB58, BB63b]. ’Zur [Bor10a, Bor15a, Bor18e, Bor19f, Bor20g, Bor20h, Bor20i, BB21a, BR31, Bor59b, Bor61a, Bor63g, Bor63e, Bor63v, Bor63z, Bor63-50, Hab19, KED^{+}60, Bor10d, Bor10e, Bor13e, BC13b, BC13a, BvK13b, Bor14d, BB19, BB20b, BB20c, Bor21d, BB22b, BH23c, BH24b, BB25a, BB25b, BJ25b, BJ25a, BJN25, BHJ26, Bor26e, Bor26f, BO27, Bor28f, Bor29c, Bor29d, Bor29e, Bor30b, Bor32d, BM32, Bor32e, BF33, BS48, Bor55h, BL58, BHJ62, BM63, BB63e, BJ63b, Bor63-60, BJ63a, BHJ63, BL63d, BO63, Bor63-59, BF63a, Bor63-61, Bor63-63, Bor63-62, BB63j, BvK63c, Bor63-64]. Zurek [SF05]. Zusammenhang [BK23]. Zustandes [Bor23a, Bor26c]. zweiatomiger [Bor18f, Bor63-53]. Zweiatomsystems [BF33, BF63a]. zweite [Bor48b]. zwischen [Bor18e, Bor20c, Bor63-50].

References

A:1953:BRB

[A.53] R. L. A. Book review: The Restless Universe, by Max Born. Philosophy of Science, 20(4):346, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

          jstor.org/stable/pdfplus/185055.pdf.

Appleton:1953:SPP

[ABdB^{+}53] Sir Edward Appleton, David Bohm, Louis de Broglie, Richard Courant, Albert Einstein, Pascual Jordan, Theodore V. Kármán, S. S. Penner, Alfred Landé, Erwin Schrödinger, and Hermann Weyl, editors. Scientific papers presented to Max Born on his retirement from the Tait Chair of Natural Philosophy in the University of Edinburgh. Oliver and Boyd, Edinburgh, UK; London, UK, 1953. vi + 94 pp. LCCN QC71 .S35.

Adler:1955:ANW

[ABH^{+}55] Kurt Adler, Max Born, W. Heisenberg, et al. Achtzehn Nobelpreisträger warnen vor Atomwaffen. (German) [Eighteen Nobel laureates warn against nuclear weapons]. Keesing’s Archiv der Gegenwart, 25(??):5251H–??, ???? 1955. The original typescript, entitled “Mainauer Kundgebung” [Mainau rally], signed by authors, is reproduced in [ABH^{+}87].

Adler:1987:ANW

[ABH^{+}87] Kurt Adler, Max Born, W. Heisenberg, et al. Achtzehn Nobelpreisträger warnen vor Atomwaffen. (German) [Eighteen Nobel laureates warn against nuclear weapons]. In ????, editor, 30 Jahre Göttinger Erklärung. Nachdenken über die Rolle des Wissenschaftlers in der Gesellschaft. (German) [30 years Göttingen declaration. Reflection on the role of the scientist in the company], page 42. ????, Munich, West Germany, 1987. Reproduction of original signed typescript of [ABH^{+}55].

Anandakrishnan:2011:GBA

[ADO11] Ramu Anandakrishnan, Mayank Daga, and Alexey V. Onufriev. An n log n generalized Born approximation. Journal of Chemical Theory and Computation, 7(3):544–559, 2011. CODEN JCTCCE. ISSN 1549-9618 (print), 1549-9626 (electronic).

Auffeves:2020:DBR

[AG20] Alexia Auffèves and Philippe Grangier. Deriving Born’s rule from an inference to the best explanation. Foundations of Physics, 50(12):1781–1793, December 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-020-00326-8.

Alberdi:1971:BRB

[Alb71] Inés Alberdi. Book review: Ciencia y conciencia en la era atómica, by Max Born and Hedwing Born. Revista española de la opinión pública, 25(??):564, September 1971. CODEN ???? ISSN 0034-

9429. URL http://www.jstor.org/stable/40199657;   http:
//www.jstor.org/stable/pdfplus/40199657.pdf.

Ali:2005:FTD

[Ali05] Abdulmuhsen H. Ali. A field theoretic derivation of the Max Born hypothesis. Journal of Mathematical Physics, 46(10):103302, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p103302_s1.

Anderson:1981:BRB

[And81] David L. Anderson. Book reviews: My Life: Recollections of a Nobel Laureate, by Max Born and Memoirs of a Physicist in the Atomic Age, by Walter M. Elsasser. American Journal of Physics, 49(1):94–95, January 1981. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v49/i1/p94_s1.

Anonymous:1946:BRB

[Ano46] Anonymous. Book review: A Photoelectric Transformer, by Max Born, R. Furth, and R. W. Pringle. Mathematical Tables and Other Aids to Computation, 2(14):89, April 1946. CODEN MTTCAS.

ISSN 0891-6837. URL http://www.jstor.org/stable/pdfplus/
2002537.pdf.

Anonymous:1952:SBA

[Ano52] Anonymous. Seventieth birthday anniversaries: Profs. J. Franck and M. Born. Nature, 170(4336):953, December 6, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v170/
n4336/pdf/170953b0.pdf.

Anonymous:1956:BBT

[Ano56] Anonymous. Books: Experiment and Theory in Physics, by Max Born. Bulletin of the Atomic Scientists, 12(7):280, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:MBG

[Ano63] Anonymous. Max Born zum 80. Geburtstag. (German) [Max Born’s 80th birthday]. Zeitschrift für Physik, 171(1):1, February 1963. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01379332.

Anonymous:1967:CEN

[Ano67] Anonymous. Comments: Editor’s note once again: Space. Bulletin of the Atomic Scientists, 23(3):30, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ure67, Bor68b].

Anonymous:1970:OPM

[Ano70] Anonymous. Obituary: Professor Max Born. Nature, 225(5233): 669–671, February 14, 1970. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v225/n5233/pdf/225669a0.pdf.

Anonymous:1982:MBC

[Ano82] Anonymous. Max Born’s centenary. Nature, 300(5892):468, December 9, 1982. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v300/n5892/pdf/300468a0.pdf.

Anonymous:2001:AS

[Ano01] Anonymous. Astrology and science. Physics in Perspective (PIP), 3(4):424, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2002:EWR

[Ano02] Anonymous. Errata: “The Wide-Ranging Family History of Max Born”. Notes and Records of the Royal Society of London, 56 (3):403, September 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.org/

stable/pdfplus/3557748.pdf. See [Bor02].

Anonymous:2006:REM

[Ano06] Anonymous. The Russell–Einstein Manifesto. The SPS Observer,

23(2):??, Summer 2006. URL http://www.spsobserver.org/

2006/observer_ethics.pdf. The Manifesto is signed by Max Born, Percy W. Bridgman, Albert Einstein, Leopold Infeld, Frederic Joliot-Curie, Herman J. Muller, Linus Pauling, Cecil F. Powell, Joseph Rotblat, Bertrand Russell, and Hideki Yukawa.

Anonymous:2010:QMB

[Ano10] Anonymous. Quantum mechanics born to be linear. Science, 329 (5990):367, July 23, 2010. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/329/5990/367.4.full.pdf.

Anonymous:20xx:MBS

[Anoxx]  Anonymous.  Max Born: Studium Sommersemester 1902

in Heidelberg gemeinsam mit seinem Freund Ernst Hellinger aus Breslau. (German) [Max Born: Summer semester studies 1902 in Heidelberg, together with his friend Ernst Hellinger from Wrocław]. Web site, 20xx. URL http://www.rzuser.

uni-heidelberg.de/~f25/homo-heid/rett-born.htm;   http:
//www.ub.uni-heidelberg.de/helios/fachinfo/www/math/homo-
heid/born.htm.

Ash:1996:FMS

[AS96] Mitchell G. Ash and Alfons Söllner, editors. Forced migration and scientific change: emigré German-speaking scientists and scholars after 1933. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 1996. ISBN 0-521-49741-8 (hardcover). xviii + 301 pp.

LCCN E184.G3 F73 1996. URL http://www.loc.gov/catdir/
description/cam027/95024894.html;   http://www.loc.gov/
catdir/samples/cam031/95024894.html;   http://www.loc.
gov/catdir/toc/cam026/95024894.html.

Baker:1983:BI

[Bak83] Lionel R. Baker. Born, the innovator. In Colles and Swift [CS83], pages 2–7. ISBN 0-89252-404-9 (paperback). LCCN QC350 .M39

1982. URL http://link.spie.org/PSISDG/0369/.

Born:1919:GZG

[BB19] Max Born and Elisabeth Bormann. Zur Gittertheorie der Zinkblende. (German) [Toward a lattice theory of sphalerite]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(21–22): 733–741, December 5, 1919. CODEN VDPEAZ. ISSN 0372-5448.

Born:1920:ESG

[BB20a] Max Born and Elisabeth Bormann. Die Elektronenaffinität des Schwefelatoms. (German) [The electron affinity of the sulfur atom]. Zeitschrift für Physik, 1(3):250–255, June 1920. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01329169.

Born:1920:GZG

[BB20b] Max Born and Elisabeth Bormann. Zur Gittertheorie der Zinkblende. (German) [Toward a lattice theory of sphalerite]. Annalen der Physik (1900), 367(11):218–246, ???? 1920. ISSN 1521-3889.

Born:1920:GZI

[BB20c] Max Born and Elisabeth Bormann. Zur Gittertheorie der Zinkblende. II. (German) [Toward a lattice theory of sphalerite. II]. Annalen der Physik (1900), 62(??):218–246, ???? 1920. CO-DEN ???? ISSN ????

Born:1921:IMB

[BB21a] Max Born and Elisabeth Bormann. Über eine interferometrische Methode zur Bestimmung der Dicke dünner Schichten. (German) [An interferometric method for determination of the thickness of thin layers]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 2(??):54–??, ???? 1921. CODEN VDPEAZ. ISSN 0372-5448.

Born:1921:SMS

[BB21b] Max Born and E. Brody. Über die Schwingungen eines mechanischen Systems mit endlicher Amplitude und ihre Quantelung. (German) [On the vibrations of a mechanical system with finite amplitude and its quantization]. Zeitschrift für Physik, 6(1):140–152, December 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01327973.

Born:1921:SWF

[BB21c] Max Born and E. Brody. Über die specifische Wärme fester Körper bei hohen Temperaturen. (German) [On the specific heat of solids at high temperatures]. Zeitschrift für Physik, 6(1):132–139, December 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01327972.

Born:1922:BUA

[BB22a] Max Born and E. Brody. Bemerkungen zu unseren Abhandlungen “Über die Schwingungen eines mechanischen Systems mit endlicher Amplitude und ihre Quantelung” und “Über die specifische Wärme fester Körper bei hohen Temperaturen”. (German) [Comments on our papers “On the vibrations of a mechanical system with finite amplitude and its quantization” and “The specific heat of solids at high temperatures”]. Zeitschrift für Physik, 8(1):205–207, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01329595.

Born:1922:TKI

[BB22b] Max Born and E. Brody. Zur Thermodynamik der Kristallgitter. II. (German) [Toward a thermodynamics of crystal lattice. II]. Zeitschrift für Physik, 11(1):327–352, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01328426.

Born:1924:ETR

[BB24] Max Born and Henry L. (Henry Leopold) Brose. Einstein’s Theory of Relativity. Methuen and Co. Ltd., London, UK, third edition, 1924. xi + 293 pp. LCCN QC6 .B66; 530 B73; QC6 .B645rE 1924.

Born:1925:GAG

[BB25a] Max Born and O. F. Bollnow. Zur Gittertheorie des Anatas. (German) [Toward a lattice theory of Anatas]. Nachr. Ges. Wiss. Göttingen, ??(??):18–21, ???? 1925. CODEN ???? ISSN ????

Born:1925:GRG

[BB25b] Max Born and O. F. Bollnow. Zur Gittertheorie des Rutils. (German) [Toward the lattice theory of rutile]. Naturwissenschaften, 13 (??):559, ???? 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1927:AFM

[BB27] Max Born and O. F. Bollnow. Der Aufbau der festen Materie: Theoretische Grundlagen. (German) [The structure of solid matter: theoretical foundations]. In Handbuch der Physik, volume 24, pages 370–465. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., ???? 1927.

Born:1933:FRG

[BB33] Max Born and M. Blackman. Über die Feinstruktur der Reststrahlen. (German) [On the fine structure of residual radiation]. Zeitschrift für Physik, 82(9–10):551–558, September 1933. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01338328.

Blackman:1942:RTS

[BB42a] M. Blackman and M. Born. Raman’s theory of specific heat of crystals. Nature, 150(3793):55, July 11, 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v150/n3793/pdf/150055a0.pdf.

Born:1942:RTS

[BB42b] Max Born and M. Blackman. Raman’s theory of specific heat of crystals. Nature, 150(??):55–??, ???? 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Begbie:1943:TSX

[BB43a] G. H. Begbie and M. Born. Thermal scattering of X-rays by crystals. Nature, 152(3844):19–20, July 3, 1943. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v152/n3844/pdf/152019a0.pdf.

Born:1943:TSX

[BB43b] Max Born and G. H. Begbie. Thermal scattering of X-rays by crystals. Nature, 152(??):19–??, ???? 1943. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1943:TCLb

[BB43c] Max Born and Mary Bradburn. The thermodynamics of crystal lattices. II. Calculation of certain lattice sums occurring in thermodynamics. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 39(2):104–113, July 1943. CO-DEN PCPSA4. ISSN 0008-1981.

Born:1945:RER

[BB45] Max Born and Mary Bradburn. The Raman effect in rock salt. Nature, 156(3967):567, November 10, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v156/n3967/pdf/156567a0.pdf.

Born:1946:SE

[BB46] Max Born and L. J. F. Brimble. Science in Egypt. Nature, 158 (4002):43–46, July 13, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v158/n4002/pdf/158043a0.pdf.

Begbie:1947:TSX

[BB47a] G. H. Begbie and Max Born. Thermal scattering of X-rays by crystals. I. Dynamical foundations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 188(1013):179–188, January 30, 1947. CODEN PRLAAZ. ISSN

0080-4630. URL http://rspa.royalsocietypublishing.org/
content/188/1013/179.

Born:1947:TRE

[BB47b] Max Born and Mary Bradburn. The theory of the Raman effect in crystals, in particular rock salt. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 188(1013):161–178, January 30, 1947. CODEN PRLAAZ. ISSN

0080-4630. URL http://rspa.royalsocietypublishing.org/
content/188/1013/161.

Born:1955:LSM

[BB55] Max Born and E. Bauer. L’interprétation statistique de la mécanique quantique — (conférence nobel, 1954). (French) [The statistical interpretation of quantum mechanics (Nobel lecture, 1954)]. Journal de Physique et le Radium, 16(10):737–743, October 1955. CODEN JPRAAJ. ISSN 0368-3842.

Born:1958:UGC

[BB58] Max Born and W. Biem. Zum Uhrenparadoxon. (German) [On the clock paradox]. Proc. Amst., B61(2):110–120, ???? 1958. CODEN ???? ISSN ????

Born:1963:BBS

[BB63a] Hedwig Born and Max Born. Books: The Scientist Speculator: An Anthology of Partly-Baked Ideas, edited by I. J. Good. Bulletin of the Atomic Scientists, 19(5):30–32, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This is a strongly critical review of the book. The review begins: “The publication of Books: The Scientist Speculator: An Anthology of Partly-Baked Idea is nothing less than a crime against the ethical code, unwritten but vital, of the community of scientists.” See rebuttal in [Dys63].

Born:1963:UG

[BB63b] Max Born and W. Biem. Zum Uhrenparadoxon. (German) [On the clock paradox]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 7, pages 220–230. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 20, 110–120 (1958).

Born:1963:FRG

[BB63c] Max Born and M. Blackman. Über die Feinstruktur der Reststrahlen. (German) [On the fine structure of residual radiation]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 31, pages 529–536. LCCN ???? Reprinted from Z. Physik 82, 551–558 (1933).

Born:1963:ESG

[BB63d] Max Born and Elisabeth Bormann. Die Elektronenaffinität des Schwefelatoms. (German) [The electron affinity of the sulfur atom]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 22, pages 404–409. LCCN ???? Reprinted from Z. Physik 1, 250–255 (1920).

Born:1963:GZG

[BB63e] Max Born and Elisabeth Bormann. Zur Gittertheorie der Zinkblende. (German) [Toward a lattice theory of sphalerite]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 21, pages 395–403. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 21, 733–741 (1919).

Born:1963:TRE

[BB63f] Max Born and Mary Bradburn. The theory of the raman-effect in crystals, in particular rock-salt. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 39, pages 636–653. LCCN ???? Reprinted from Proc. Roy. Soc. A 188, 161–178 (1947).

Born:1963:BUA

[BB63g] Max Born and E. Brody. Bemerkungen zu unseren Abhandlungen ,,Über die Schwingungen eines mechanischen Systems mit endlicher Amplitude und ihre Quantelung” und ,,Über die spezifische Wärme fester Körper bei hohen ‘Temperaturen”’. (German) [Comments on our papers, “On the Vibrations of a Finite Mechanical Mechanical System and its Quantization”, and “On the specific heat of solid bodies at high temperatures”]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 25, pages 431– 433. LCCN ???? Reprinted from Z. Physik 8, 205–207 (1922).

Born:1963:SMS

[BB63h] Max Born and E. Brody. Über die Schwingungen eines mechanischen Systems mit endlicher Amplitude und ihre Quantelung. (German) [On the vibrations of a mechanical system with finite amplitude and its quantization]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 24, pages 418–430. LCCN ???? See [BB63g]. Reprinted from Z. Physik 6, 140–152 (1921).

Born:1963:SWF

[BB63i] Max Born and E. Brody. Über die spezifische Wärme fester Körper bei hohen Temperaturen. (German) [On the specific heat of solids at high temperatures]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 23, pages 410–417. LCCN ???? See [BB63g]. Reprinted from Z. Physik 6, 132–139 (1921).

Born:1963:TKI

[BB63j] Max Born and E. Brody. Zur Thermodynamik der Kristallgitter II. (German) [Toward a thermodynamics of crystal lattices II]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 28, pages 483–508. LCCN ???? Reprinted from Z. Physik 11, 327–352 (1922).

Born:1968:DQT

[BB68] Max Born and Walter Biem. Dualism in quantum theory. Physics Today, 21(8):51–55, August 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/link.aip.org/link/?PTO/21/51/1;   http://physicstoday.
org/resource/1/phtoad/v21/i8/p51_s1.

Born:1969:AEH

[BB69a] Max Born and Hedwig Born, editors. Albert Einstein — Hedwig und Max Born. Briefwechsel 1916–55. (German) [Albert Einstein — Hedwig and Max Born. Correspondence 1916–55]. Nymphenburger Verlagshandlung, Munich, Germany, 1969. 329 pp. Preface by Bertrand Russell and foreword by Werner Heisenberg.

Born:1969:LGG

[BB69b] Max Born and Hedwig Born. Der Luxus des Gewissens. (German) [The Luxury of Conscience]. Nymphenburger Verlagshandlung GmbH, München, West Germany, 1969. ???? pp. LCCN ????

Born:1971:BELa

[BB71a] Max Born and Hedwig Born, editors. The Born–Einstein letters: correspondence between Albert Einstein and Max and Hedwig Born from 1916–1955. MacMillan Publishing Company, New York, NY, USA, 1971. ISBN 0-333-11267-9. xi + 240 + 5 pp. LCCN QC16.E5 A4513 1971b. Preface by Bertrand Russell and foreword by Werner Heisenberg. Translated by Irene Born.

Born:1971:BELb

[BB71b] Max Born and Hedwig Born, editors. The Born–Einstein letters; correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1971. ISBN 0-8027-0326-7. x + 240 pp. LCCN QC16.E5 A4513.

Born:1971:CCE

[BB71c] Max Born and Hedwig Born. Ciencia y conciencia en la era atómica. (Spanish) [Science and conscience in the atomic age]. Alianza Editorial, Madrid, Spain, 1971. ISBN 84-206-1313-4. 206 pp. LCCN ???? Translation of Der Luxus des Gewissens: Erlebnisse und Einsichten im Atomzeitalter.

Born:1984:CCE

[BB84] Max Born and Hedwig Born. Ciencia y conciencia en la era atómica. (Spanish) [Science and conscience in the atomic age]. Alianza Editorial, Madrid, Spain, 1984. 206 pp. LCCN ???? Translation of Der Luxus des Gewissens: Erlebnisse und Einsichten im Atomzeitalter.

Born:1990:ZEK

[BB90] Max Born and Hedwig Born. Zientzia eta kontzientzia aro atomikoan. (Basque) [Science and conscience in the atomic age]. Elhuyar, Donostia, Spain, 1990. 222 pp. LCCN ???? Translation of Der Luxus des Gewissens: Erlebnisse und Einsichten im Atomzeitalter.

Bialynicki-Birula:1999:BIN

[BB99] Iwo Białynicki-Birula. Born–Infeld nonlinear electrodynamics. Acta Physica Polonica B, 30(10):2875–2878, October 1999. CO-DEN APOBBB. ISSN 0587-4254 (print), 1509-5770 (electronic).

URL http://www.actaphys.uj.edu.pl/fulltext?series=Reg&
vol=30&page=2875.

Born:2005:BEL

[BB05] Max Born and Hedwig Born, editors. The Born–Einstein letters: friendship, politics, and physics in uncertain times: correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. MacMillan Publishing Company, New York, NY, USA, 2005. ISBN 1-4039-4496-2. xxxii + 235 pp. LCCN QC16.E5 A4 2005. URL http://

www.loc.gov/catdir/bios/hol059/2004061027.html;   http:
//www.loc.gov/catdir/description/hol054/2004061027.html;
 http://www.loc.gov/catdir/enhancements/fy0618/2004061027-

t.html. Note on the new edition by Gustav Born. New preface by Diana Buchwald and Kip S. Thorne. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. Translated by Irene Born.

Born:1950:MGB

[BBB50] Max Born, W. Brandt, and G. V. R. Born. In memoriam Gustav Born. experimental embryologist. Acta Anatomica, Basle, 10(4): 466–475, ???? 1950. CODEN ACATA5. ISSN 0001-5180 (print), 1421-9654 (electronic).

Blackett:1946:MAE

[BBD^{+}46] P. M. S. Blackett, M. Born, P. I. Dee, P. A. M. Dirac, N. Feather, E. A. Guggenheim, H. S. W. Massey, P. B. Moon, N. F. Mott, M. L. E. Oliphant, F. A. Paneth, R. E. Peierls, M. H. L. Pryce, F. E. Simon, Sir George Thompson, O. R. Frisch, and H. W. B. Skinner. Memo to the UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(12):6–8, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bopp:1957:DGN

[BBF^{+}57a] Fritz Bopp, Max Born, Rudolf Fleischmann, Walter Gerlach, Otto Hahn, Otto Haxel, Werner Heisenberg, Hans Kopfermann, Max von Laue, Heinz Maier-Leibnitz, Josef Mattauch, Friedrich-Adolf Paneth, Wolfgang Paul, Wolfgang Riezler, Fritz Strassmann, Wilhelm Walcher, Carl Friedrich von Weizsäcker, and Karl Wirtz. Declaration of the German nuclear physicists: April 13, 1957. Bulletin of the Atomic Scientists, 13(6):228, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The declaration calls for the Federal Republic of Germany (‘West Germany’) to “renounce explicitly and voluntarily the possession of atomic weapons of any kind”, and states that none of the authors is ‘ready in any way to take part in the production, the tests, or the application of atomic weapons”. See followup [vW57].

Bopp:1957:EAA

[BBF^{+}57b] Fritz Bopp, Max Born, Rudolf Fleischmann, Walter Gerlach, Otto Hahn, Otto Haxel, Werner Heisenberg, Hans Kopfermann, Max von Laue, Heinz Maier-Leibnitz, Josef Mattauch, Friedrich-Adolf Paneth, Wolfgang Paul, Wolfgang Riezler, Fritz Strassmann, Wilhelm Walcher, Carl Friedrich von Weizsäcker, and Karl Wirtz. Erklärung von Achtzehn Atomforschern. (German) [Declaration of eighteen atomic scientists]. Mitteilungen aus der Max-Planck-Gesellschaft zur Förderung der Wissenschaften, 2(??):62–64, ???? 1957.

Born:1913:MTG

[BBK13] Max Born, H. Bolza, and Th. V. Kármán. Molekularströmung und Temperatursprung. (German) [Molecular flow and temperature jump]. Nachr. Ges. Wiss. Göttingen, ??(??):221–235, ???? 1913. CODEN ???? ISSN ????

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover Publications, Inc., New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776

.B5713 1989. US$11.95. URL http://www.loc.gov/catdir/
description/dover031/89012033.html.

Born:1913:TEGb

[BC13a] Max Born and R. Courant. Zur Theorie des Eotvösschen Gesetzes. (German) [Toward the theory of the Eotvös Law]. Physikalische Zeitschrift, 14(16):731–740, August 15, 1913. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=mdp.39015021268936%3Bseq=823%3Bview=1up.

Born:1913:TEGa

[BC13b] Max Born and Richard Courant. Zur Theorie des Eotvösschen Gesetzes. (German) [Toward the theory of the Eotvös Law]. Naturwissenschaften, 1(28):674–675, July 11, 1913. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

link.springer.com/article/10.1007/BF01493978.

Born:1948:TSF

[BC48a] Max Born and K. C. Cheng. Sur la théorie de la supraconductivité. (French) [On the theory of superconductivity]. Journal de Physique et le Radium, 9(10):249–252, ???? 1948. CODEN JPRAAJ. ISSN 0368-3842.

Born:1948:TSc

[BC48b] Max Born and K. C. Cheng. The theory of superconductivity. Dokladi U.S.S.R., 62(??):313–318, ???? 1948. CODEN ???? ISSN ???? In Russian.

Born:1948:TSa

[BC48c] Max Born and Kai Chia Cheng. Theory of superconductivity. I. Nature, 161(4103):968–969, June 19, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v161/n4103/pdf/161968a0.pdf.

Born:1948:TSb

[BC48d] Max Born and Kai Chia Cheng. Theory of superconductivity. II. Nature, 161(4104):1017–1018, June 26, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v161/n4104/pdf/1611017a0.pdf.

Bowen:1980:BDQ

[BC80] Marshall Bowen and Joseph Coster. Born’s discovery of the quantum-mechanical matrix calculus. American Journal of Physics, 48(6):491–??, June 1980. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v48/i6/p491_s1.

Born:1949:RTEa

[BCG49] Max Born, K. C. Cheng, and Herbert S. Green. Reciprocity theory of electrodynamics. Nature, 164(4163):281–282, August 13, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v164/
n4163/pdf/164281b0.pdf.

Blum:1984:WHGf

[BDR84] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke: Collected Works: Series B: Scientific Review Papers, Talks, and Books: Wissenschaftliche Übersichtsartikel, Vorträge und Bücher, volume 1 of Gesammelte Werke / Collected Works. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). xx + 936 pp. LCCN

???? URL http://www.springerlink.com/content/978-3-642-
61742-3.

Beller:1990:BPI

[Bel90] Mara Beller. Born’s probabilistic interpretation: a case study of ‘concepts in flux’. Studies in History and Philosophy of Science Part A, 21(4):563–588, December 1990. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0039368190900335.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp.

LCCN QC174.13 .B45 1999. URL http://www.loc.gov/catdir/
description/uchi052/99035499.htm;   http://www.loc.gov/
catdir/toc/uchi052/99035499.htm.

Born:2003:RE

[BEP03] Max Born, Jürgen Ehlers, and Markus Pössel. Die Relativitätstheorie Einsteins. (German) [Einstein’s Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., seventh edition, 2003. ISBN 3-540-00470-X. xiv

+ 501 pp. LCCN ???? URL http://www.gbv.de/dms/ilmenau/
toc/367574594.PDF.

Bergwitz:1915:AGP

[Ber15] Karl Bergwitz, editor. Arbeiten aus den Gebieten der Physik, Mathematik, Chemie: Festschrift Julius Elster und Hans Geitel zum 60. Geburtstag. (German) [Works from the fields of physics, mathematics, chemistry: Festschrift for Julius Elster and Hans Geitel’s 60th birthday]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1915. xi + 719 pp. LCCN QD455 .F47 1915. URL http://

catalog.hathitrust.org/api/volumes/oclc/47733990.html.

Bergmann:1950:BRB

[Ber50] Gustav Bergmann. Book review: Natural Philosophy of Cause and Chance, by Max Born. Philosophy of Science, 17(2):196–199, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/184920.
pdf.

Bernstein:2004:EJR

[Ber04] Barton J. Bernstein. The enigma of J Robert Oppenheimer. [review of Oppenheimer: Portrait of an Enigma, Jeremy Bernstein, 2004 Gerald Duckworth/Ivan R Dee 240pp, £14.99/$25.00 hardcover, J Robert Oppenheimer: The American Century, David C Cassidy 2004 Pi Press/Prentice Hall 480pp]. Physics World, 12(12):36–37, December 2004. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://iopscience.iop.org/pwa/
full/pwa-pdf/17/12/phwv17i12a32.pdf.

Bernstein:2005:MBQ

[Ber05] Jeremy Bernstein. Max Born and the quantum theory. American Journal of Physics, 73(11):999–1008, November 2005. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v73/

i11/p999_s1. See erratum [Ber06].

Bernstein:2006:EMB

[Ber06] Jeremy Bernstein. Erratum: “Max Born and the quantum theory” [Am. J. Phys. 73 (11), 999–1008 (2005)]. American Journal of Physics, 74(2):160, February 2006. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v74/i2/p160_s1. See [Ber05].

Born:1928:PMG

[BEvL^{+}28] M. Born, A. Einstein, M. v. Laue, E. Schrödinger, and A. Sommerfeld. Planck-Medaille. (German) [Planck Medal]. Naturwissenschaften, 16(20):368, May 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/gg356437362203j2/fulltext.pdf.

Beyler:2006:BRN

[Bey06] Richard H. Beyler. Book review: Nancy Thorndike Greenspan: The End of the Certain World: The Life and Science of Max Born: The Nobel Physicist Who Ignited the Quantum Revolution. Isis, 97 (3):569–570, September 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/509982;   http://www.jstor.org/stable/pdfplus/
10.1086/509982.pdf.

Born:1925:BDR

[BF25a] Max Born and J. Franck. Bemerkungen über die Dissipation der Reaktionswärme. (German) [Remarks on the dissipation of reaction heat]. Annalen der Physik (1900), 76(??):225–230, ???? 1925. CODEN ???? ISSN ????

Born:1925:QMG

[BF25b] Max Born and James Franck. Quantentheorie und Molekelbildung. (German) [Quantum theory and molecular formation]. Zeitschrift für Physik, 31(1):411–429, February/April 1925. CO-DEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02980594.

Born:1928:BAG

[BF28] Max Born and V. Fock. Beweis des Adiabatensatzes. (German) [Proof of the Adiabatic Theorem]. Zeitschrift für Physik, 51(3– 4):165–180, March 1928. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01343193.

Born:1930:BPA

[BF30] Max Born and J. Franck. Beitrag zum Problem der Adsorptionskatalyse. (German) [Contribution to the problem of adsorption catalysis]. Nachr. Ges. Wiss. Göttingen, ??(??):77–89, ???? 1930. CODEN ???? ISSN ????

Born:1925:SME

[BF31] M. Born and J. Franck, editors. Struktur der Materie in Einzeldarstellungen. (German) [Structure of Matter in Monographs]. Julius Springer, Berlin, Germany, 1925–1931. ???? pp. LCCN ???? Thirteen volumes issued over seven years.

Born:1933:QZG

[BF33] Max Born and S. Flugge. Zur Quantenmechanik des Zweiatomsystems. (German) [Toward a quantum mechanics of diatomic systems]. Annalen der Physik (1900), 16(??):768–780, ???? 1933. CODEN ???? ISSN ????

Born:1938:SMC

[BF38] Max Born and Klaus Fuchs. The statistical mechanics of condensing systems. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 166(926):391–414, June 3, 1938. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/166/926/391;   http:
//www.jstor.org/stable/pdfplus/97007.pdf.

Born:1939:FERa

[BF39a] Max Born and Klaus Fuchs. On fluctuations in electromagnetic radiation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 170(941):252–265, March 21, 1939. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/170/941/252.

Born:1939:FERb

[BF39b] Max Born and Klaus Fuchs. On fluctuations in electromagnetic radiation. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 172(951):465–466, September 4, 1939. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/172/951/465.

Born:1940:MCRa

[BF40a] Max Born and Klaus Fuchs. The mass centre in Relativity. Nature, 145(3676):587, April 13, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v145/n3676/pdf/145587a0.pdf.

Born:1940:MCRb

[BF40b] Max Born and Klaus Fuchs. Mass centre in Relativity. Nature, 145(3685):933, June 15, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v145/n3685/pdf/145933b0.pdf.

Born:1940:RIS

[BF40c] Max Born and Klaus Fuchs. Reciprocity II. Scalar wave functions. Proceedings of the Royal Society of Edinburgh, 60(1):100–116, ???? 1940. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Born:1940:RIR

[BF40d] Max Born and Klaus Fuchs. Reciprocity III. Reciprocal wave functions. Proceedings of the Royal Society of Edinburgh, 60(2):141–146, ???? 1940. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Born:1940:SCLc

[BF40e] Max Born and Reinhold Fürth. The stability of crystal lattices. III. An attempt to calculate the tensile strength of a cubic lattice by purely static considerations. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36(4): 454–465, October 1940. CODEN PCPSA4. ISSN 0008-1981.

Born:1963:QZG

[BF63a] Max Born and S. Flügge. Zur Quantenmechanik des Zweiatomsystems. (German) [Toward a quantum mechanics of diatomic systems]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 67, pages 417–429. LCCN ???? Reprinted from Ann. Physik 16, 768–780 (1933).

Born:1963:BAG

[BF63b] Max Born and V. Fock. Beweis des Adiabatensatzes. (German) [Proof of the adiabatic theorem]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 62, pages 338–353. LCCN ???? Reprinted from Z. Physik 51, 165–180 (1928).

Born:1963:BPA

[BF63c] Max Born and James Franck. Beitrag zum Problem der Adsorptionskatalyse. (German) [Contribution to the problem of adsorption catalysis]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 65, pages 382–395. LCCN ???? Reprinted from Nachr. Ges. Wiss. Göttingen 1930, 77–89.

Born:1963:QMGa

[BF63d] Max Born and James Franck. Quantentheorie und Molekelbildung. (German) [Quantum theory and molecular formation]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 51, pages 78–96. LCCN ???? Reprinted from Z. Physik 81, 411–429 (1925).

Born:1946:LDB

[BFL46] M. Born, R. Fürth, and R. Ladenburg. Long duration of the Balmer spectrum in hydrogen. Nature, 157(3980):159, February 9, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v157/
n3980/pdf/157159a0.pdf.

Born:1945:PEF

[BFP45] Max Born, Reinhold Fürth, and R. W. Pringle. A photo-electric Fourier transformer. Nature, 156(3973):756–757, December 22, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v156/
n3973/pdf/156756a0.pdf.

Born:1963:PEF

[BFP63] Max Born, R. Fürth, and R. W. Pringle. A photo electric Fouriertransformer. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 5, pages 192–195. LCCN ???? Reprinted from Nature 156, 756 (1945).

Born:1921:EGGb

[BG21a] Max Born and Walter Gerlach. Elektronenaffinität und Gittertheorie. (German) [Electron affinity and lattice theory]. Zeitschrift für Physik, 5(5–6):433–441, September 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01327680.

Born:1921:ZLG

[BG21b] Max Born and Walter Gerlach. Über die Zerstreuung des Lichtes in Gasen. (German) [On the scattering of light in gases]. Zeitschrift für Physik, 5(5–6):374–375, September 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01327671.

Born:1946:GKT

[BG46] Max Born and Herbert S. Green. A general kinetic theory of liquids. I. The molecular distribution functions. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 188(1012):10–18, December 31, 1946. CODEN PRLAAZ.

ISSN 0080-4630. URL http://rspa.royalsocietypublishing.
org/content/188/1012/10.

Born:1947:GKTc

[BG47a] Max Born and Herbert S. Green. A general kinetic theory of liquids. III. Dynamical properties. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 190 (1023):455–474, September 9, 1947. CODEN PRLAAZ. ISSN

0080-4630. URL http://rspa.royalsocietypublishing.org/
content/190/1023/455.

Born:1947:GKTd

[BG47b] Max Born and Herbert S. Green. A general kinetic theory of liquids. IV. Quantum mechanics of fluids. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 191(1025):168–181, November 11, 1947. CODEN PRLAAZ. ISSN

0080-4630. URL http://rspa.royalsocietypublishing.org/
content/191/1025/168.

Born:1947:KTL

[BG47c] Max Born and Herbert S. Green. A kinetic theory of liquids. Nature, 159(4034):251–254, February 22, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v159/n4034/pdf/159251a0.pdf.

Born:1947:QTL

[BG47d] Max Born and Herbert Sydney Green. Quantum theory of liquids. Nature, 159(4048):738–739, May 31, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v159/n4048/pdf/159738a0.pdf.

Born:1948:KBT

[BG48] Max Born and Herbert S. Green. The kinetic basis of thermodynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 192(1029):166–180, February 4, 1948. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/192/1029/166.

Born:1949:GKT

[BG49] Max Born and Herbert S. Green. A General Kinetic Theory of Liquids. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1949. ???? pp. LCCN ????

Born:1961:ES

[BG61] Max Born and G. Glaser. Erwin Schrödinger/Wilhelm Keil 70 Jahre. Physikalische Blätter, 17(2):85–87, February 1961. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19610170205/abstract.

Born:1963:GKT

[BG63] Max Born and H. S. Green. A general kinetic theory of liquids. I. The molecular distribution functions. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 46, pages 710–?? LCCN ???? Reprinted from Proc. Roy. Soc. A 188, 10–18 (1946).

Born:1949:QTR

[BGCR49] Max Born, Herbert S. Green, K. C. Cheng, and A. E. Rodriguez. Quantum theory of rest masses. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and physical sciences, 62(4): 470–488, ???? 1949. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic). With appendices by K. C. Cheng and A. E. Rodriguez.

Born:1990:QMI

[BGĬ90] M. Born, V. Geĭzenberg, and P. Ĭordan. Quantum mechanics. II (a fragment). In Einstein collection, 1986–1990 (Russian), pages 146– 157. “Nauka”, Moscow, Russia, 1990. Translated from the German by A. N. Temchin.

Born:1950:GKT

[BGK50] Max Born, Herbert S. Green, and John G. Kirkwood. A general kinetic theory of liquids. Physics Today, 3(10):35–37, October 1950. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
3/35/2;   http://physicstoday.org/resource/1/phtoad/v3/
i10/p35_s2.

Born:1933:DGKb

[BGM33a] Max Born and M. Göppert-Mayer. Dynamische Gittertheorie der Kristalle. (German) [Dynamical lattice theory of crystals]. In Handbuch der Physik, volume 24/2, pages 623–794. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1933.

Born:1933:DGKa

[BGM33b] Max Born and Maria Göppert-Mayer. Dynamische Gittertheorie Kristalle. (German) [Dynamical crystal lattice theory]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933. ???? pp. LCCN ????

Born:1923:EIA

[BH23a] Max Born and Werner Heisenberg. Die Elektronenbahnen im angeregten Heliumatom. (German) [The electron trajectories in the excited helium atom]. Zeitschrift für Physik, 16(1):229–243, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01327396;   http://www.springerlink.com/content/
t75836n61671070k/.

Born:1923:PBM

[BH23b] Max Born and Werner Heisenberg. Über Phasenbeziehungen bei den Bohrschen Modellen von Atomen und Molekeln. (German) [On phase relations in the Bohr model of atoms and molecules]. Zeitschrift für Physik, 14(1):44–55, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://www.springerlink.com/content/v71005762t157n89/.

Born:1923:QMM

[BH23c] Max Born and Erich Hückel. Zur Quantentheorie mehratomiger Molekeln. (German) [Toward a quantum theory of polyatomic molecules]. Physikalische Zeitschrift, 24(1):1–12, January 1, 1923. CODEN PHZTAO. ISSN 0369-982X. URL https://

jscholarship.library.jhu.edu/handle/1774.2/214.

Born:1924:EDI

[BH24a] Max Born and Werner Heisenberg. Über den Einfluß der Deformierbarkeit der Ionen auf optische und chemische Konstanten. I. (German) [On the influence of the deformability of ions on optical and chemical constants. I]. Zeitschrift für Physik, 23(1):388–410, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://www.springerlink.com/content/
h0588704m33v5l60/.

Born:1924:QMG

[BH24b] Max Born and Werner Heisenberg. Zur Quantentheorie der Molekeln. (German) [Toward a quantum theory of molecules]. Annalen der Physik (1900), 379(9):1–31, ???? 1924. CODEN ???? ISSN 1521-3889.

Born:1925:VAG

[BH25] Max Born and Friedrich Hund. Vorlesungen über Atommechanik. (German) [Lectures on atom mechanics], volume 2 of Struktur der Materie in Einzeldarstellungen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1925. x + 358 pp. LCCN ???? English translation in [BH27, Bor60c].

Born:1927:MA

[BH27] Max Born and Friedrich Hund. Mechanics of the Atom. Bell, London, UK, 1927. xvi + 317 pp. LCCN ???? Reprinted in [Bor60c].

Born:1928:MQF

[BH28] Max Born and Werner Heisenberg. La mécanique des quanta. (French) [The mechanics of quanta]. In ????, editor, Electrons et Photons: Rapports et Discussions du Cinquième Conseil de Physique Tenu à Bruxelles du 24 au 29 Octobre 1927 sous les Auspices de l’Institut International de Physique Solvay, pages 143–184. Gauthier-Villars et cie, Paris, France, 1928. LCCN ????

Born:1954:DTC

[BH54] Max Born and Kun Huang. Dynamical Theory of Crystal Lattices. International Series of Monographs on Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1954. xii + 420 pp. LCCN ????

Born:1955:SDM

[BH55] Max Born and D. J. Hooton. Statistische Dynamik mehrfach periodischer Systeme. (German) [Statistical dynamics of multiperiodic systems]. Zeitschrift für Physik, 142(2):201–218, April 1955. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01329422.

Born:1956:SDM

[BH56] Max Born and D. J. Hooton. Statistical dynamics of multiplyperiodic systems. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 52(2):287–300, April 1956. CODEN PCPSA4. ISSN 0008-1981.

Born:1963:EIA

[BH63a] Max Born and Werner Heisenberg. Die Elektronenbahnen im angeregten Heliumatom. (German) [Electron orbits in the excited helium atom]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 48, pages 23–37. LCCN ???? Reprinted from Z. Physik 16, 229–243 (1923).

Born:1963:EDI

[BH63b] Max Born and Werner Heisenberg. Über den Einfluß der Deformierbarkeit der Ionen auf optische und chemische Konstanten I. (German) [On the influence of the deformability of ions on optical and chemical constants I]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 49, pages 38–60. LCCN ???? Reprinted from Z. Physik 23, 388–410 (1924).

Born:1975:QM

[BH75] Max Born and Werner Heisenberg. 5. quantum mechanics. In Mehra [Meh75], chapter 6, pages 146–149. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Born:1984:MQ

[BH84] Max Born and Werner Heisenberg. La mécanique des quanta. (French) [The mechanics of quanta ]. In Blum et al. [BDR84], pages 58–99. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_7/.

Born:1988:DTC

[BH88] Max Born and Kun Huang. Dynamical theory of crystal lattices. Oxford Classic Texts in the Physical Sciences. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1988. xii + 420 pp. UK£20.00.

Born:1926:QIG

[BHJ26] Max Born, Werner Heisenberg, and Pascual Jordan. Zur Quantenmechanik II. (German) [Toward quantum mechanics. II]. Zeitschrift für Physik, 35(8–9):557–615, August 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01379806;   http://
www.springerlink.com/content/r46w151t550pl624/. English

translation in [BHJ67].

Born:1962:BMG

[BHJ62] Max Born, Werner Heisenberg, and Pascual Jordan, editors. Zur Begründung der Matrizenmechanik. (German) [Toward the Foundations of Matrix Mechanics], volume 2 of Dokumente der Naturwissenschaft. Abteilung Physik. E. Battenberg, Stuttgart, Germany, 1962. 135 pp. LCCN QC174.3 .Z8.

Born:1963:QIG

[BHJ63] Max Born, Werner Heisenberg, and Pascal Jordan. Zur Quantenmechanik II. (German) [Toward quantum mechanics II]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 54, pages 155–213. LCCN ???? Reprinted from Z. Physik 85, 557–615 (1926).

Born:1967:QMc

[BHJ67] Max Born, Werner Heisenberg, and Pascual Jordan. On quantum mechanics II. In van der Waerden [vdW67], pages 321–385. LCCN QC174.12 S655. English translation of [BHJ26].

Born:1921:RSE

[BHvL21] Max Born, David Hilbert, and Max von Laue. Ruhm, Soldner, und Einstein. (German) [Fame, Soldner, and Einstein]. Frankfurter Zeitung, November 6, 1921. Critical comments on [vS21].

Born:1933:LEM

[BI33a] Max Born and L. Infeld. Letter: Electromagnetic mass. Nature, 132(3347):970, December 23, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v132/n3347/pdf/132970a0.pdf.

Born:1933:LFN

[BI33b] Max Born and Leopold Infeld. Letter: Foundations of the new field theory. Nature, 132(3348):1004, December 30, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v132/
n3348/pdf/1321004b0.pdf.

Born:1934:FNF

[BI34a] Max Born and L. Infeld. Foundations of the new field theory. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 144(852):425–451, March 29, 1934. CODEN PRLAAZ. ISSN 0080-4630. URL http:

//rspa.royalsocietypublishing.org/content/144/852/425.
full.pdf+html.

Born:1934:QNF

[BI34b] Max Born and L. Infeld. On the quantisation of the new field equations. I. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 147(862):522–546, December 1, 1934. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/147/862/522.

Born:1934:RPF

[BI34c] Max Born and L. Infeld. Remarks on the paper by Frenkel on Born’s theory of the electron. Proceedings of the Royal Society of London.

Series A, Mathematical and physical sciences, 146(859):935, October 15, 1934. CODEN PRLAAZ. ISSN 0080-4630. URL http://

rspa.royalsocietypublishing.org/content/146/859/935. See

[Fre34].

Born:1934:DEO

[BI34d] Max Born and Leopold Infeld. Déduction de l’équation d’ondes de Dirac à partir de l’électrodynamique quantique. (French) [Derivation of the Dirac wave equation from quantum electrodynamics]. Comptes rendus de l’Académie des sciences, Paris, 199(??):1596– 1598, ???? 1934. CODEN ???? ISSN ????

Born:1934:PNE

[BI34e] Max Born and Leopold Infeld. Principes de la nouvelle électrodynamique quantique. (French) [Principles of the new quantume electrodynamics]. Comptes rendus de l’Académie des sciences, Paris, 199(??):1297–1299, ???? 1934. CODEN ???? ISSN ????

Born:1935:QNF

[BI35] Max Born and L. Infeld. On the quantization of the new field theory. II. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150(869):141–166, May 1, 1935. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/150/869/141.

Born:1963:FNF

[BI63a] Max Born and Leopold Infeld. Foundations of the new field theory. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 74, pages 514–540. LCCN ???? Reprinted from Proc. Roy. Soc. A 144, 425–451 (1934).

Born:1963:QNF

[BI63b] Max Born and Leopold Infeld. On the quantisation of the new field equations I. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 75, pages 541–559. LCCN ???? Reprinted from Proc. Roy. Soc. A 147, 522–546 (1934).

Born:1967:EEG

[BI67] Max Born and Leopold Infeld. Erinnerungen an Einstein. (German) [Memories of Einstein]. Union-Verlag, Berlin, West Germany, 1967. ?? pp. LCCN ????

Born:1969:EEG

[BI69] Max Born and Leopold Infeld. Erinnerungen an Einstein. (German) [Memories of Einstein]. Union-Verlag, Berlin, West Germany, 1969. 74 pp. LCCN ????

Born:1925:QGT

[BJ25a] Max Born and Pascual Jordan. Zur Quantenmechanik. (German) [Toward quantum mechanics]. Zeitschrift für Physik, 34(1): 858–888, December 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/ article/10.1007/BF01328531. English translation of pages 858– 882 in [BJ67].

Born:1925:QAV

[BJ25b] Max Born and Pascual Jordan. Zur Quantentheorie aperiodischer Vorgänge. (German) [Toward the quantum theory of aperiodic processes]. Zeitschrift für Physik, 33(1):479–505, December 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01328329.

Born:1930:EQI

[BJ30] Max Born and Pascual Jordan. Elementare Quantenmechanik. (II. Band der Vorlesungen über Atommechanik). (German) [Elementary quantum mechanics. II. Volume of lectures on atom mechanics], volume 9 of Struktur der Materie in Einzeldarstellungen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1930. xi + 434 pp. LCCN ????

Born:1963:QGQb

[BJ63a] Max Born and Pascal Jordan. Zur Quantenmechanik. (German) [Toward quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 53, pages 124–154. LCCN ???? Reprinted from Z. Physik 34, 858–888 (1925).

Born:1963:QAV

[BJ63b] Max Born and Pascal Jordan. Zur Quantentheorie aperiodischer Vorgänge. (German) [Toward the quantum theory of aperiodic processes]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 52, pages 97–123. LCCN ???? Reprinted from Z. Physik 83, 479–505 (1925).

Born:1967:QMb

[BJ67] Max Born and Pascual Jordan. On quantum mechanics. In van der Waerden [vdW67], pages 277–306. LCCN QC174.12 S655. English translation of pages 858–882 of [BJ25a].

Born:1977:QM

[BJ77] M. Born and P. Jordan. Quantum mechanics. Uspekhi Fizicheskikh Nauk, 122(4):586–611, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Born:1925:TSA

[BJN25] Max Born, P. Jordan, and L. Nordheim. Zur Theorie der Stossanregung von Atomen und Molekülen. (German) [Toward the theory of collision excitation of atoms and molecules]. Naturwissenschaften, 13(??):969–970, ???? 1925. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1912:SRG

[BK12] Max Born and Th. V. Kármán. Über Schwingungen in Raumgittern. (German) [On vibrations in space lattices]. Physikalische Zeitschrift, 13(8):297–309, April 15, 1912. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=mdp.39015023176806%3Bseq=361%3Bview=1up.

Born:1923:ZDS

[BK23] Max Born and H. Kornfeld. Über den Zusammenhang des Dipolmoments und der Sublimationswärme der Halogenwasserstoffe. (German) [On the link between the dipole moment and heat of sublimation of the hydrogen halides]. Physikalische Zeitschrift, 24(6):121– 124, March 15, 1923. CODEN PHZTAO. ISSN 0369-982X. URL

https://jscholarship.library.jhu.edu/handle/1774.2/214.

Born:1958:AAN

[BKCK58] Max Born, H. Kortum, T. Cumme, and Karl Kromphardt. Aussprache über Atomgefahren: Über die Notwendigkeit des Umdenkens. (German) [Debate on nuclear dangers: on the necessity of rethinking]. Physikalische Blätter, 14(1):26–35, January 1958. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19580140106/abstract.

Born:1911:VEA

[BL11] Max Born and Rudolf Ladenburg. Über das Verhältnis von Emissions- und Absorptionsvermögen bei stark absorbierenden Körpern. (German) [On the relationship between emission and absorbance in strongly absorbing bodies]. Physikalische Zeitschrift, 12(6):198–202, March 15, 1911. CODEN PHZTAO. ISSN

0369-982X. URL http://babel.hathitrust.org/cgi/pt?id=
mdp.39015011417709%3Bseq=234%3Bview=1up.

Born:1918:KBA

[BL18a] Max Born and A. Landé. Kristallgitter und Bohrsches Atommodell. (German) [Crystal lattice and Bohr’s atomic model]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(21–24):202–209, December 30, 1918. CODEN VDPEAZ. ISSN 0372-5448.

Born:1918:ABK

[BL18b] Max Born and A. Landé. Über die absolute Berechnung der Kristalleigenschaften mit Hilfe Bohrscher Atommodelle. (German) [On the calculation of absolute crystal properties using Bohr atomic models]. S. B. preuss. Akad. Wiss. Berlin, ??(??):1048–1068, ???? 1918. CODEN ???? ISSN ????

Born:1918:BKR

[BL18c] Max Born and A. Landé. Über die Berechnung der Kompressibilität regulärer Kristalle aus der Gittertheorie. (German) [On the calculation of the compressibility of regular crystals from lattice theory]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(21–24):210–216, December 30, 1918. CODEN VDPEAZ. ISSN 0372-5448.

Born:1919:ABH

[BL19] Max Born and A. Landé. Antwort auf die Bemerkungen des Herrn L. Vegard zu unseren Arbeiten über Kristallgitter und Bohrsches Atommodell. (German) [Response to the remarks of Mr. L. Vegard to our studies on the crystal lattice and Bohr atomic model]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(11–12): 385–387, June 30, 1919. CODEN VDPEAZ. ISSN 0372-5448.

Born:1942:TVD

[BL42] Max Born and Kathleen Lonsdale. Temperature variation of diffuse scattering of X-rays by crystals. Nature, 150(3808):490, October 24, 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v150/
n3808/pdf/150490a0.pdf.

Born:1943:DFL

[BL43] Max Born and Walter Ledermann. Density of frequencies in lattice dynamics. Nature, 151(3824):197–198, February 13, 1943. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v151/
n3824/pdf/151197a0.pdf.

Born:1958:QKT

[BL58] Max Born and W. Ludwig. Zur Quantenmechanik des kräftefreien Teilchens. (German) [Toward a quantum mechanics of an inertial particle]. Zeitschrift für Physik, 150(1):106–117, February 1958. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01338519.

Born:1963:KBA

[BL63a] Max Born and A. Landé. Kristallgitter und Bohrsches Atommodell. (German) [Crystal lattice and Bohr’s atomic model]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 16, pages 348–355. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 20, 202–209 (1918).

Born:1963:ABK

[BL63b] Max Born and A. Landé. Über die absolute Berechnung der Kristalleigenschaften mit Hilfe Bohrscher Atommodelle. (German) [On the absolute calculation of crystal properties using Bohrian atom models]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 15, pages 327–347. LCCN ???? Reprinted from S. B. Preuß. Akad. Wiss. Berlin 1918, 1048–1068.

Born:1963:BKR

[BL63c] Max Born and A. Landé. Über die Berechnung der Kompressibilität regulärer Kristalle aus der Gittertheorie. (German) [On calculating the compressibility of regular crystals from lattice theory]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 17, pages 356–362. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 20, 210–216 (1918).

Born:1963:QKT

[BL63d] Max Born and W. Ludwig. Zur Quantenmechanik des kräftefreien Teilchens. (German) [Toward a quantum mechanics of an inertial particle]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 69, pages 442–453. LCCN ???? Reprinted from Z. Physik 150, 106–117 (1958).

Blanche:1958:BRB

[Bla58] R. Blanché. Book review: L’expérience et la théorie en physique, by Max Born and J.-P. Mathieu. Revue Philosophique de la France et de l’Étranger, 148(??):123, ???? 1958. CODEN ???? ISSN 0035-3833 (print), 2104-385X (electronic). URL http:/

/www.jstor.org/stable/41089692;   http://www.jstor.org/
stable/pdfplus/41089692.pdf.

Blagonravov:1967:CSR

[Bla67] Anatoly Blagonravov. Comments: Space research justified by needs of science. Bulletin of the Atomic Scientists, 23(8):40–42, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bor68b].

Blanchard:2008:CSE

[Bla08] Ph. Blanchard. Composite systems: Entanglement, envariance, events and Born’s rule. European Physical Journal — Special Topics, 159(??):37–45, ???? 2008. CODEN EPJSAC. ISSN 1951-6355 (print), 1951-6401 (electronic).

Born:1942:QTDa

[BLS42] Max Born, Kathleen Lonsdale, and Helen M. J. Smith. Quantum theory and diffuse X-ray reflexions. Nature, 149(??):402–??, ???? 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1932:GIG

[BM32] Max Born and Joseph E. Mayer. Zur Gittertheorie der Ionenkrystalle. (German) [Toward a lattice theory of ionic crystals]. Zeitschrift für Physik, 75(1–2):1–18, January 1932. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01340511.

Born:1940:SCLd

[BM40] Max Born and Rama Dhar Misra. On the stability of crystal lattices. IV. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36(4):466–478, October 1940. CODEN PCPSA4. ISSN 0008-1981.

Born:1963:GIG

[BM63] Max Born and J. E. Mayer. Zur Gittertheorie der Ionenkristalle. (German) [Toward a lattice theory of ionic crystals]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 29, pages 509–526. LCCN ???? Reprinted from Z. Physik 76, 1–18 (1932).

Born:1936:NTLa

[BN36a] Max Born and N. S. Nagendra Nath. The neutrino theory of light. Proceedings — Indian Academy of Sciences, Section A, 3(??):318– 337, ???? 1936. CODEN PISAA7. ISSN 0370-0089.

Born:1936:NTLb

[BN36b] Max Born and N. S. Nagendra Nath. The neutrino theory of light. II. Proceedings — Indian Academy of Sciences, Section A, 4(??): 611–620, ???? 1936. CODEN PISAA7. ISSN 0370-0089.

Badino:2013:RPH

[BN13] Massimiliano Badino and Jaume Navarro. Research and Pedagogy: a History of Quantum Physics Through Its Textbooks, volume 2 of Max Planck research library for the history and development of knowledge. Studies. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5871-X. 302 pp. LCCN QC173.98. URL http://

www.edition-open-access.de/studies/2/.

Born:1907:VWG

[BO07] Max Born and Erich Oettinger. Variationsprinzipe der Wärmetheorie. (German) [Variational principles of thermodynamics]. Physikalische Zeitschrift, 8(17):572–580, September 1, 1907. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://babel.hathitrust.org/
cgi/pt?id=mdp.39015068320954%3Bseq=604%3Bview=1up.

Born:1927:QMG

[BO27] Max Born and R. Oppenheimer. Zur Quantentheorie der Molekeln. (German) [Toward a quantum theory of molecules]. Annalen der Physik (1900), 84(20):457–484, ???? 1927. CODEN AN-PYA2. ISSN 0003-3804 (print), 1521-3889 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/andp.19273892002/

abstract. This paper introduces the now widely-used Born– Oppenheimer approximation that allows separation of electronic and nuclear motion. See also English translation [BO98].

Born:1963:QMGb

[BO63] Max Born and J. Robert Oppenheimer. Zur Quantentheorie der Molekeln. (German) [Toward a quantum theory of molecules]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 61, pages 310–337. LCCN ???? Reprinted from Ann. Physik. 84, 457–484 (1927).

Born:1998:QTM

[BO98] M. Born and R. Oppenheimer. On the quantum theory of molecules. Rev. Acad. Colombiana Cienc. Exact. Fís. Natur., 22 (84):375–391, 1998. CODEN ???? ISSN 0370-3908. Translated from the 1927 German original [BO27] by Arnulfo Poveda, José Luis Villaveces and Gloria Esperanza Moyano.

Bonitz:2005:KFE

[Bon05] Manfred Bonitz. Klaus Fuchs — the enduring contribution to physics from his British period. Scientometrics, 62(3):343–350, March 2005. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/10.
1007/s11192-005-0026-4.

Born:1906:USE

[Bor06] Max Born. Untersuchungen über die Stabilität der elastischen Linie in Ebene und Raum unter verschiedenen Grenzbedingungen. [Studies on the stability of the elastic line in the plane and in space under different limiting conditions]. Ph.D. dissertation, Universität Göttingen, Göttingen, Germany, 1906. URL http:

//genealogy.math.ndsu.nodak.edu/id.php?id=18344. Preiss-

chrift und Inaugural-Dissertation. [Prize-winning inaugural dissertation.].

Born:1909:BAT

[Bor09a] Max Born. Berichtigung zu der Arbeit: Die Theorie des starren Elektrons in der Kinematik des Relativitätsprinzips. (German) [Correction to the work: The theory of the electron in the rigid kinematics of the Principle of Relativity]. Annalen der Physik (1900), 335(15):840, ???? 1909. ISSN 1521-3889. See [Bor09b].

Born:1909:TSE

[Bor09b] Max Born. Die Theorie des starren Elektrons in der Kinematik des Relativitätsprinzips. (German) [The theory of the rigid electron in the kinematics of the Principle of Relativity]. Annalen der Physik (1900), 30(??):1–56, ???? 1909. CODEN ???? ISSN ???? See correction [Bor09a].

Born:1909:TMR

[Bor09c] Max Born. Die träge Masse und das Relativitätsprinzip. (German) [The inertial mass and the Principle of Relativity]. Annalen der Physik (1900), 28(3):571–584, ???? 1909. CODEN ???? ISSN ????

Born:1909:TAH

[Bor09d] Max Born. Über das Thomsonsche Atommodell (Habilitationsvortrag, Göttingen). (German) [On Thomson’s atomic model (habilitation lecture, Göttingen)]. Physikalische Zeitschrift, 10(25): 1031–1034, December 15, 1909. CODEN PHZTAO. ISSN

0369-982X. URL http://babel.hathitrust.org/cgi/pt?id=
mdp.39015023919049%3Bseq=1107%3Bview=1up.

Born:1909:DEKa

[Bor09e] Max Born. Über die Dynamik des Elektrons in der Kinematik des Relativitätsprinzips. (German) [On the dynamics of the electron in the kinematics of the Principle of Relativity]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 11(21):617–623, November 15, 1909. CODEN VDPEAZ. ISSN 0372-5448. Also in [Bor09f].

Born:1909:DEKb

[Bor09f] Max Born. Über die Dynamik des Elektrons in der Kinematik des Relativitätsprinzips. (German) [On the dynamics of the electron in the kinematics of the Principle of Relativity]. Physikalische Zeitschrift, 10(22):814–817, November 10, 1909. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://babel.hathitrust.org/
cgi/pt?id=mdp.39015023919049%3Bseq=868%3Bview=1up. Also

in [Bor09e].

Born:1909:VEK

[Bor09g] Max Born. Über eine Verallgemeinerung der Eulerschen Knickformel. (German) [A generalization of Euler’s buckling formula].

Physikalische Zeitschrift, 10(11):383–387, June 1, 1909. CODEN

PHZTAO. ISSN 0369-982X. URL http://babel.hathitrust.
org/cgi/pt?id=mdp.39015023919049%3Bseq=419%3Bview=1up.

Born:1910:BMA

[Bor10a] Max Born. Berichtigung zu meiner Arbeit ’Zur Elektrodynamik bewegter Körper’. (German) [Correction to my work ’on the electrodynamics moving bodies’]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 12(18–19):730, October 15, 1910. CO-DEN VDPEAZ. ISSN 0372-5448. See [Bor10d].

Born:1910:AGE

[Bor10b] Max Born. Eine Ableitung der Grundgleichungen für die elektromagnetischen Vorgänge in bewegten Körpern vom Standpunkte der Elektronentheorie (Aus dem Nachlass von Hermann Minkowski bearbeitet von Max Born). (German) [A derivation of the basic equations for electromagnetic processes in moving bodies from the standpoint of the theory of electrons (from the estate of Hermann Minkowski, edited by Max Born)]. Mathematische Annalen, 68(??): 526–551, ???? 1910. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Born:1910:DSK

[Bor10c] Max Born. Über die Definition des starren Körpers in der Kinematik des Relativitäts-prinzips. (German) [On the definition of a rigid body in the kinematics of the Principle of Relativity]. Physikalische Zeitschrift, 11(6):233–234, March 15, 1910. CODEN

PHZTAO. ISSN 0369-982X. URL http://babel.hathitrust.
org/cgi/pt?id=nyp.33433090815337%3Bseq=285%3Bview=1up.

Born:1910:EBK

[Bor10d] Max Born. Zur Elektrodynamik bewegter Körper. (German) [Toward the electrodynamics of moving bodies]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 12(10):457–467, May 30, 1910. CODEN VDPEAZ. ISSN 0372-5448. See correction [Bor10a].

Born:1910:KSK

[Bor10e] Max Born. Zur Kinematik des starren Körpers im System des Relativitätsprinzips. (German) [Kinematics of rigid bodies in the system of the Principle of Relativity]. Nachr. Ges. Wiss. Göttingen, ??(??):161–179, ???? 1910. CODEN ???? ISSN ????

Born:1911:ERG

[Bor11] Max. Born. Elastizitätstheorie und Relativitätstheorie. (German) [Elasticity theory and Relativity Theory]. Physikalische Zeitschrift, 12(14):569–575, July 15, 1911. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=mdp.39015011417709%3Bseq=623%3Bview=1up.

Born:1912:LBM

[Bor12a] Max Born. Lichtfortpflanzung in bewegten Medien. (German) [Light propagation in moving media]. In Handwörterb. Naturwiss, volume 6, pages 287–294. Gustav Fischer, Jena, Germany, 1912.

Born:1912:PP

[Bor12b] Max Born. Prinzipien der Physik. (German) [Principles of physics]. In Handwörterb. d. Naturwiss, volume 6, pages 1118–1126. Gustav Fischer, Jena, Germany, 1912.

Born:1913:TWQ

[Bor13a] Max Born. Die Theorie der Wärmestrahlung und die Quantenhypothese. (German) [The theory of heat radiation and the quantum hypothesis]. Naturwissenschaften, 1(??):499–504, ???? 1913. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1913:IGI

[Bor13b] Max Born. Infinitesimalrechnung. (German) [Infinitesimal calculation]. In Handwörterb. Naturwiss, volume 5, pages 413–425. Gustav Fischer, Jena, Germany, ???? 1913.

Born:1913:RGS

[Bor13c] Max Born. Raum. (German) [Space]. In Handwürterb. Naturwiss, volume 7, pages 120–?? Gustav Fischer, Jena, Germany, ???? 1913. CODEN ???? ISSN ????

Born:1913:REH

[Bor13d] Max Born. Zum Relativitätsprinzip: Entgegnung auf Herrn Gehrcke’s Artikel ’Die gegen die Relativitätstheorie erhobenen Einwande’. (German) [On the Principle of Relativity: reply to Mr. Gehrcke’s article ’the Theory of Relativity for the moving charges’]. Naturwissenschaften, 1(??):92–94, 191–192, ???? 1913. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1913:KTM

[Bor13e] Max Born. Zur kinetischen Theorie der Materie. (German) [Toward the kinetic theory of matter]. Naturwissenschaften, 1(??):297–299, ???? 1913. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1914:IES

[Bor14a] Max Born. Der Impuls-Energie-Satz in der Elektrodynamik von Gustav Mie. (German) [The pulse-energy theorem in Gustav Mie’s electrodynamics]. Nachr. Ges. Wiss. Göttingen, ??(??):23–37, ???? 1914. CODEN ???? ISSN ????

Born:1914:EBE

[Bor14b] Max Born. Die elektronentheoretische Begründung der Elektrodynamik bewegter Körper. (German) [The theoretical foundation of the electrodynamics of moving bodies]. Jb. Radioakt, 11(??):301– 307, ???? 1914. CODEN ???? ISSN ????

Born:1914:MET

[Bor14c] Max Born. Über die Methode der Eigenschwingungen in der Theorie der specifischen Wärme. (German) [On the method of natural oscillations in the theory of specific heat]. Physikalische Zeitschrift, 15(4):185–191, February 15, 1914. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=uc1.c2831834%3Bseq=217%3Bview=1up.

Born:1914:RDG

[Bor14d] Max Born. Zur Raumgittertheorie des Diamanten. (German) [Toward the space lattice theory of diamond]. Annalen der Physik (1900), 44(??):605–642, ???? 1914. CODEN ???? ISSN ????

Born:1915:BMH

[Bor15a] Max Born. Bemerkungen zu der Mitteilung des Herrn Karl Czukor: Zur Theorie der Dielektrika. (German) [Reponse to the communication of Mr. Charles Czukor: “Toward the theory of dielectrics”]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 17(10): 204, May 30, 1915. CODEN VDPEAZ. ISSN 0372-5448.

Born:1915:DKG

[Bor15b] Max Born. Dynamik der Kristallgitter. (German) [Dynamics of crystal lattices]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1915. vii + 122 pp. LCCN ????

Born:1915:DDP

[Bor15c] Max Born. Über die Dispersion der Drehung der Polarisationsebene in Kristallen. (German) [On the dispersion of the rotation of the plane of polarization in crystals]. Physikalische Zeitschrift, 16(23):437–438, December 1, 1915. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=njp.32101054770928%3Bseq=487%3Bview=1up.

Born:1915:NOA

[Bor15d] Max Born. Über die natürliche optische Aktivität von Flüssigkeiten und Gasen. (German) [On the natural optical activity of liquids and gases]. Physikalische Zeitschrift, 16(13–14):251–258, July 15, 1915. CODEN PHZTAO. ISSN 0369-982X. URL http://babel.

hathitrust.org/cgi/pt?id=njp.32101054770928%3Bseq=299%
3Bview=1up.

Born:1915:OAK

[Bor15e] Max Born. Über die optische Aktivität der Kristalle. (German) [On the optical activity of crystals]. In Bergwitz [Ber15], pages 391–403.

LCCN QD455 .F47 1915. URL http://catalog.hathitrust.
org/api/volumes/oclc/47733990.html.

Born:1916:ETG

[Bor16a] Max Born. Einstein’s Theorie der Gravitation und der allgemeinen Relativität. (German) [Einstein’s Theory of Gravitation and General Relativity]. Physikalische Zeitschrift, 17(4): 51–59, February 15, 1916. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015010783705%3Bseq=75%3Bview=1up.

Born:1916:AFG

[Bor16b] Max Born. Über anisotrope Flüssigkeiten. (German) [On anisotropic fluids]. S. B. preuss. Akad. Wiss. Berlin, ??(??):614– 650, ???? 1916. CODEN ???? ISSN ????

Born:1917:ZLS

[Bor17] Max Born. Über die Zerstreuung des Lichtes in Substanzen mit anisotropen Molekeln. (German) [On the scattering of light in substances with anisotropic molecules]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 19(21–22):243–264, November 30, 1917. CODEN VDPEAZ. ISSN 0372-5448.

Born:1918:EMK

[Bor18a] Max Born. Die elektromagnetische Masse der Kristalle. (German) [The electromagnetic mass of crystals]. S. B. preuss. Akad. Wiss. Berlin, ??(??):712–718, ???? 1918. CODEN ???? ISSN ????

Born:1918:ENO

[Bor18b] Max Born. Elektronentheorie des natürlichen optischen Drehungsvermögens isotroper und anisotroper Flüssigkeiten. (German) [Electron theory of natural optical rotatory power of isotropic and anisotropic liquids]. Annalen der Physik (1900), 55(??):177–240, ???? 1918. CODEN ???? ISSN ????

Born:1918:HH

[Bor18c] Max Born. H. Herkner. Naturwissenschaften, 6(??):179–??, ???? 1918. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1918:BAK

[Bor18d] Max Born. Über die Berechnung der absoluten Kristalldimensionen. (German) [On the calculation of absolute crystal dimensions]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(21– 24):224–229, December 30, 1918. CODEN VDPEAZ. ISSN 0372-5448.

Born:1918:MBZ

[Bor18e] Max Born. Über die Maxwellsche Beziehung zwischen Brechungsindex und Dielektrizitätskonstante und über eine Methode zur Bestimmung der Ionenladung in Kristallen. (German) [On the relationship between Maxwell’s refractive index and dielectric constant and a method for the determination of ionic charge of crystals]. S. B. preuss. Akad. Wiss. Berlin, ??(??):604–613, ???? 1918. CODEN ???? ISSN ????

Born:1918:UEZ

[Bor18f] Max Born. Über die ultraroten Eigenschwingungen zweiatomiger Kristalle. (German) [On the infra-red natural frequencies of diatomic crystals]. Physikalische Zeitschrift, 19(24):539–548, December 15, 1918. CODEN PHZTAO. ISSN 0369-982X. URL http://

babel.hathitrust.org/cgi/pt?id=mdp.39015056706941%3Bseq=
625%3Bview=1up.

Born:1918:ZLH

[Bor18g] Max Born. Über die Zerstreuung des Lichtes in H_{2}, O_{2} und N_{2}. (German) [On the scattering of light in H_{2}, O_{2} and N_{2}]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(1–4):16–32, February 28, 1918. CODEN VDPEAZ. ISSN 0372-5448.

Born:1918:KAG

[Bor18h] Max Born. Über kubische Atommodelle. (German) [On cubic atomic models]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 20(21–24):230–239, December 30, 1918. CODEN VD-PEAZ. ISSN 0372-5448.

Born:1919:EHG

[Bor19a] Max Born. Die elektronenaffinität der Halogenatome. (German) [The electron affinity of halogen atoms]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(21—22):679–685, December 5, 1919. CODEN VDPEAZ. ISSN 0372-5448.

Born:1919:REI

[Bor19b] Max Born. Die Relativitätstheorie Einsteins und ihre physikalischen Grundlagen. (German) [Einstein’s Relativity Theory and its Physical Basis]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1919. ???? pp. LCCN ????

Born:1919:TAG

[Bor19c] Max Born. Eine thermochemische Anwendung der Gittertheorie. (German) [A thermochemical application of lattice theory]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(1–2):13– 24, January 30, 1919. CODEN VDPEAZ. ISSN 0372-5448. This paper, with Haber’s independent work [Hab19], is the origin of the famous “Born–Haber cycle” in thermodynamics.

Born:1919:ENKa

[Bor19d] Max Born. Über die elektrische Natur der Kohäsionskräfte fester Körper. (German) [On the electrical nature of the cohesive forces of solids]. Annalen der Physik (1900), 61(??):87–106, ???? 1919. CODEN ???? ISSN ????

Born:1919:ENKb

[Bor19e] Max Born. Über die elektrische Natur der Kohäsionskräfte fester Körper. (German) [On the electrical nature of the cohesive forces of solids]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(15–16):533–538, August 30, 1919. CODEN VDPEAZ. ISSN 0372-5448.

Born:1919:MAE

[Bor19f] Max Born. Vom mechanischen Äther zur elektrischen Materie. (German) [From mechanical ether to electrical matter]. Naturwissenschaften, 7(??):136–141, ???? 1919. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1920:AGA

[Bor20a] Max Born. Das Atom. (German) [The atom]. Naturwissenschaften, 8(??):213–226, ???? 1920. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1920:AMD

[Bor20b] Max Born. Der Aufbau der Materie. Drei Aufsätze über moderne Atomistik Elektronentheorie. (German) [The structure of matter. Three essays on modern atomic electron theory]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1920. ???? pp. LCCN ????

Born:1920:BZC

[Bor20c] Max Born. Die Brücke zwischen Chemie und Physik. (German) [The bridge between chemistry and physics]. Naturwissenschaften, 8(??):373–382, ???? 1920. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1920:REI

[Bor20d] Max Born. Die Relativitätstheorie Einsteins und ihre physikalischen grundlagen gemeinverständlich dargestellt von Max Born. (German) [Einstein’s Theory of Relativity and its physical foundations]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1920. x + 242 pp. LCCN QC6 .B65.

Born:1920:DMF

[Bor20e] Max Born. Eine direkte Messung der freien Weglänge neutraler Atome. (German) [A direct measurement of the free path of neutral atoms]. Physikalische Zeitschrift, 21(21–22):578– 581, November 1–15, 1920. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015086723080%3Bseq=644%3Bview=1up.

Born:1920:EMG

[Bor20f] Max Born. Enzyklopädie der Mathematik. (German) [Encyclopedia of Mathematics]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1920. ???? pp. LCCN ????

Born:1920:KBTa

[Bor20g] Max Born. Kritische Betrachtungen zur traditionellen Darstellung der Thermodynamik. (German) [Critical considerations for traditional representation of thermodynamics]. Physikalische Zeitschrift, 22(7):218–224, April 1, 1920. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015020056829%3Bseq=254%3Bview=1up.

Born:1920:KBTb

[Bor20h] Max Born. Kritische Betrachtungen zur traditionellen Darstellung der Thermodynamik. (German) [Critical considerations for traditional representation of thermodynamics]. Physikalische Zeitschrift, 22(8):249–254, April 15, 1920. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015020056829%3Bseq=287%3Bview=1up.

Born:1920:KBTc

[Bor20i] Max Born. Kritische Betrachtungen zur traditionellen Darstellung der Thermodynamik. (German) [Critical considerations for traditional representation of thermodynamics]. Physikalische Zeitschrift, 22(9):282–286, May 1, 1920. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015020056829%3Bseq=322%3Bview=1up.

Born:1920:BEIa

[Bor20j] Max Born. Über die Beweglichkeit der elektrolytischen Ionen. (German) [On the mobility of the electrolytic ion]. Zeitschrift für Elektrochemie, 26(??):401–403, ???? 1920. CODEN ZEELAI. ISSN 0372-8382.

Born:1920:BEIb

[Bor20k] Max Born. Über die Beweglichkeit der elektrolytischen Ionen. (German) [On the mobility of the electrolytic ion]. Zeitschrift für Physik, 1(3):221–249, June 1920. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01329168.

Born:1920:VHI

[Bor20l] Max Born. Volumen und Hydratationswärme der Ionen. (German) [Volume and heat of hydration of ions]. Zeitschrift für Physik, 1 (1):45–48, February 1920. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01881023.

Born:1921:PNR

[Bor21a] Max Born. Die physikalische Natur der Röntgenstrahlen. (German) [The physical nature of X-rays]. Umschau, 25(??):29–32, ???? 1921. CODEN ???? ISSN ????

Born:1921:DMN

[Bor21b] Max Born. Über einen direkten mechanischen Nachweis des Dipolcharakters von Flüssigkeitsmolekeln. (German) [On direct mechanical dipole character of fluid molecules]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 2(??):53, ???? 1921. CO-DEN VDPEAZ. ISSN 0372-5448.

Born:1921:EGGa

[Bor21c] Max Born. Über elektrostatische Gitterpotentiale. (German) [On electrostatic lattice potentials]. Zeitschrift für Physik, 7(1):124–140, December 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01332783.

Born:1921:TKG

[Bor21d] Max Born. Zur Thermodynamik der Kristallgitter. (German) [Toward a thermodynamics of crystal lattices]. Zeitschrift für Physik, 7(1):217–248, December 1921. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01332792.

Born:1922:REI

[Bor22a] Max Born. Die Relativitätstheorie Einsteins und ihre physikalischen grundlagen. Elementar dargestellt. (German) [Einstein’s Theory of Relativity and its physical basis. Elementary treatment]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1922. xi + 267 pp. LCCN QC6 .B65.

Born:1922:HPG

[Bor22b] Max Born. Hilbert und die Physik. (German) [Hilbert and physics]. Naturwissenschaften, 10(4):88–93, January 27, 1922. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1922:TRE

[Bor22c] Max Born. La teoría de la relatividad de Einstein y sus fundamentos fisicos: Exposición elemental. (Spanish) [Einstein’s Theory of Relativity and its fundamental physics: elementary introduction]. Calpe, Madrid, Spain, 1922. 384 pp. LCCN ????

Born:1922:MWG

[Bor22d] Max Born. Über das Modell der Wasserstoffmolekel. (German) [On the model of the hydrogen molecule]. Naturwissenschaften, 10(??): 667–678, ???? 1922. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1922:NOA

[Bor22e] Max Born. Über die naturliche optische Aktivität der Kristalle. (German) [On the natural optical activity of crystals]. Zeitschrift für Physik, 8(1):390–417, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01329609.

Born:1922:TPG

[Bor22f] Max Born. Über die Temperaturabhängigkeit der Pyroelektrizität. (German) [On the temperature dependence of pyroelectricity]. Physikalische Zeitschrift, 23(5):125–128, March 1, 1922. CODEN

PHZTAO. ISSN 0369-982X. URL http://babel.hathitrust.
org/cgi/pt?id=mdp.39015086723239%3Bseq=155%3Bview=1up.

Born:1923:AFZ

[Bor23a] Max Born. Atomtheorie des festen Zustandes (Dynamik der Kristallgitter). (German) [Atomic theory of the solid state (dynamics of crystal lattices)]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1923. ???? pp. LCCN ????

Born:1923:CM

[Bor23b] Max Born. The constitution of matter, volume ?? Methuen and Co. Ltd., London, UK, 1923. ???? pp. LCCN ????

Born:1923:QSGa

[Bor23c] Max Born. Quantentheorie und Störungsrechnung. (German) [Quantum theory and perturbation calculation]. Naturwissenschaften, 11(27):537–542, July 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1923:QSGb

[Bor23d] Max Born. Quantentheorie und Störungsrechnung. (German) [Quantum theory and perturbation calculation]. Naturwissenschaften, 10(??):677–678, ???? 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1924:EIP

[Bor24a] Max Born. Der Einfluß der Ionendeformation auf physikalische und chemische Konstanten. (German) [The influence of the deformability of ions on optical and chemical constants.]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 5(??):8–9, ???? 1924. CO-DEN VDPEAZ. ISSN 0372-5448.

Born:1924:CBD

[Bor24b] Max Born. Die chemische Bindung als dynamisches Problem. (German) [The chemical bond as a dynamic problem]. Naturwissenschaften, 12(??):1199–1207, ???? 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1924:RRT

[Bor24c] Max Born. Recherches récentes sur la théorie de l’affinité chimique. (French) [Recent research on the theory of chemical affinity]. Scientia (Milan), 36(??):155–168, ???? 1924. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic).

Born:1924:AGA

[Bor24d] Max Born. Über Atomtheorie. (German) [On atomic theory]. Elektrotech. Z., 45(??):889–897, ???? 1924. CODEN ???? ISSN ????

Born:1924:EDC

[Bor24e] Max Born. Über die elektrische Deutung der chemischen Krafte. (German) [On the electrical interpretation of chemical forces]. Zeitschrift für Elektrochemie, 30(??):382–386, ???? 1924. CODEN ZEELAI. ISSN 0372-8382.

Born:1924:QGQa

[Bor24f] Max Born. Über Quantenmechanik. (German) [On quantum mechanics]. Zeitschrift für Physik, 26(1):379–395, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01327341. Kemmer and Schlapp [KS71, p. 33] report that this may the earliest published paper with “quantum mechanics” in its title. English translation in [Bor67c].

Born:1926:PAD

[Bor26a] M. Born. Problems of atomic dynamics. I: The structure of the atom (20 lectures). II: The lattice theory of rigid bodies (10 lectures). MIT Press, Cambridge, MA, USA, 1926. xiv + 200

pp. URL http://mitpress.mit.edu/about/history;   http:
//mitpress.mit.edu/books/problems-atomic-dynamics.

Born:1926:AQG

[Bor26b] Max Born. Das Adiabatenprinzip in der Quantenmechanik. (German) [The adiabatic principle in quantum mechanics]. Zeitschrift für Physik, 40(3–4):167–192, March 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01400360.

Born:1926:PAS

[Bor26c] Max Born. Probleme der Atomdynamik. I. Die Struktur des Atoms. II. Die Gittertheorie des festen Zustandes. (German) [Problems of atomic dynamics. I. The structure of the atom. II. The lattice theory of the solid state]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1926. 180 pp. LCCN ????

Born:1926:QSGb

[Bor26d] Max Born. Quantenmechanik der Stoßvorgänge. (German) [Quantum mechanics of collision processes]. Zeitschrift für Physik, 38(11– 12):803–827, November 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01397184.

Born:1926:QSGa

[Bor26e] Max Born. Zur Quantenmechanik der Stoßvorgänge. (German) [Toward a quantum mechanics of collision processes]. Zeitschrift für Physik, 37(12):863–867, December 1926. CODEN ZEPYAA.

ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01397477.

Born:1926:WSG

[Bor26f] Max Born. Zur Wellenmechanik der Stoßvorgänge. (German) [Toward a wave mechanics of collision processes]. Nachr. Ges. Wiss. Göttingen, ??(??):146–160, ???? 1926. CODEN ???? ISSN ????

Born:1927:AKG

[Bor27a] Max Born. Atomkonstanten und Körpereigenschaften. (German) [Atomic constants and physical properties]. Z. math. naturw. Unterr, 40(??):241–253, ???? 1927. CODEN ???? ISSN ????

Born:1927:PAQ

[Bor27b] Max Born. Physical aspects of quantum mechanics. Nature, 119 (2992):354–357, March 5, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v119/n2992/pdf/119354a0.pdf.

Born:1927:QSG

[Bor27c] Max Born. Quantenmechanik und Statistik. (German) [Quantum mechanics and statistics]. Naturwissenschaften, 15(??):238–242, ???? 1927. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1928:AL

[Bor28a] Max Born. Antoon Lorentz. Nachr. Ges. Wiss. Göttingen, ??(??): 69–73, ???? 1928. CODEN ???? ISSN ????

Born:1928:KAR

[Bor28b] Max Born. Kongress der Assoziation der russischen Physiker. (German) [Congress of the Association of Russian Physicists]. Naturwissenschaften, 16(??):741–743, ???? 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1928:SBS

[Bor28c] Max Born. Sommerfeld als Begründer einer Schule. (German) [Sommerfeld as founder of a school]. Naturwissenschaften, 16(49):1035– 1036, December 7, 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1928:SPT

[Bor28d] Max Born. Über den Sinn der physikalischen Theorien. (German) [On the meaning of physical theories]. Nachr. Ges. Wiss. Göttingen, ??(??):51–70, ???? 1928. CODEN ???? ISSN ????

Born:1928:BSV

[Bor28e] Max Born. Über die Bedeutung der Stoßvorgänge für das Verständnis der Quantenmechanik. (German) [On the importance of collision processes for the understanding of quantum mechanics]. In Atti Congr. Intern. dei Fisici Como-Pavia-Roma 1927, volume 2, pages 443–447. ????, ????, 1928.

Born:1928:TRG

[Bor28f] Max Born. Zur Theorie des Ramaneffektes. (German) [Toward a theory of the Raman effect]. Naturwissenschaften, 16(??):673, ???? 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1929:QTC

[Bor29a] Max Born. The quantum theory of chemical valence. Nature, 126(??):205–??, ???? 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1929:SPT

[Bor29b] Max Born. Über den Sinn der physikalischen Theorien. (German) [On the meaning of physical theories]. Naturwissenschaften, 17(??): 109–118, ???? 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1929:QCK

[Bor29c] Max Born. Zur Quantentheorie der chemischen Krafte. (German) [Toward a quantum theory of chemical forces]. Zeitschrift für Physik, 64(11–12):729–740, September 22, 1929. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01773002.

Born:1929:TKGa

[Bor29d] Max Born. Zur Theorie des Kernzerfalls. (German) [Toward the theory of nuclear disintegration]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 10(??):31–32, ???? 1929. CODEN VD-PEAZ. ISSN 0372-5448.

Born:1929:TKGb

[Bor29e] Max Born. Zur Theorie des Kernzerfalls. (German) [Toward the theory of nuclear disintegration]. Zeitschrift für Physik, 58 (5–6):306–321, May 1929. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01340382.

Born:1930:QTC

[Bor30a] M. Born. The quantum theory of chemical valence free. Nature, 126(3171):205, August 9, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v126/n3171/full/126205d0.html.

Born:1930:QCB

[Bor30b] Max Born. Zur Quantentheorie der chemischen Bindung. (German) [Toward a quantum theory of chemical bonding]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 11(??):34–??, ???? 1930. CODEN VDPEAZ. ISSN 0372-5448.

Born:1931:ATH

[Bor31a] Max Born. The application of the theory of homopolar valency to polyatomic molecules. In Chemistry at the Centenary Meeting of the British Association, pages 249–254. Heffer, Cambridge, UK, 1931. CODEN ???? ISSN ????

Born:1931:CBQ

[Bor31b] Max Born. Chemische Bindung und Quantenmechanik. (German) [Chemical binding and quantum mechanics]. Ergebnisse der Exakten Naturwissenschaften, 10(??):387–444, ???? 1931. CODEN EENAA3. ISSN 0367-0325.

Born:1931:GPC

[Bor31c] Max Born. Das Grenzgebiet von Physik und Chemie. (German) [The border region of physics and chemistry]. Umschau, 35(??): 509–511, ???? 1931. CODEN ???? ISSN ????

Born:1931:FJB

[Bor31d] Max Born. Faraday-Jahrhundertfeier und British Association im London. (German) [Faraday centenary and the British Association London]. Naturwissenschaften, 19(??):932–934, ???? 1931. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1931:QPM

[Bor31e] Max Born. Quelques problèmes de mécanique quantique. (French) [Some problems with quantum mechanics]. Ann. Inst. Poincaré, 1 (??):205–263, ???? 1931. CODEN ???? ISSN ????

Born:1931:WCK

[Bor31f] Max Born. Was sind die chemischen Kräfte?. (German) [What are the chemical forces?]. Umschau (1897), 35(??):532–534, ???? 1931. CODEN UMSCAS. ISSN 0372-4409.

Born:1932:UEA

[Bor32a] Max Born. Die ultraroten Eigenfrequenzen der Alkalihalogenidkristalle. (German) [The infra-red natural frequencies of alkali halide crystals]. Zeitschrift für Physik, 76(7–8):559–560, July 1932. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01336737.

Born:1932:BEG

[Bor32b] Max Born. Eine Bemerkung über den Elektronenradius. (German) [A note on the electron radius]. Naturwissenschaften, 20(??):269, ???? 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1932:GUA

[Bor32c] Max Born. Geometrische und undulatorische Abbildungsfehler der optischen Instrumente. (German) [Geometrical and undulatory aberrations in optical instruments]. Naturwissenschaften, 20(??): 921–923, ???? 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1932:DUA

[Bor32d] Max Born. Zur Deutung der ultravioletten Absorptionsbanden der Alkalihalogenide. (German) [Toward the interpretation of the ultraviolet absorption bands of the alkali halides]. Zeitschrift für Physik, 79(1–2):62–68, January 1932. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01349501.

Born:1932:THV

[Bor32e] Max Born. Zur Theorie der homöopolaren Valenz bei mehratomigen Molekülen. (German) [Toward the theory of homopolar valence in polyatomic molecules]. Zeitschrift für Angewandte Chemie, 45 (??):6–8, ???? 1932. CODEN ???? ISSN ????

Born:1933:MPS

[Bor33a] Max Born. Moderne Physik: Sieben Vorträge über Materie und Strahlung. (German) [Modern Physics: Seven Lectures on Matter and Radiation]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933. vii + 272 pp. LCCN ????

Born:1933:MFE

[Bor33b] Max Born. Modified field equations with a finite radius of the electron. Nature, 132(3329):282, August 19, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v132/n3329/pdf/132282a0.pdf.

Born:1933:OLE

[Bor33c] Max Born. Optik: Ein Lehrbuch der elektromagnetischen Lichttheorie. (German) [Optics: a textbook of the electromagnetic theory of light]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1933. vii + 591 pp. LCCN ????

Born:1934:CRN

[Bor34a] Max Born. Cosmic rays and the new field theory. Nature, 133 (3350):63–64, January 13, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v133/n3350/pdf/133063b0.pdf.

Born:1934:QTE

[Bor34b] Max Born. On the quantum theory of the electromagnetic field. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 143(849):410–437, January 1, 1934. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/96101;   http://www.jstor.org/stable/
pdfplus/96101.pdf.

Born:1934:QE

[Bor34c] Max Born. Quantum electrodynamics. In Papers and Discussions of the International of Conference on Physics. London. Vol. I, pages 19–27. ????, ????, 1934. CODEN ???? ISSN ????

Born:1935:AP

[Bor35a] Max Born. Atomic physics. Blackie, Glasgow, Scotland, 1935. ???? pp. LCCN ????

Born:1935:MN

[Bor35b] Max Born. The mysterious number 137. Proceedings — Indian Academy of Sciences, Section A, 2(??):533–561, ???? 1935. CO-DEN PISAA7. ISSN 0370-0089.

Born:1935:TOA

[Bor35c] Max Born. On the theory of optical activity. I. General theory of a system of coupled isotropic oscillators. II. Molecules with a binary axis of symmetry. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150(869):84– 105, May 1, 1935. CODEN PRLAAZ. ISSN 0080-4630. URL

http://rspa.royalsocietypublishing.org/content/150/869/
84;   http://www.jstor.org/stable/pdfplus/96325.pdf. See

erratum [Bor36a].

Born:1935:QFT

[Bor35d] Max Born. Quantised field theory and the mass of the proton. Nature, 136(3450):952–953, December 14, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v136/n3450/pdf/136952a0.pdf.

Born:1935:QE

[Bor35e] Max Born. Quantum electrodynamics. In Internat. Conf. on Physics, volume 1, pages 19–27. ????, ????, 1935.

Born:1935:RU

[Bor35f] Max Born. The Restless Universe. Blackie, Glasgow, Scotland, 1935. ???? pp. LCCN ????

Born:1936:ETO

[Bor36a] Max Born. Erratum: On the theory of optical activity. I. General theory of a system of coupled isotropic oscillators; II. Molecules with a binary axis of symmetry. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 153(879): 339–??, January 1, 1936. CODEN PRLAAZ. ISSN 0080-4630. See [Bor35c].

Born:1936:LED

[Bor36b] Max Born. On the linearization of the energy density of the electromagnetic field. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 32(1):102–107, January 1936. CODEN PCPSA4. ISSN 0008-1981.

Born:1936:SPA

[Bor36c] Max Born. Some philosophical aspects of modern physics (Inaugural Lecture as Tait Professor of Natural Philosophy, University of Edinburgh). Proceedings of the Royal Society of Edinburgh, 57 (??):1–18, ???? 1936. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Born:1936:UTFa

[Bor36d] Max Born. Unitary theory of field and matter I. Classical treatment. Charged particle with magnetic rest-moment. Proc. Indian Acad. Sci, 3(??):8–24, ???? 1936. CODEN ???? ISSN ????

Born:1936:UTFb

[Bor36e] Max Born. Unitary theory of field and matter. II. Classical treatment. Charged particle with electric and magnetic moment. Proceedings — Indian Academy of Sciences, Section A, 3(??):85–97, ???? 1936. CODEN PISAA7. ISSN 0370-0089.

Born:1937:BRB

[Bor37a] Max Born. Book review: The Theory of the Properties of Metals and Alloys, by N. F. Mott and H. Jones. Mathematical Gazette, 21(242):61, February 1937. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.jstor.org/

stable/pdfplus/3605761.pdf.

Born:1937:SPA

[Bor37b] Max Born. Some philosophical aspects of modern physics. Proceedings of the Royal Society of Edinburgh, 57(??):1–18, January 1937. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Born:1937:SMC

[Bor37c] Max Born. The statistical mechanics of condensing systems. Physica, 4(??):1034–1044, ???? 1937. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Born:1937:WMC

[Bor37d] Max Born. Wave mechanics of couples (neutron–neutrino). Nature, 139(3506):68, January 9, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v139/n3506/pdf/139068a0.pdf.

Born:1938:ARN

[Bor38a] Max Born. Application of “reciprocity” to nuclei. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 166(927):552–557, June 16, 1938. CODEN PRLAAZ.

ISSN 0080-4630. URL http://rspa.royalsocietypublishing.
org/content/166/927/552.

Born:1938:RQT

[Bor38b] Max Born. Relativity and quantum theory. Nature, 141(3564): 327–328, February 19, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v141/n3564/pdf/141327a0.pdf.

Born:1938:SRR

[Bor38c] Max Born. Some remarks on reciprocity. Proceedings — Indian Academy of Sciences, Section A, 8(??):309–314, ???? 1938. CO-DEN PISAA7. ISSN 0370-0089.

Born:1938:SUQ

[Bor38d] Max Born. A suggestion for unifying quantum theory and relativity. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 165(921):291–303, April 5, 1938. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/165/921/291.

Born:1938:TNL

[Bor38e] Max Born. Théorie non-linéaire du champ électromagnetique. (French) [Nonlinear theory of the electromagnetic field]. Ann. Inst. Poincaré, 7(??):155–265, ???? 1938. CODEN ???? ISSN ????

Born:1938:TNK

[Bor38f] Max Born. The Twenty-Ninth Kelvin Lecture: “Statistical Laws of Nature”. Journal of the Institution of Electrical Engineers, 83(504): 802–813, ???? 1938. CODEN JISEAL. ISSN 0368-2692. Lecture delivered before The Institution, 28th April, 1938.

Born:1939:CPEb

[Bor39a] Max Born. Cause, purpose, and economy in natural laws (minimum principles in physics). Proceedings of the Royal Institution of Great Britain, 30(3):596–628, February 10, 1939. CODEN PIG-

BAI. ISSN 0035-8959. URL http://hdl.handle.net/2027/uc1.
32106020416217?urlappend=%3Bseq=202.

Born:1939:CPEa

[Bor39b] Max Born. Cause, purpose and economy of natural laws — minimum principles in physics. Nature, 143(3618):357–361, March 4, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v143/
n3618/pdf/143357a0.pdf.

Born:1939:GBS

[Bor39c] Max Born. Gittertheoretische Behandlung des Schmelzprozesses. (German) [Lattice-theoretical treatment of melting processes]. C. R. Inst. Sci. Roum., 3(??):405–411, ???? 1939. CODEN ???? ISSN ????

Born:1939:RN

[Bor39d] Max Born. Reciprocity and the number 137. I. Proceedings of the Royal Society of Edinburgh, 59(??):219–223, ???? 1939. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Born:1939:TCM

[Bor39e] Max Born. Thermodynamics of crystals and melting. Journal of Chemical Physics, 7(8):591–603, ???? 1939. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://jcp.

aip.org/resource/1/jcpsa6/v7/i8/p591_s1;   http://link.
aip.org/link/?JCP/7/591/1.

Born:1940:OPO

[Bor40a] M. Born. Obituary: Prof. Otto Toeplitz. Nature, 145(3677): 617, April 20, 1940. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v145/n3677/pdf/145617a0.pdf.

Born:1940:ATP

[Bor40b] Max Born. Aspects of theoretical physics [review of five books]. Nature, 145(??):528–530, ???? 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1940:LTP

[Bor40c] Max Born. Book reviews: (1) Lehrbuch der theoretischen Physik (2) Lectures on Quantum Mechanics (3) Atombau und Spektrallinien (4) La Chimie Mathématique (5) Hand- und Jahrbuch der chemischen Physik. Nature, 145(3675):528–530, April 6, 1940.

CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v145/
n3675/pdf/145528a0.pdf.

Born:1940:OSJ

[Bor40d] Max Born. Obituary: Sir J. J. Thomson, O.M., F.R.S. Nature, 146(3698):356–357, September 14, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v146/n3698/pdf/146356b0.pdf.

Born:1940:SCLa

[Bor40e] Max Born. On the stability of crystal lattices. I. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36(2):160–172, April 1940. CODEN PCPSA4. ISSN 0008-1981.

Born:1940:OT

[Bor40f] Max Born. Otto Toeplitz. Nature, 145(??):617–??, ???? 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1940:RBB

[Bor40g] Max Born. Relation between breaking and melting. Nature, 145 (3680):741–742, May 11, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v145/n3680/pdf/145741b0.pdf.

Born:1940:SJJ

[Bor40h] Max Born. Sir J. J. Thomson. Nature, 146(??):356–??, ???? 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1941:DRX

[Bor41a] Max Born. Diffuse reflexion of X-rays. Nature, 147(3735):674, May 31, 1941. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v147/n3735/pdf/147674a0.pdf.

Born:1941:SJJ

[Bor41b] Max Born. Sir J. J. Thomson. Proceedings of the Physical Society, London, 53(1):305–310, May 1, 1941. CODEN PPSOAU. ISSN ????

Born:1942:DAB

[Bor42a] Max Born. Dr Arnold Berliner. Nature, 150(??):284–??, ???? 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1942:ETV

[Bor42b] Max Born. Effect of thermal vibrations on the scattering of Xrays. III. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 180(983):397–413, July 3, 1942. CODEN PRLAAZ. ISSN 0080-4630. URL http://rspa.

royalsocietypublishing.org/content/180/983/397.

Born:1942:LDX

[Bor42c] Max Born. Lattice dynamics and X-ray scattering. Proceedings of the Physical Society, London, 54(4):362–376, July 1, 1942. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (electronic). URL

http://iopscience.iop.org/0959-5309/54/4/304.

Born:1942:ODA

[Bor42d] Max Born. Obituaries [Dr. Arnold Berliner]. Nature, 150(3801): 284–285, September 5, 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v150/n3801/pdf/150284b0.pdf.

Born:1942:SCL

[Bor42e] Max Born. On the stability of crystal lattices. IX. Covariant theory of lattice deformations and the stability of some hexagonal lattices. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 38(1):82–99, January 1942. CODEN PCPSA4. ISSN 0008-1981. See corrigenda [Bor44c].

Born:1942:TTD

[Bor42f] Max Born. On the theory of temperature diffuse scattering. Physical Review, 61(5–6):377–378, March 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.61.377;   http://prola.aps.
org/abstract/PR/v61/i5-6/p377_1.

Born:1942:QTDb

[Bor42g] Max Born. Quantum theory and diffuse X-ray reflexions. Nature, 149(??):403–??, ???? 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1942:TTP

[Bor42h] Max Born. The teaching of theoretical physics in universities. Reports on Progress in Physics, 8(??):1–10, ???? 1942. CODEN RP-PHAG. ISSN 0034-4885 (print), 1361-6633 (electronic).

Born:1943:ETP

[Bor43a] Max Born. Experiment and theory in physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1943. ???? pp. LCCN ????

Born:1943:OLE

[Bor43b] Max Born. Optik: Ein Lehrbuch der elektromagnetischen Lichttheorie. (German) [Optics: a textbook of the electromagnetic theory of light]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1943. vii + 591 pp. LCCN ????

Born:1943:TIR

[Bor43c] Max Born. Theoretical investigations on the relation between crystal dynamics and X-ray scattering. Reports on Progress in Physics, 9(??):294–333, ???? 1943. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (electronic).

Born:1943:TCLa

[Bor43d] Max Born. The thermodynamics of crystal lattices. I. Discussion of the methods of calculation. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 39(2):100– 103, July 1943. CODEN PCPSA4. ISSN 0008-1981.

Born:1944:BRSa

[Bor44a] Max Born. Book review: Schrödinger: Statistical thermodynamics. Nature, 154(??):782–??, ???? 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1944:BRSb

[Bor44b] Max Born. Book review: Schrödinger: Statistical thermodynamics. Nature, 157(??):825–??, ???? 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1944:CPS

[Bor44c] Max Born. Corrigenda to the paper “On the stability of crystal lattices. IX”. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 40(3):262–263, November 1944. CODEN PCPSA4. ISSN 0008-1981. See [Bor42e].

Born:1944:RPP

[Bor44d] Max Born. Recent progress in the physics of the solid state. Coal Research, ??(??):39–47, September 1944. CODEN ???? ISSN ????

Born:1944:ST

[Bor44e] Max Born. Statistical thermodynamics. Nature, 154(3921):782– 783, December 23, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v154/n3921/pdf/154782a0.pdf.

Born:1944:UTP

[Bor44f] Max Born. Unification of the theories of photon and meson. Nature, 154(3920):764–765, December 16, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v154/n3920/pdf/154764a0.pdf.

Born:1945:TP

[Bor45a] M. Born. Theoretical physics. Nature, 156(3959):325–326, September 15, 1945. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v156/n3959/pdf/156325a0.pdf.

Born:1945:DC

[Bor45b] Max Born. Dublin Colloquium, 1945. Nature, 156(3972):704–706, December 15, 1945. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v156/n3972/pdf/156704a0.pdf.

Born:1945:QTP

[Bor45c] Max Born. On the quantum theory of pyroelectricity. Reviews of Modern Physics, 17(2–3):245–251, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.17.245.

Born:1945:TPU

[Bor45d] Max Born. Theoretical physics in the U.S.S.R. Nature, 156(??): 325–??, ???? 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1946:AP

[Bor46a] Max Born. Atomic physics. Hafner, New York, NY, USA, ???? edition, 1946.

Born:1946:BRB

[Bor46b] Max Born. Book review: Brillouin: Wave propagation in periodic structures. Nature, 158(??):926–??, ???? 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1946:BRS

[Bor46c] Max Born. Book reviews: Schrödinger: Statistical thermodynamics; Heitler: Elementary wave mechanics. Nature, 157(??):825–??, ???? 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1946:ECD

[Bor46d] Max Born. Elastic constants of diamond. Nature, 157(3992): 582, May 4, 1946. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v157/n3992/pdf/157582a0.pdf.

Born:1946:ECI

[Bor46e] Max Born. Elastic constants of ice. Nature, 158(4023):830–831, December 7, 1946. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v158/n4023/pdf/158830b0.pdf.

Born:1946:FPT

[Bor46f] Max Born. The free path for the transfer of energy in crystals. Proc. Math. Phys. Soc. Egypt, 3(??):35–41, ???? 1946. CODEN ???? ISSN ????

Born:1946:OPH

[Bor46g] Max Born. Obituary: Prof. Heinrich Rausch v. Traubenberg. Nature, 157(3985):328, March 16, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v157/n3985/pdf/157328a0.pdf.

Born:1946:PHR

[Bor46h] Max Born. Professor Heinrich Rausch von Traubenberg. Nature, 157(??):328–??, ???? 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1946:RER

[Bor46i] Max Born. Raman effect in rock-salt. Nature, 157(3998):810, June 15, 1946. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v157/n3998/pdf/157810a0.pdf.

Born:1946:STE

[Bor46j] Max Born. Statistical thermodynamics elementary wave mechanics. Nature, 157(3999):825, June 22, 1946. CODEN NATUAS.

ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v157/n3999/pdf/157825c0.pdf.

Born:1946:WPP

[Bor46k] Max Born. Wave propagation in periodic structures. Nature, 158 (4026):926, December 28, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v158/n4026/pdf/158926a0.pdf.

Born:1947:LE

[Bor47a] M. Born. [letters to editor]. Nature, 159(4034):266–267, February 22, 1947. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v159/n4034/pdf/159266b0.pdf.

Born:1947:OPV

[Bor47b] M. Born. Obituary: Prof. V. M. Goldschmidt, For.Mem.R.S. Nature, 159(4047):701, May 24, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v159/n4047/pdf/159701a0.pdf.

Born:1947:AEU

[Bor47c] Max Born. Atomic energy and its use in war and peace (Lecture). Fouad I. Univ. Press, Cairo, Egypt, ???? 1947. ???? pp.

Born:1947:BRP

[Bor47d] Max Born. Book review: Pauli: Meson theory of nuclear forces. Nature, 160(??):418–??, ???? 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1947:MTN

[Bor47e] Max Born. Meson theory of nuclear forces. Nature, 160(4065): 418, September 27, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v160/n4065/pdf/160418b0.pdf.

Born:1947:PVM

[Bor47f] Max Born. Professor V. M. Goldschmidt, For. Mem. R. S. Nature, 159(??):700–??, ???? 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1947:RSR

[Bor47g] Max Born. Raman spectrum of rock-salt under high resolution. Nature, 159(??):266–??, ???? 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1948:MKE

[Bor48a] M. Born. Max Karl Ernst Ludwig Planck. 1858–1947. Obituary Notices of Fellows of the Royal Society, 6(17):161–188, November 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (elec-

tronic). URL http://jstor.org/doi/10.2307/768916;   http:
//www.jstor.org/stable/pdfplus/768916.pdf.

Born:1948:QZH

[Bor48b] Max Born. Die Quantenmechanik und der zweite Hauptsatz der Thermodynamik. (German) [Quantum mechanics and the Second Law of Thermodynamics]. Annalen der Physik (Berlin), 438(1): 107–114, ???? 1948. CODEN ANPYA2. ISSN 0003-3804 (print), 1521-3889 (electronic).

Born:1948:MKP

[Bor48c] Max Born. Max K. E. L. Planck. Obituary Notices of Fellows of the Royal Society, 6(??):161–188, ???? 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Born:1948:RQM

[Bor48d] Max Born. Relativistic quantum mechanics and the principle of reciprocity. In Rep. Internat. Conf. ’Fundamental particles’, volume 1, pages 14–21. Physical Society, London, UK, ???? 1948.

Born:1949:L

[Bor49a] M. Born. The lectures. Nature, 164(4161):165, July 30, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v164/
n4161/pdf/164165a0.pdf.

Born:1949:EST

[Bor49b] Max Born. Einstein’s statistical theories. In Schilpp [Sch49a], chapter 5, pages 161–177. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Born:1949:EPP

[Bor49c] Max Born. Elementary particles and the principle of reciprocity. Nature, 163(4136):207–208, February 5, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v163/n4136/pdf/163207a0.pdf.

Born:1949:FQS

[Bor49d] Max Born. The foundations of quantum statistics. Nuovo Cimento, Supplemento, 6(??):161–166, ???? 1949. CODEN NUCUAF. ISSN 0550-3868.

Born:1949:SPT

[Bor49e] Max Born. Le second principe de la thermodynamique déduit de la théorie des quanta. (French) [The Second Law of Thermodynamics derived from quantum theory]. Annales de l’Institut Henri Poincaré, 11(1):1–13, ???? 1949. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic).

Born:1949:NPC

[Bor49f] Max Born. Natural Philosophy of Cause and Chance. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1949. ???? pp. LCCN ????

Born:1949:RTEb

[Bor49g] Max Born. Reciprocity theory of elementary particles. Reviews of Modern Physics, 21(3):463–473, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.21.463.

Born:1949:SOR

[Bor49h] Max Born. The second-order Raman effect in crystals, in particular diamond. J. Chim. phys, 46(??):6–8, ???? 1949. CODEN ???? ISSN ????

Born:1949:TTT

[Bor49i] Max Born. Two topics in theoretical physics (review of two lectures by W. Heisenberg). Nature, 164(4161):165, July 30, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v164/
n4161/pdf/164165a0.pdf.

Born:1950:ER

[Bor50a] Max Born. Einstein and Relativity. Nature, 166(4227):751, November 4, 1950. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v166/n4227/pdf/166751a0.pdf.

Born:1950:ERR

[Bor50b] Max Born. Einstein and Relativity: Review of Einstein: The Meaning of Relativity, 4th edition. Nature, 166(??):651–??, ???? 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1950:NLF

[Bor50c] Max Born. Non-localizable fields and reciprocity. Nature, 165 (4190):269–270, February 18, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v165/n4190/pdf/165269a0.pdf.

Born:1950:PLW

[Bor50d] Max Born. A physicist looks at world affairs: Review of Einstein: Out of my later years. Nature, 166(4235):1085, December 30, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v166/
n4235/pdf/1661085a0.pdf.

Born:1950:PHC

[Bor50e] Max Born. Physics (a half-century review). Scientific American, 183(3):28–31, September 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v183/n3/pdf/
scientificamerican0950-28.pdf.

Born:1950:PMJa

[Bor50f] Max Born. Physics and metaphysics (Joule Memorial Lecture). Science News, 17(??):9–27, ???? 1950. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Born:1950:PMJb

[Bor50g] Max Born. Physics and metaphysics (Joule Memorial Lecture). Mem. Manch. lit. and phil. Soc, 91(??):35–53, ???? 1950. CODEN ???? ISSN ????

Born:1951:ASa

[Bor51a] Max Born. Arnold Sommerfeld. Proceedings of the Physical Society, London, 64A(??):1148–1149, ???? 1951. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (electronic).

Born:1951:ASb

[Bor51b] Max Born. Arnold Sommerfeld. Proceedings of the Physical Society, London, 64B(??):1097–1098, ???? 1951. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171 (electronic).

Born:1951:AP

[Bor51c] Max Born. Atomic physics. Blackie, Glasgow, Scotland, fifth edition, ???? 1951. xiv + 437 pp. Revised by the author from the original translation of John Dougall.

Born:1951:BRS

[Bor51d] Max Born. Book review: Schrödinger: Space-time structure. Nature, 167(??):786–??, ???? 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Born:1951:GTI

[Bor51e] Max Born. Die Gültigkeitsgrenze der Theorie der idealen Kristalle und ihre Überwindung. (German) [The limits of validity of the theory of ideal crystals, and its surmounting]. In Festschr. Akad. Wiss. Göttingen 1951, Math.-Phys. Kl., pages 1–16. Akad. Wiss. Göttingen, Göttingen, Germany, ???? 1951. CODEN ???? ISSN ????

Born:1951:FYP

[Bor51f] Max Born. Fifty years of physics. Science News, 19(??):46–60, ???? 1951. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Born:1951:KEK

[Bor51g] Max Born. Kopplung der Elektronen- und Kernbewegung in Molekeln und Kristallen. (German) [Coupling of electron and nuclear motion in molecules and crystals]. Nachr. Ges. Wiss. Göttingen, 6(??):??, ???? 1951. CODEN ???? ISSN ????

Born:1951:PLF

[Bor51h] Max Born. Physics in the last fifty years. Nature, 168(4276):625– 630, October 13, 1951. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v168/n4276/pdf/168625a0.pdf.

Born:1951:RU

[Bor51i] Max Born. The restless universe. Dover Publications, Inc., New York, NY, USA, second edition, 1951. 315 pp. LCCN ????

Born:1951:SS

[Bor51j] Max Born. Space–time structure. Nature, 167(4255):786–787, May 19, 1951. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v167/n4255/pdf/167786b0.pdf.

Born:1952:SGG

[Bor52a] Hedwig Born. Stille Gänge. (German) [Silent Corridors]. Leonard Friedrich Verlag, Bad Pyrmont, West Germany, 1952. 27 pp.

Born:1952:AJW

[Bor52b] Max Born. Arnold Johannes Wilhelm Sommerfeld. 1868–1951. Obituary Notices of Fellows of the Royal Society, 8(21):274–296, November 1952. CODEN ???? ISSN 1479-571X (print), 2053-

9118 (electronic). URL http://www.jstor.org/stable/768813;
 http://www.jstor.org/stable/pdfplus/768813.pdf.

Born:1952:DNT

[Bor52c] Max Born. Dirac’s new theory of the electron. Nature, 169(4313): 1105, June 28, 1952. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v169/n4313/pdf/1691105a0.pdf.

Born:1952:NPV

[Bor52d] Max Born. Un nouveau point de vue sur la théorie de la matière condensée. C. R. 2. Réunion Chim. phys. Paris, 2(??):334–337, ???? 1952. CODEN ???? ISSN ????

Born:1953:AR

[Bor53a] Max Born. Astronomical recollections. In Vistas in astronomy, page ?? Pergamon, New York, NY, USA, 1953.

Born:1953:BRB

[Bor53b] Max Born. Book review: Philosophic Problems of Nuclear Science. Eight Lectures, by W. Heisenberg and F. C. Hayes. The Philosophical Quarterly, 3(10):88–90, January 1953. CODEN ???? ISSN 0031-8094 (print), 1467-9213 (electronic). URL http://www.

jstor.org/stable/pdfplus/2216712.pdf.

Born:1953:CSP

[Bor53c] Max Born. The conceptual situation in physics and the prospects of its future development (37th Guthrie Lecture to the Physical Society). Proceedings of the Physical Society, London, Section A, 66(6):501–513, June 1, 1953. CODEN PPSAAM. ISSN 0370-1298 (print), 1747-3829 (electronic).

Born:1953:IQM

[Bor53d] Max Born. The interpretation of quantum mechanics. The British Journal for the Philosophy of Science, 4(14):95–106, August 1953. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/
IV/14/95.full.pdf+html;   http://www.jstor.org/stable/
685986;   http://www.jstor.org/stable/pdfplus/685986.pdf.

Born:1953:OLE

[Bor53e] Max Born. Optik: Ein Lehrbuch der elektromagnetischen Lichttheorie. (German) [Optics: a textbook of the electromagnetic theory of light]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1953. vii + 591 pp. LCCN ????

Born:1953:PR

[Bor53f] Max Born. Physical reality. The Philosophical Quarterly, 3(11): 139–149, April 1953. CODEN ???? ISSN 0031-8094 (print), 1467-

9213 (electronic). URL http://www.jstor.org/stable/2216882;
 http://www.jstor.org/stable/pdfplus/2216882.pdf.

Born:1953:TBF

[Bor53g] Max Born. Theoretische Bemerkungen zu Freundlichs Formeln für die stellare Rotverschiebung. (German) [Theoretical remarks on Freundlich’s formula for the stellar red shift]. Nachr. Akad. Wiss. Göttingen, Math.-Phys. Kl., Math.-Phys.-Chem. Abt., ??(??):102– 108, ???? 1953. CODEN ???? ISSN ????

Born:1954:BRB

[Bor54a] Max Born. Book review: A History of the Theories of Aether and Electricity. The Modern Theories, 1900–1926, by Edmund Whittaker. The British Journal for the Philosophy of Science, 5(19):261– 263, November 1954. CODEN BJPIA5. ISSN 0007-0882 (print),

1464-3537 (electronic). URL http://bjps.oxfordjournals.org/
content/V/19/261.full.pdf+html;   http://www.jstor.org/

stable/685722. Special issue on The Age of the Universe.

Born:1954:BSP

[Bor54b] Max Born. Die begriffliche Situation in der Physik. (German) [The conceptual situation in physics]. Physikalische Blätter, 10(5):193– 201, May 1954. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19540100501/abstract.

Born:1954:IFR

[Bor54c] Max Born. On the interpretation of Freundlich’s red shift formula. Proceedings of the Physical Society, London, Section A, 67(2):193– 194, February 1, 1954. CODEN PPSAAM. ISSN 0370-1298 (print),

1747-3829 (electronic). URL http://iopscience.iop.org/0370-
1298/67/2/115.

Born:1955:AEL

[Bor55a] Max Born. Albert Einstein und das Lichtquantum. (German) [Albert Einstein and the quantum of light]. Naturwissenschaften, 42 (15):425–431, August 1, 1955. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1955:CDR

[Bor55b] Max Born. Continuity, determinism, and reality. Kongelige Danske Videnskabernes Selskab, Matematiskfysiske Meddelelser, 30(2):1– 26, ???? 1955. CODEN ???? ISSN ????

Born:1955:EME

[Bor55c] Max Born. ’Das entscheidende Motiv’, eine Erwiderung. (German) [’The main reason why’, a reply]. Merkur, ??(??):??, October 1955. CODEN ???? ISSN ????

Born:1955:SDQ

[Bor55d] Max Born. Die statistische Deutung der Quantenmechanik. (German) [The statistical interpretation of quantum mechanics]. In Les Prix Nobel en 1954 Stockholm, pages 79–90. ????, ????, 1955.

Born:1955:EWA

[Bor55e] Max Born. Entwicklung und Wesen des Atomzeitalters. (German) [Development and nature of the nuclear age]. Merkur, ??(??):??, August 1955. CODEN ???? ISSN ????

Born:1955:KMT

[Bor55f] Max Born. Ist die klassische Mechanik tatsächlich deterministisch?. (German) [Is classical mechanics really deterministic?]. Physikalische Blätter, 11(2):49–54, February 1955. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/phbl.19550110201/
abstract.

Born:1955:SIQ

[Bor55g] Max Born. Statistical interpretation of quantum mechanics. Science (New Series), 122(3172):675–679, October 14, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/pdfplus/1752079.pdf.

Born:1955:FDG

[Bor55h] Max Born. Zur Frage des Determinismus. (German) [Toward the question of determinism]. Physikalische Blätter, 11(7):314–315, July 1955. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19550110705/abstract.

Born:1956:EAE

[Bor56a] Max Born. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. Mathematische und Naturwissenschaftliche Unterricht, 9(3):97–105, ???? 1956. CODEN MNWUAL. ISSN 0025-5866.

Born:1956:ETP

[Bor56b] Max Born. Experiment and theory in physics. Dover Publications, Inc., New York, NY, USA, 1956. 43 pp. LCCN ????

Born:1956:PMG

[Bor56c] Max Born. Physics in My Generation. Pergamon, New York, NY, USA, 1956. viii + 232 pp. LCCN ????

Born:1956:PRG

[Bor56d] Max Born. Physik und Relativität. (German) [Physics and Relativity]. Naturwissenschaftliche Rundschau, 11(??):417–424, ???? 1956. CODEN NARSAC. ISSN 0028-1050.

Born:1957:EBN

[Bor57a] Max Born. Europaische Betrachtungen eines Naturforschers. (German) [European reflections of a naturalist]. Die Sammlung, 12(??): 593–604, ???? 1957. CODEN ???? ISSN ????

Born:1957:MA

[Bor57b] Max Born. Man and the atom. Bulletin of the Atomic Scientists, 13 (6):186–194, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bor63-27]. See note in [BR57].

Born:1957:MP

[Bor57c] Max Born. Max Planck 1858–1947. In Hermann Heimpel, editor, Die Grossen Deutschen: deutsche Biographie, page ?? Propylaen-Verlag bei Ullstein, Berlin, Germany, 1957.

Born:1957:MFB

[Bor57d] Max Born. [My former belief in the superiority of science]. Bulletin of the Atomic Scientists, 13(1):31, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1957:PIW

[Bor57e] Max Born. Physik im Wandel meiner Zeit. (German) [Physics in Flux in My Time], volume 111 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1957. vii + 252 pp. LCCN QC71.

Born:1958:CRP

[Bor58a] Max Born. The concept of reality in physics. Bulletin of the Atomic Scientists, 14(8):313–321, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1958:RPG

[Bor58b] Max Born. Der Realitätsbegriff in der Physik. (German) [The concept of reality in physics]. Die Sammlung, 346(??):13–??, ???? 1958. CODEN ???? ISSN ????

Born:1958:BRU

[Bor58c] Max Born. Ein Besuch bei den Raumfahrern und das Uhrenparadoxon. (German) [A visit with astronauts and the clock paradox]. Physikalische Blätter, 14(5):207–212, May 1958. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/phbl.19580140503/
abstract.

Born:1958:ES

[Bor58d] Max Born. Europe and science. Bulletin of the Atomic Scientists, 14(2):73–79, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1958:PIW

[Bor58e] Max Born. Physik im Wandel meiner Zeit. (German) [Physics in Flux in My Time], volume 111 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1958. vii + 252 pp. LCCN ????

Born:1958:RCS

[Bor58f] Max Born. Reader’s comment: Space flight for what? Bulletin of the Atomic Scientists, 14(6):234, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1958:SFS

[Bor58g] Max Born. Sir Francis Simon, F.R.S. Z. phys. Chem, 16(??):ix–xvii,

???? 1958. CODEN ???? ISSN ????

Born:1958:VKM

[Bor58h] Max Born. Vorhersagbarkeit in der klassischen Mechanik. (German) [Predictability in classical mechanics]. Zeitschrift für Physik, 153(3):372–388, June 1958. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.com/

article/10.1007/BF01329043.

Born:1959:QMM

[Bor59a] Max Born. Dans quelle mesure la mécanique classique peut-elle prédire les trajectoires?. (French) [To what extent can classical mechanics predict trajectories?.]. Journal de Physique et le Radium, 20(1):43–50, ???? 1959. CODEN JPRAAJ. ISSN 0368-3842.

Born:1959:EHM

[Bor59b] Max Born. Erinnerungen an Hermann Minkowski zur 50 Wiederkehr seines Todestages. (German) [Memories of Hermann Minkowski on the 50th anniversary of his death]. Naturwissenschaften, 46(17):501–505, September 1, 1959. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1959:EMV

[Bor59c] Max Born. Erinnerungen an Max v. Laues Entdeckung der Beugung von Röntgenstrahlen durch Kristalle. (German) [Memories of Max von Laue’s discovery of the diffraction of X-rays by crystals]. Zeitschrift für Kristallographie, 112(??):1–??, ???? 1959. CODEN ZEKRDZ. ISSN 0044-2968.

Born:1959:PIW

[Bor59d] Max Born. Physik im Wandel meiner Zeit. (German) [Physics in Flux in My Time], volume 111 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, third edition, 1959. vii + 252 pp. LCCN ????

Born:1959:VKM

[Bor59e] Max Born. Voraussagbarkeit in der klassischen Mechanik. (German) [Predictability in classical mechanics]. Physikalische Blätter, 15(8):342–349, August 1959. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19590150802/abstract.

Born:1960:PIG

[Bor60a] Max Born. Die Physik und die Ismen. (German) [Physics and isms]. Physikalische Blätter, 16(4):147–148, April 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19600160402/abstract.

Born:1960:LMI

[Bor60b] Max Born. L’universo in movimento. (Italian) [The restless universe]. G. Feltrinelli, Milano, Italy, 1960. 244 pp. LCCN ????

Born:1960:MA

[Bor60c] Max Born. Mechanics of the atom. F. Ungar Publishing Co., London, UK, 1960. xvi + 317 pp. LCCN ????

Born:1960:PP

[Bor60d] Max Born. Physics and politics. Bulletin of the Atomic Scientists, 16(6):194–200, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1960:PPG

[Bor60e] Max Born. Physik und Politik. (German) [Physics and politics]. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1960. ???? pp. LCCN ????

Born:1961:BSD

[Bor61a] Max Born. Bemerkungen zur statistischen Deutung der Quantenmechanik. (German) [Remarks on the statistical interpretation of quantum mechanics]. In W. Heisenberg und die Physik unserer Zeit. (German) [W. Heisenberg and the physics of our time], pages 103– 118. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961.

Born:1961:FNT

[Bor61b] Max Born. La fisica e il nostro tempo. (Italian) [Physics and Our Time]. Sansoni, Firenze, Italy, 1961. 401 + 16 pp. LCCN ????

Born:1961:MAL

[Bor61c] Max Born. More about Lorentz transformation equations. The British Journal for the Philosophy of Science, 12(46):150–151, August 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/
XII/46/150.full.pdf+html;   http://www.jstor.org/stable/
685494;   http://www.jstor.org/stable/pdfplus/685494.pdf.

Born:1962:PPU

[Bor62a] Max Born. Die Physik in der Problematik unseres Zeitalters. (German) [The physics of the problem of our age]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 17(??):??, ???? 1962. CODEN UNIVA8. ISSN 0041-9079.

Born:1962:ETR

[Bor62b] Max Born. Einstein’s Theory of Relativity. Dover Publications, Inc., New York, NY, USA, 1962. vii + 376 pp. LCCN QC6 .B66 1962. With the collaboration of Günther Leibfried and Walter Biem.

Born:1962:JF

[Bor62c] Max Born. Jacob Frenkel. Soviet Physics. Uspekhi, 5(??):194–??, ???? 1962. CODEN SOPUAP. ISSN 0038-5670.

Born:1962:PPa

[Bor62d] Max Born. Physics and politics. Oliver and Boyd, Edinburgh, UK; London, UK, 1962. ???? pp. LCCN ????

Born:1962:PPb

[Bor62e] Max Born. Physics and politics. Basic Books, New York, NY, USA, 1962. vii + 86 pp. LCCN QC71 .B6653.

Born:1962:TPW

[Bor62f] Max Born. Translation of a poem by Wilhelm Busch. Edinburgh University Journal, ??(??):??, Spring 1962. CODEN ???? ISSN ????

Born:1963:ALG

[Bor63a] Max Born. Antoon Lorentz. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 81, pages 607–611. LCCN ???? Reprinted from Nachr. Ges. Wiss. Göttingen 1928, 69–73.

Born:1963:AJW

[Bor63b] Max Born. Arnold Johannes Wilhelm Sommerfeld 1868–1951. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 89, pages 647–659. LCCN ???? Reprinted from Obit. Notices Roy. Soc. 8, 275–287 (1952).

Born:1963:AAGa

[Bor63c] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 1. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. xxiv + 718 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:AAGb

[Bor63d] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:BSD

[Bor63e] Max Born. Bemerkungen zur statistischen Deutung der Quantenmechanik. (German) [Comments on the statistical interpretation of quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 70, pages 454–468. LCCN ???? Reprinted from Werner Heisenberg und die Physik unserer Zeit. Vieweg, Braunschweig 1961, pp. 103–118.

Born:1963:BTS

[Bor63f] Max Born. Berichtigung: Die Theorie des starren Elektrons in der Kinematik des Relativitätsprinzips. (German) [Correction: Theory of the rigid electron in the kinematics of the Principle of Relativity].

In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 2, page 160. LCCN ???? See [Bor63n]. Reprinted from Ann. Physik (4), 30, 840 (1909).

Born:1963:BFG

[Bor63g] Max Born. Betrachtungen zur Farbenlehre. (German) [Considerations of color theory]. Naturwissenschaften, 50(??):29–39, ???? 1963. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1963:CDR

[Bor63h] Max Born. Continuity, determinism and reality. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 6, pages 196–219. LCCN ???? Reprinted from Dan. Mat. Fys. Medd. 80, No. 2 (1955), 26 S..

Born:1963:CSC

[Bor63i] Max Born. Corrigenda: On the Stability of Crystal Lattices IX. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 34, pages 581–582. LCCN ???? See [Bor63-32]. Reprinted from Proc. Camb. Phil. Soc. 40, 262–263 (1944).

Born:1963:AQG

[Bor63j] Max Born. Das Adiabatenprinzip in der Quantenmechanik. (German) [Adiabatic principle in quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 58, pages 258–283. LCCN ???? Reprinted from Z. Physik 40, 167–192 (1926).

Born:1963:IES

[Bor63k] Max Born. Der Impuls-Energie-Satz in der Elektrodynamik von Gustav Mie. (German) [The pulse-energy theorem in Gustav Mie’s electrodynamics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 71, pages 470–483. LCCN ???? Reprinted from Nachr. Ges. Wiss. Göttingen 1914, 23–37.

Born:1963:EHG

[Bor63l] Max Born. Die Elektronenaffinität der Halogenatome. (German) [The electron affinity of halogen atoms]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 19, pages 375– 382. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 21, 679–685 (1919).

Born:1963:SDQ

[Bor63m] Max Born. Die statistische Deutung der Quantenmechanik. (German) [The statistical interpretation of quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 68, pages 430–441. LCCN ???? Reprinted from Nobelvortrag, gehalten am 11. Dezember 1954, Les Prix Nobel en 1954, Stockholm 1955, pp. 79–90.

Born:1963:TSE

[Bor63n] Max Born. Die Theorie des starren Elektrons in der Kinematik des Relativitätsprinzips. (German) [The theory of the rigid electron in the kinematics of the Principle of Relativity]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 2, pages 105–159. LCCN ???? See correction [Bor63f]. Reprinted from Ann. Physik (4), 30, 1–56 (1909).

Born:1963:UEA

[Bor63o] Max Born. Die ultraroten Eigenfrequenzen der Alkalihalogenidkristalle. (German) [The infra-red natural frequencies of alkali halide crystals]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 30, pages 527–528. LCCN ???? Reprinted from Z. Physik 76, 559–560 (1932).

Born:1963:DAB

[Bor63p] Max Born. Dr. Arnold Berliner. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 84, page 620. LCCN ???? Reprinted from Nature 150, 284 (1942).

Born:1963:BEG

[Bor63q] Max Born. Eine Bemerkung über den Elektronenradius. (German) [A note on the electron radius]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 72, pages 484–485. LCCN ???? Reprinted from Naturwiss. 20, 269 (1932).

Born:1963:DMF

[Bor63r] Max Born. Eine direkte Messung der freien Weglänge neutraler Atome. (German) [A direct measurement of the free path of neutral atoms]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 43, pages 694–697. LCCN ???? Reprinted from Physik. Zschr. 21, 578–581 (1920).

Born:1963:TAG

[Bor63s] Max Born. Eine thermochemische Anwendung der Gittertheorie. (German) [A thermochemical application of lattice theory]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 18, pages 363–374. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 21, 13–24 (1919).

Born:1963:ECD

[Bor63t] Max Born. Elastic constants of diamond. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 38, pages 634– 635. LCCN ???? Reprinted from Nature 157, 582 (1946).

Born:1963:EAE

[Bor63u] Max Born. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 90, pages 660–668. LCCN ???? Reprinted from Der Mathematische und Naturwissenschaftliche Unterricht IX 97–105 (1956).

Born:1963:EHM

[Bor63v] Max Born. Erinnerungen an Hermann Minkowski zur 50. Wiederkehr seines Todestages. (German) [Memories of Hermann Minkowski on the 50th anniversary of his death]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 92, pages 678–690. LCCN ???? Reprinted from Naturwiss. 46, 501–505 (1959).

Born:1963:ES

[Bor63w] Max Born. Erwin Schrödinger. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 93, pages 691–694. LCCN ???? Reprinted from Physik. Bl. 17, 85–87 (1961).

Born:1963:HPG

[Bor63x] Max Born. Hilbert und die Physik. (German) [Hilbert and physics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 78, pages 584–598. LCCN ???? Reprinted from Naturwiss. 10, 88–93 (1922).

Born:1963:KEK

[Bor63y] Max Born. Kopplung der Elektronen- und Kernbewegung in Molekeln und Kristallen. (German) [Coupling of electron and nuclear motion in molecules and crystals]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 40, pages 654– 655. LCCN ???? Reprinted from Nachr. Akad. Wiss. Göttingen, 1951, Nr. 6.

Born:1963:KBT

[Bor63z] Max Born. Kritische Betrachtungen zur traditionellen Darstellung der Thermodynamik. (German) [Critical considerations on the traditional representation of thermodynamics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 3, pages 161–176. LCCN ???? Reprinted from Physik. Zschr. 22, 218–224, 249–254, 282–286 (1921).

Born:1963:MAa

[Bor63-27] Max Born. Man and the atom. In Grodzins and Rabinowitch [GR63], pages 590–601. LCCN D842 .B78. Reprint of [Bor57b].

Born:1963:MKE

[Bor63-28] Max Born. Max Karl Ernst Ludwig Planck 1858–1947. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 88, pages 626–646. LCCN ???? Reprinted from Obit. Notices Roy. Soc. 6, 161–181 (1948).

Born:1963:QTP

[Bor63-29] Max Born. On the quantum theory of pyroelectricity. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 37, pages 627–633. LCCN ???? Reprinted from Rev. Mod. Physics 17, 245–251 (1945).

Born:1963:QTE

[Bor63-30] Max Born. On the quantum theory of the electromagnetic field. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 73, pages 486–513. LCCN ???? Reprinted from Proc. Roy. Soc. A 143, 410–437 (1934).

Born:1963:SCLa

[Bor63-31] Max Born. On the stability of crystal lattices I. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 33, pages 550–562. LCCN ???? Reprinted from Proc. Camb. Phil. Soc. 36, 160–172 (1940).

Born:1963:SCLb

[Bor63-32] Max Born. On the stability of crystal lattices IX. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 34, pages 563–580. LCCN ???? See [Bor63i]. Reprinted from Proc. Camb. Phil. Soc. 38, 82–99 (1942).

Born:1963:POT

[Bor63-33] Max Born. Prof. Otto Toeplitz. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 82, pages 612–613. LCCN ???? Reprinted from Nature 145, 617 (1940).

Born:1963:PHR

[Bor63-34] Max Born. Professor Heinrich Rausch von Traubenberg. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 86, pages 623–624. LCCN ???? Reprinted from Nature 157, 328 (1946).

Born:1963:PVM

[Bor63-35] Max Born. Professor V. M. Goldschmidt, For.Mem.R.S. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 87, page 625. LCCN ???? Reprinted from Nature 159, 701 (1947).

Born:1963:QSVb

[Bor63-36] Max Born. Quantenmechanik der Stoßvorgänge. (German) [quantum mechanics of collision processes]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 57, pages 233– 257. LCCN ???? Reprinted from Z. Physik / 38, 803–827 (1926).

Born:1963:QSG

[Bor63-37] Max Born. Quantenmechanik und Statistik. (German) [Quantum mechanics and statistics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 60, pages 299–309. LCCN ???? Reprinted from Naturwiss. 15, 238–242 (1927).

Born:1963:RNP

[Bor63-38] Max Born. Reciprocity and the number 137, Part I. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 77, pages 579–582. LCCN ???? Reprinted from Proc. Roy. Soc. Edinb. 59, 219–223 (1939).

Born:1963:SFS

[Bor63-39] Max Born. Sir Francis Simon, F.R.S. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 91, pages 669– 677. LCCN ???? Reprinted from Z. physik. Chemie N.F. 16, Simon-Gedenkheft, pp. ix–xvii (1958).

Born:1963:SJJ

[Bor63-40] Max Born. Sir J. J. Thomson, O.M., F.R.S. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 83, pages 614–619. LCCN ???? Reprinted from Proc. Physic. Soc. Lond. 53, 305–310 (1941).

Born:1963:SBS

[Bor63-41] Max Born. Sommerfeld als Begründer einer Schule. (German) [Sommerfeld as the founder of a school]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 80, pages 604–606. LCCN ???? Reprinted from Naturwiss. 16, 1035–1036 (1928).

Born:1963:STR

[Bor63-42] Max Born. Special Theory of Relativity. Nature, 197(4874):1287, March 30, 1963. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v197/n4874/pdf/1971287a0.pdf.

Born:1963:SEL

[Bor63-43] Max Born. Stabilität der elastischen Linie in Ebene und Raum. (German) [Stability of the elastic line in the plane and space]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 1, pages 1–104. LCCN ???? Reprinted from Preisschrift und Dissertation, Göttingen, Dieterichsche Universitäts Buchdruckerei Göttingen, 1906.

Born:1963:SMC

[Bor63-44] Max Born. The statistical mechanics of condensing systems. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 45, pages 699–709. LCCN ???? Reprinted from Physica 4, 1034–1044 (1937).

Born:1963:SUQ

[Bor63-45] Max Born. A suggestion for unifying quantum theory and relativity. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 76, pages 560–578. LCCN ???? Reprinted from Proc. Roy. Soc. A 165, 291–303 (1938).

Born:1963:TIR

[Bor63-46] Max Born. Theoretical investigations on the relation between crystal dynamics and X-ray scattering. In Ausgewählte Abhandlungen.

(German) [Selected works] [Bor63c], chapter 36, pages 587–626. LCCN ???? Reprinted from Rep. Progr. Physics 9, 294–333 (1943).

Born:1963:TCL

[Bor63-47] Max Born. The thermodynamics of crystal lattices. I. Discussion of the methods of calculation. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 35, pages 583–586. LCCN ???? Reprinted from Proc. Camb. Phil. Soc. 39, 100–103 (1943).

Born:1963:TCM

[Bor63-48] Max Born. Thermodynamics of crystals and melting. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 32, pages 537–549. LCCN ???? Reprinted from J. Chem. Physics 7, 591–603 (1939).

Born:1963:BEI

[Bor63-49] Max Born. Über die Beweglichkeit der elektrolytischen Ionen. (German) [On mobility of electrolytic ions]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 42, pages 665– 693. LCCN ???? Reprinted from Z. Physik 1, 221–249 (1920).

Born:1963:MBZ

[Bor63-50] Max Born. Über die Maxwellsche Beziehung zwischen Brechungsindex und Dielektrizitätskonstante und über eine Methode zur Bestimmung der Ionenladung von Kristallen. (German) [On the Maxwellian relationship between refractive index and dielectric constant and on a method for determining the ionic charge of crystals]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 13, pages 307–316. LCCN ???? Reprinted from S. B. Preuß. Akad. Wiss. Berlin 1918, 604–613.

Born:1963:NOA

[Bor63-51] Max Born. Über die natürliche optische Aktivität von Flüssigkeiten und Gasen. (German) [On the natural optical activity of liquids and gases]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 41, pages 657–664. LCCN ???? Reprinted from Physik. Zschr. 16, 251–258 (1915).

Born:1963:OAK

[Bor63-52] Max Born. Über die optische Aktivität der Kristalle. (German) [On the optical activity of crystals]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 12, pages 294– 306. LCCN ???? Reprinted from Elster-Geitel-Festschrift, Vieweg, Braunschweig 1915, S. 391–403.

Born:1963:UEZ

[Bor63-53] Max Born. Über die ultraroten Eigenschwingungen zweiatomiger Kristalle. (German) [On the infra-red natural frequencies of diatomic crystals]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 14, pages 317–326. LCCN ???? Reprinted from Physik. Zschr. 19, 539–548 (1918).

Born:1963:DMN

[Bor63-54] Max Born. Über einen direkten mechanischen Nachweis des Dipolcharakters von Flüssigkeitsmolekeln. (German) [On a direct mechanical proof of the dipole character of liquid molecules]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 44, page 698. LCCN ???? Reprinted from Verh. Dtsch. Physik. Ges. 2, 53 (1921).

Born:1963:EGG

[Bor63-55] Max Born. Über elektrostatische Gitterpotentiale. (German) [On electrostatic lattice potentials]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 26, pages 434–450. LCCN ???? Reprinted from Z.Physik 7, 124–140 (1921).

Born:1963:QGQa

[Bor63-56] Max Born. Über Quantenmechanik. (German) [On quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 50, pages 61–77. LCCN ???? Reprinted from Z. Physik 26, 379–395 (1924).

Born:1963:VWS

[Bor63-57] Max Born. Verzeichnis der wissenschaftlichen Schriften von Max Born. (German) [List of scientific writings of Max Born]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], pages 695–?? LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:WBN

[Bor63-58] Max Born. Was bleibt noch zu hoffen?. (German) [What’s left to hope for?]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 18(4):337–346, ???? 1963. CODEN UNIVA8. ISSN 0041-9079.

Born:1963:QSVa

[Bor63-59] Max Born. Zur Quantenmechanik der Stoßvorgänge. (German). [Toward a quantum mechanics of collision processes]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 56, pages 228–232. LCCN ???? Reprinted from Z. Physik 37, 863–867 (1926).

Born:1963:QCK

[Bor63-60] Max Born. Zur Quantentheorie der chemischen Kräfte. (German) [Toward the quantum theory of chemical forces]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 64, pages 370–781. LCCN ???? Reprinted from Z. Physik 64, 729–740 (1930).

Born:1963:RDG

[Bor63-61] Max Born. Zur Raumgittertheorie des Diamanten. (German) [Toward the space lattice theory of diamond]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 11, pages 256–293. LCCN ???? Reprinted from Ann. Physik (4) 44, 605–642 (1914).

Born:1963:TKGb

[Bor63-62] Max Born. Zur Theorie des Kernzerfalls. (German) [Toward the theory of nuclear disintegration]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 63, pages 354–369. LCCN ???? Reprinted from Z. Physik 58, 306–321 (1929).

Born:1963:TKGa

[Bor63-63] Max Born. Zur Thermodynamik der Kristallgitter. (German) [Toward a thermodynamics of crystal lattices]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 27, pages 451–482. LCCN ???? Reprinted from Z. Physik 7, 217–248 (1921).

Born:1963:WSV

[Bor63-64] Max Born. Zur Wellenmechanik der Stoßvorgänge. (German) [Toward a wave mechanics of collision processes]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 59, pages 284–298. LCCN ???? Reprinted from Nachr. Ges. Wiss. Göttingen 1926, 146–160.

Born:1964:MA

[Bor64a] Max Born. Der Mensch und das Atom. (German) [Man and the atom]. In Die Friedensrundschau (Hamburg). (German) [The peace review (Hamburg)], pages 16–23. ????, ????, 1964.

Born:1964:REG

[Bor64b] Max Born. Die Relativitaätstheorie Einsteins. (German) [Einstein’s Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1964. xi + 328 pp. LCCN QC6 .B65 1964. With the collaboration of Walter Biem.

Born:1964:MMB

[Bor64c] Max Born. Message from Max Born. Reviews of Modern Physics, 36(2):509, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://rmp.aps.org/

abstract/RMP/v36/i2/p509_1. Open letter to Robert Oppen-

heimer.

Born:1964:RMW

[Bor64d] Max Born. Reminiscences of my work on the dynamics of crystal lattices. In Proceedings of the International Conference Lattice Dynamics at Copenhagen, page ?? Pergamon, New York, NY, USA, 1964.

Born:1964:SW

[Bor64e] Max Born. Symbol und Wirklichkeit. (German) [Symbol and reality]. Physik Journal, 20(12):554–555, December 1964. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Born:1964:SWG

[Bor64f] Max Born. Symbol und Wirklichkeit. (German) [Symbol and reality]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 19 (8):817–834, ???? 1964. CODEN UNIVA8. ISSN 0041-9079.

Born:1964:WLH

[Bor64g] Max Born. What is left for hope for? Bulletin of the Atomic Scientists, 20(4):2–5, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1965:ETR

[Bor65a] Max Born. Einstein’s theory of relativity. Dover Publications, Inc., New York, NY, USA, revised edition, 1965. ISBN 0-486-60769-0. vii + 376 pp. LCCN QC6 .B653 1965. URL http://www.loc.

gov/catdir/description/dover032/65001214.html;   http:/
/www.loc.gov/catdir/enhancements/fy1318/65001214-t.html.

Prepared with the collaboration of Günther Leibfried and Walter Biem.

Born:1965:EEG

[Bor65b] Max Born. Erinnerungen an Einstein. (German) [Memories of Einstein]. Physikalische Blätter, 21(7):297–306, July 1965. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19650210701/abstract.

Born:1965:FAP

[Bor65c] Max Born. Física Atómica. (Portuguese) [Atomic Physics]. Fundação Calouste Gulbenkian, Lisboa, Portugal, 1965. 513 pp. LCCN ????

Born:1965:OLE

[Bor65d] Max Born. Optik: Ein Lehrbuch der elektromagnetischen Lichttheorie. (German) [Optics: a textbook of the electromagnetic theory of light]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1965. viii + 591 pp. LCCN ????

Born:1965:RTE

[Bor65e] Max Born. Reciprocity theory of elementary particles. Progress of Theoretical Physics (Osaka), ??(??):31–55, ???? 1965. CODEN ???? ISSN ???? Supplementary Number.

Born:1965:RMBa

[Bor65f] Max Born. Recollections of Max Born. I. How I became a physicist. Bulletin of the Atomic Scientists, 21(7):3–6, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1965:RMBb

[Bor65g] Max Born. Recollections of Max Born. II. What i did as a physicist. Bulletin of the Atomic Scientists, 21(8):9–13, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1965:RMBc

[Bor65h] Max Born. Recollections of Max Born. III. Reflections. Bulletin of the Atomic Scientists, 21(9):3–6, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wil67].

Born:1965:SWV

[Bor65i] Max Born. Symbol und Wirklichkeit I: Ein Versuch, auf naturwissenschaftliche Weise zu philosophieren — nicht eine Philosophie der Naturwissenschaften. (German) [Symbol and reality I: An attempt to philosophize in a scientific manner not a philosophy of science]. Physik Journal, 21(2):53–63, February 1965.

Born:1965:SWI

[Bor65j] Max Born. Symbol und Wirklichkeit. II. (German) [Symbol and reality. II]. Physikalische Blätter, 21(3):106–108, March 1965. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19650210302/abstract.

Born:1965:VNG

[Bor65k] Max Born. Von der Verantwortung des Naturwissenschaftlers. (German) [On the responsibility of scientists]. Nymphenburger Verlagshandlung GmbH, München, West Germany, 1965. ???? pp. LCCN ????

Born:1966:SR

[Bor66a] M. Born. Symbol and reality. Dialectica: International Review of Philosophy of Knowledge, 20(2):143–157, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Born:1966:BES

[Bor66b] Max Born. Blessings and evils of space travel. Bulletin of the Atomic Scientists, 22(8):12–14, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translated by Charles W. Scherr from the German original, published in Universitas and in Neue Deutsche Hefte, 106, 113 (1965), as well as in the author’s 1965 book of essays, Von der Verantwortung Naturwissenschaftlers.

Born:1966:HVG

[Bor66c] Max Born. Hoffnung auf die Vernunft. (German) [Hope for the reason]. Europa, ??(??):7–8, ???? 1966. CODEN ???? ISSN ????

Born:1966:MRB

[Bor66d] Max Born. The Moon race and beyond. Discovery, 5(??):27–??, ???? 1966. CODEN ???? ISSN ????

Born:1966:PIW

[Bor66e] Max Born. Physik im Wandel meiner Zeit. (German) [Physics in Flux in My Time], volume 111 of Die Wissenschaft. Friedrich Vieweg und Sohn, Braunschweig, Germany, fourth expanded edition, 1966. 299 pp. LCCN ????

Born:1966:OS

[Bor66f] Max Born. Über Objektivität und Subjektivität. (German) [On objectivity and subjectivity]. In Dialog des Abendlandes — Physik und Philosophie. (German) [Dialogue of the West — physics and philosophy], page ?? ????, München, Germany, 1966.

Born:1966:SUW

[Bor66g] Max Born. Vom Segen und Unsegen der Weltraumfahrt. (German) [The blessing and curse of space travel]. Deutsches Adelsblatt, ?? (??):165–166, ???? 1966. CODEN ???? ISSN ????

Born:1967:MA

[Bor67a] Max Born. The mechanics of the atom. F. Ungar Pub. Co., New York, NY, USA, 1967. 317 pp. LCCN ???? Translated from the German by J. W. Fisher and revised by D. R. Hartree.

Born:1967:P

[Bor67b] Max Born. Prolog. In Walter R. Fuchs, editor, El libro de la física moderna. (Spanish) [The book of modern physics], page 363. Omega, Barcelona, Spain, 1967. LCCN ????

Born:1967:QMa

[Bor67c] Max Born. Quantum mechanics. In van der Waerden [vdW67], pages 181–198. LCCN QC174.12 S655. English translation of [Bor24f].

Born:1968:BFR

[Bor68a] Max Born. Book and film reviews: An extraordinary achievement: My life and my views. The Physics Teacher, 6(6):310–312, September 1968. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928

(electronic). URL http://link.aip.org/link/?PTE/6/310/1;
http://tpt.aapt.org/resource/1/phteah/v6/i6/p310_s1.

Born:1968:CCM

[Bor68b] Max Born. Current comments: From Max Born. Bulletin of the Atomic Scientists, 24(2):29, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ure67, Bla67].

Born:1968:EGP

[Bor68c] Max Born. Erinnerungen und Gedanken eines Physikers. (German) [Memories and thoughts of a physicist]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 23(??):249–276, ???? 1968. CODEN UNIVA8. ISSN 0041-9079.

Born:1968:MLV

[Bor68d] Max Born. My life and views. Charles Scribner’s Sons, New York, NY, USA, 1968. ???? pp. LCCN ????

Born:1969:EGG

[Bor69a] Max Born. Aus ,,Erinnerungen und Gedanken”. (German) [From “Memories and thoughts”]. Physik Journal, 25(7):289–295, July 1969. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Born:1969:BEB

[Bor69b] Max Born. Born–Einstein Briefwechsel. (German) [The Born– Einstein Letters]. Nymphenburger Verlagshandlung GmbH, München, West Germany, 1969. ???? pp. LCCN ????

Born:1969:REG

[Bor69c] Max Born. Die Relativitaätstheorie Einsteins. (German) [Einstein’s Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 1969. xii + 328 pp. LCCN QC6 .B65 1969. With the collaboration of Walter Biem.

Born:1969:FIS

[Bor69d] Max Born. El futuro inmediato. (Spanish) [The near future]. Plaza & Janes, Barcelona, Spain, 1969. 349 pp. LCCN ????

Born:1969:ETP

[Bor69e] Max Born. Experiment und Theorie in der Physik. (German) [Experiment and Theory in Physics]. Mosbach, ????, 1969. ???? pp. LCCN ????

Born:1969:FCG

[Bor69f] Max Born. Fizica în concepţia generaţiei mele. (Romanian) [Physics in Flux in My Time]. Editura Ştiinţifică, Bucureşti, Romania, 1969. 379 pp. LCCN ????

Born:1969:PMG

[Bor69g] Max Born. Physics in My Generation. Heidelberg science library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1969. vii + 172 pp. LCCN QC71 .B66 1969.

Born:1970:PMGa

[Bor70a] Max Born. Physics in My Generation. Longmans, Harlow, UK, 1970. vi + 172 pp. LCCN ????

Born:1970:PMGb

[Bor70b] Max Born. Physics in My Generation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1970. vi + 172 pp. LCCN ????

Born:1970:PAD

[Bor70c] Max Born. Problems of Atomic Dynamics. MIT Press paperbacks in the history of science and technology. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-52019-2. xiv + 200 pp. LCCN QC173 .B7363 1970.

Born:1971:FAI

[Bor71] Max Born. Fisica atomica. (Italian) [Atomic physics]. Boringhieri, Torino, Italy, seventh edition, 1971. 529 pp. LCCN ????

Born:1972:OLE

[Bor72] Max Born. Optik: Ein Lehrbuch der elektromagnetischen Lichttheorie. (German) [Optics: a textbook of the electromagnetic theory of light]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1972. ISBN 0-387-05954-7 (New York), 3-540-05954-7 (Berlin). vii + 591 pp. LCCN QC357.2 .B67 1972.

Born:1975:MLE

[Bor75] Max Born. Mein Leben. Die Erinnerungen des Nobelpreisträgers. (German) [My Life. Recollections of a Nobel Laureate]. Nymphenburger Verlagshandlung GmbH, München, West Germany, 1975.

400 pp. Translation to German by Helmut Degner with the collaboration of Ilse Krewinkel qnd Werner Rau.

Born:1977:QMC

[Bor77] M. Born. Quantum mechanics of collision processes. Uspekhi Fizicheskikh Nauk, 122(4):632–651, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Born:1978:MLR

[Bor78] Max Born. My life: recollections of a Nobel laureate. Scribner, New York, NY, USA, 1978. ISBN 0-684-15662-8. xi + 308 pp. LCCN QC16.B643 A3213 1978. US$17.50.

Born:1979:MBA

[Bor79] G. V. R. Born. Max Born: Another impression. Science (New Series), 206(4419):636, November 9, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/pdfplus/1749028.pdf;   http://www.
sciencemag.org/content/206/4419/636.1.full.pdf.

Born:1981:POT

[Bor81] Max Born. Professor Otto Toeplitz. Integral Equations and Operator Theory, 4(2):278–280, June 1981. CODEN ???? ISSN 0378-620X (print), 1420-8989 (electronic). Reprint of [Bor40f].

Born:1983:PIW

[Bor83] Max Born. Physik im Wandel meiner Zeit. (German) [Physics in Flux in My Time], volume 9 of Facetten der Physik. Friedrich Vieweg und Sohn, Braunschweig, Germany, fourth expanded edition, 1983. ISBN 3-528-08539-8. xxv + 299 pp. LCCN QC71 .B66x 1983. Reprint of [Bor66e].

Born:2001:YA

[Bor01a] Max Born. 100 and 50 years ago. Nature, 411(6835):249, May 17, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v411/
n6835/full/411249a0.html.

Born:2001:REG

[Bor01b] Max Born. Die Relativitätstheorie Einsteins. (German) [Einstein’s Theory of Relativity]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., sixth edition, 2001. ISBN 3-540-67904-9. xvii + 470 pp. LCCN QC6 .B74 2001. With comments and expansion by Jürgen Ehlers and Markus Pössel.

Born:2002:WRF

[Bor02] G. V. R. Born. The wide-ranging family history of Max Born. Notes and Records of the Royal Society of London, 56(2):219–262, May 22, 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-

0178 (electronic). URL http://www.jstor.org/stable/pdfplus/

3557669.pdf. See errata [Ano02].

Born:2007:MEL

[Bor07] Max Born. The Momentum–Energy Law in the electrodynamics of Gustav Mie. In Renn et al. [RSS^{+}07], pages 745–756. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 no. 250 v. 4; QC173.6 .G469

2007. URL https://link.springer.com/chapter/10.1007/

978-1-4020-4000-9_39. Originally published as “Der Impuls– Energie-Satz in der Elektrodynamik von Gustav Mie”, Nachrichten von der Königlichen Gesellschaft der Wissenschaften zu Göttingen 1, 23–36 (1914).

Bowley:1969:LTM

[Bow69] W. W. Bowley. Legendre transforms, Maxwell’s relations, and the Born diagram in fluid dynamics. American Journal of Physics, 37 (10):1066–??, October 1969. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bowman:2008:BBO

[Bow08] Joel M. Bowman. Beyond Born–Oppenheimer. Science, 319(5859): 40–41, January 4, 2008. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/319/5859/40.full.pdf.

Born:1922:QGM

[BP22] Max Born and W. Pauli, Jr. Über die Quantelung gestörter mechanischer Systeme. (German) [On the quantization of constrained mechanical systems]. Zeitschrift für Physik, 10(1):137–158, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/10.
1007/BF01332555;   http://www.springerlink.com/content/
l3686r1744052961/fulltext.pdf.

Born:1942:VTV

[BP42] Max Born and H. Ll. D. Pugh. Vibrations of a thin vertical cantilever caused by damped harmonic disturbance of the ground. J. Instn. Civil Engrs, 18(??):279–293, ???? 1942. CODEN ???? ISSN ????

Born:1944:QMFa

[BP44a] Max Born and H. W. Peng. Quantum mechanics of fields. I. Pure fields. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and physical sciences, 62(1):40–57, ???? 1944. CO-DEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Born:1944:QMFb

[BP44b] Max Born and H. W. Peng. Quantum mechanics of fields. II. Statistics of pure fields. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and physical sciences, 62(1):92–102, ???? 1944. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Born:1944:SMF

[BP44c] Max Born and H. W. Peng. Statistical mechanics of fields and the ’Apeiron’. Nature, 153(3875):164–165, February 5, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v153/
n3875/pdf/153164a0.pdf.

Born:1946:QMF

[BP46] Max Born and H. W. Peng. Quantum mechanics of fields. III. Electromagnetic field and electron field in interaction. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and physical sciences, 62(2):127–137, January 1946. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Born:1963:QGM

[BP63a] Max Born and W. Pauli, Jr. Über die Quantelung gestörter mechanischer Systeme. (German) [On the quantization of constrained mechanical systems]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 47, pages 1–22. LCCN ???? Reprinted from Z. Physik 10, 137–158 (1922).

Born:1963:VTV

[BP63b] Max Born and H. Ll. D. Pugh. Vibration of a thin vertical cantilever caused by damped harmonie disturbance of the ground. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 4, pages 177–191. LCCN ???? Reprinted from J. Instn. Civil Engrs. 18, 279–293 (1942).

Born:1931:AQG

[BR31] Max Born and Georg Rumer. Ansätze zur Quantenelektrodynamik. (German) [Approaches to quantum electrodynamics]. Zeitschrift für Physik, 69(3–4):141–152, May 1, 1931. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01798119.

Born:1949:MMP

[BR49] Max Born and Antonio E. Rodriguez. Meson masses and the Principle of Reciprocity. Nature, 163(4139):320–321, February 26, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v163/
n4139/pdf/163320b0.pdf.

Born:1957:LEW

[BR57] Max Born and Eugene I. Rabinowitch. Letter to the Editor: We’re sorry. Bulletin of the Atomic Scientists, 13(7):273, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bolton:1995:CLS

[BR95] H. C. Bolton and Alan Roberts. On the comparison of literary and scientific styles: The letters and articles of Max Born, F.R.S. Notes and Records of the Royal Society of London, 49(2):295–302, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/532016.
pdf.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Born:1971:FAP

[BRB71] Max Born, J. M. Radcliff, and R. J. Blin Stoyle. Física Atómica. (Portuguese) [Atomic Physics]. Fundação Calouste Gulbenkian, Lisboa, Portugal, third edition, 1971. xv + 357 pp. LCCN ???? Translation to Portuguese by Egídio Namorado.

Bretscher:1970:BVP

[Bre70] Egon Bretscher. Born’s view of physics. Nature, 227(5255): 311, July 18, 1970. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v227/n5255/pdf/227311b0.pdf.

Bredig:1973:BRB

[Bre73] Max A. Bredig. Book review: The Born–Einstein Letters. Science, 180(4091):1118, June 15, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/180/4091/1118.3.full.pdf.

Brown:1968:CWS

[Bro68] Frederic Joseph Brown. Chemical warfare: a study in restraints. Princeton University Press, Princeton, NJ, USA, 1968. xix + 355 pp. LCCN UG447 .B73.

Bromberg:1972:BRT

[Bro72] Joan Bromberg. Book review: Twentieth Century The Born– Einstein Letters. Correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Trans. by Irene Born. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. London: Macmillan, 1971. Pp. xi + 240. £3.85. British Journal for the History of Science, 6 (2):222–223, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4025309.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:SSR

[Bro73b] Jacob Bronowski. The sub-structure of reality: Max Born. In The Ascent of Man [Bro73a], pages 360–364. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Brown:2006:CWS

[Bro06] Frederic Joseph Brown. Chemical warfare: a study in restraints. Transaction Publishers, New Brunswick, NJ, USA, 2006. ISBN 1-4128-0495-7. xxxiii + 355 pp. LCCN UG447 .B73 2006.

Born:1916:AFI

[BS16] Max Born and F. Stumpf. Über anisotrope Flüssigkeiten. II. (German) [On anisotropic fluids. II]. S. B. preuss. Akad. Wiss. Berlin, ??(??):1043–1060, ???? 1916. CODEN ???? ISSN ????

Born:1919:OKI

[BS19] Max Born and O. Stern. Über die Oberflächenenergie der Kristalle und ihren Einfluß auf die Kristallgestalt. (German) [On the surface energy of crystals and their influence on crystal shape]. Sitzungsberichte der Preußischen Akademie der Wissenschaften: Physikalisch-Mathematische Klasse, 48(??):901–913, November 27, 1919. CODEN SPWPAI. ISSN ????

Born:1935:AFC

[BS35] Max Born and Erwin Schrödinger. The absolute field constant in the new field theory. Nature, 135(3409):342, March 2, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v135/
n3409/pdf/135342a0.pdf.

Born:1941:ETV

[BS41] Max Born and Kathleen Sarginson. The effect of thermal vibrations on the scattering of X-rays. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 179(976):69–93, August 29, 1941. CODEN PRLAAZ. ISSN 0080-4630. URL http:

//rspa.royalsocietypublishing.org/content/179/976/69.

Born:1945:PZ

[BS45] Max Born and R. Schlapp. Pieter Zeeman. In R. Soc. Edinb. Year Book, pages 25–27. Royal Society of Edinburgh, Edinburgh, Scotland, ???? 1945.

Born:1948:TSG

[BS48] Max Born and Kaj-Sia Seng. Zur Theorie der Supraleitfähigkeit. (German) [Toward a theory of superconductivity]. Dokl. Akad. Nauk SSSR, n. Ser., 62(??):313–318, ???? 1948. CODEN ???? ISSN ????

Born:1950:DPF

[BS50] Max Born and Alexander Schönberg. Demonstrability of the photochemical formation of biradicals by magnetic methods. Nature, 166(4216):307, August 19, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v166/n4216/pdf/166307a0.pdf.

Born:1963:PZH

[BS63a] Max Born and R. Schlapp. Pieter Zeemann, Hon. F.R.S.E. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 85, pages 621–622. LCCN ???? Reprinted from Roy. Soc. Edinb. Year Book 1943–44 25–27.

Born:1963:OKI

[BS63b] Max Born and O. Stern. Über die Oberflächenenergie der Kristalle und ihren Einfluß auf die Kristallgestalt. (German) [On the surface energy of crystals and their influence on crystal shape]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 20, pages 382–394. LCCN ???? Reprinted from S. B. Preuß. Akad. Wiss. Berlin 1919, 901–913.

Bishop:1985:BBO

[BS85] D. M. Bishop and S. A. Solunac. Breakdown of the Born– Oppenheimer approximation in the calculation of electric hyperpolarizabilities. Physical Review Letters, 55(19):1986–1988, November 1985. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://adsabs.harvard.edu/abs/
1985PhRvL..55.1986B.

Born:1934:NSF

[BT34] Max Born and J. H. C. Thompson. A note on the spectrum of the frequencies of a polar crystal lattice. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 147(862):594–599, December 1, 1934. CODEN PRLAAZ.

ISSN 0080-4630. URL http://rspa.royalsocietypublishing.
org/content/147/862/594.

Busch:1965:KMG

[Bus65] Wilhelm Busch. Klecksel der Maler. (German) [Klecksel the painter]. Frederick Ungar, New York, 1965. 76 pp. LCCN ???? Translation to English by Max Born from the German original.

Bacciagaluppi:2009:QTC

[BV09] Guido Bacciagaluppi and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay conference.

Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9 (hardcover). LCCN QC173.96 .B33 2009. URL http://

www.loc.gov/catdir/toc/ecip0817/2008019585.html.

Born:1913:VEP

[BvK13a] Max Born and Th. v. Kármán. Über die Verteilung der Eigenschwingungen von Punktgittern. (German) [On the distribution of the natural vibrations of point grids]. Physikalische Zeitschrift, 14(2):65–71, January 15, 1913. CODEN PHZTAO. ISSN 0369-

982X. URL http://babel.hathitrust.org/cgi/pt?id=mdp.
39015021268936%3Bseq=101%3Bview=1up.

Born:1913:TSW

[BvK13b] Max Born and Th. v. Kármán. Zur Theorie der specifischen Wärme. (German) [Toward the theory of specific heat]. Physikalische Zeitschrift, 14(1):15–19, January 1, 1913. CODEN PHZTAO.

ISSN 0369-982X. URL http://babel.hathitrust.org/cgi/pt?
id=mdp.39015021268936%3Bseq=51%3Bview=1up.

Born:1963:VEP

[BvK63a] Max Born and Th. v. Kármán. Über die Verteilung der Eigenschwingungen von Punktgittern. (German) [On the distribution of the natural vibrations of point grids]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 10, pages 249– 255. LCCN ???? Reprinted from Physik. Zschr. 14, 65–71 (1913).

Born:1963:SRG

[BvK63b] Max Born and Th. v. Kármán. Über Schwingungen in Raumgittern. (German) [On vibrations in space lattices]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 8, pages 231–243. LCCN ???? Reprinted from Physik. Zschr. 13, 297–309 (1912).

Born:1963:TSW

[BvK63c] Max Born and Th. v. Kármán. Zur Theorie der spezifischen Wärme. (German) [Toward the theory of specific heat]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63c], chapter 9, pages 244–248. LCCN ???? Reprinted from Physik. Zschr. 14, 15–19 (1913).

Born:1923:MA

[BvL23] Max Born and Max von Laue. Max Abraham. Physikalische Zeitschrift, 24(3):49–53, February 1, 1923. CODEN PHZTAO.

ISSN 0369-982X. URL https://jscholarship.library.jhu.
edu/handle/1774.2/214.

Born:1963:MAb

[BvL63] Max Born and Max von Laue. Max Abraham. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 79, pages 599–603. LCCN ???? Reprinted from Naturwiss. 16, 1035–1036 (1928).

Born:1923:BGR

[BvLM23] M. Born, M. v. Laue, and Lise Meitner. Besprechungen. (German) [Review]. Naturwissenschaften, 11(47):937–938, November 1923. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Born:1926:NFQ

[BW26a] Max Born and Norbert Wiener. Eine neue Formulierung der Quantengesetze für periodische und nichtperiodische Vorgänge. (German) [A new formulation of quantum laws for periodic and aperiodic processes]. Zeitschrift für Physik, 36(3):174–187, March 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01382261.

Born:1926:NFL

[BW26b] Max Born and Norbert Wiener. A new formulation of the laws of quantization of periodic and aperiodic phenomena. Journal of mathematics and physics / Massachusetts Institute of Technology, 5(??):84–98, ???? 1926. CODEN JMPHA9. ISSN 0097-1421.

Born:1931:EQA

[BW31a] Max Born and V. Weisskopf. Erratum: Quantenmechanik der Adsorptionskatalyse. (German) [Erratum: Quantum mechanics of adsorption catalysis]. Z. Phys. Chem., B12(??):478, ???? 1931. CO-DEN ???? ISSN ???? See [BW31b].

Born:1931:QAG

[BW31b] Max Born and V. Weisskopf. Quantenmechanik der Adsorptionskatalyse. (German) [Quantum mechanics of adsorption catalysis]. Z. Phys. Chem., B12(??):206–227, ???? 1931. CODEN ???? ISSN ???? See erratum [BW31a].

Born:1959:POE

[BW59] Max Born and Emil Wolf, editors. Principles of Optics, Electromagnetic Theory of Propagation, Interference and Diffraction of Light. Pergamon, New York, NY, USA, 1959. xxvi + 803 pp. LCCN ????

Born:1963:QAG

[BW63a] Max Born and Victor Weißkopf. Quantenmechanik der Adsorptionskatalyse. (German) [Quantum mechanics of adsorption catalysis]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 66, pages 395–416. LCCN ???? Reprinted from Z. physik. Chem. B 12, 206–227 (1931).

Born:1963:NFQ

[BW63b] Max Born and Norbert Wiener. Eine neue Formulierung der Quantengesetze für periodische und nichtperiodische Vorgänge. (German) [A new formulation of quantum laws for periodic and nonperiodic operations]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63d], chapter 55, pages 214–227. LCCN ???? Reprinted from Z. Physik 36, 174–187 (1926).

Born:1964:JF

[BW64a] Max Born and Wilhelm H. Westphal. James Franck. Physikalische Blätter, 20(7):324–328, July 1964. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/phbl.19640200704/
abstract.

Born:1964:POE

[BW64b] Max Born and Emil Wolf. Principles of Optics; Electromagnetic Theory of Propagation, Interference and Diffraction of Light. MacMillan Publishing Company, New York, NY, USA, second edition, 1964. xxviii + 808 pp. LCCN QC355 .B63 1964a. With contributions by A. B. Bhatia and others.

Born:1965:POE

[BW65] Max Born and Emil Wolf. Principles of Optics; Electromagnetic Theory of Propagation, Interference and Diffraction of Light. Pergamon, New York, NY, USA, third edition, 1965. xxviii + 808 pp. LCCN QC355 .B63 1965. With contributions by A. B. Bhatia and others.

Born:1969:POE

[BW69] Max Born and Emil Wolf. Principles of Optics; Electromagnetic Theory of Propagation, Interference and Diffraction of Light. Pergamon, New York, NY, USA, fourth edition, 1969. ISBN 0-08-013987-6. xxviii + 808 pp. LCCN QC355 .B63 1969. With contributions by A. B. Bhatia and others.

Born:1975:POE

[BW75] Max Born and Emil Wolf. Principles of Optics; Electromagnetic Theory of Propagation, Interference, and Diffraction of Light. Pergamon, New York, NY, USA, fifth edition, 1975. ISBN 0-08-018018-3. xxviii + 808 pp. LCCN QC355.2 .B67 1975. With contributions by A. B. Bhatia and others.

Born:1980:POE

[BW^{+}80] Max Born, Emil Wolf, et al. Principles of Optics: Electromagnetic Theory of Propagation, Interference and Diffraction of Light. Pergamon, New York, NY, USA, sixth edition, 1980. ISBN 0-08-026481-6 (paperback), 0-08-026482-4. xxvii + 808 pp. LCCN QC355.2 .B67 1980. With contributions by A. B. Bhatia and others.

Born:1997:POE

[BW97] Max Born and Emil Wolf. Principles of Optics: Electromagnetic Theory of Propagation, Interference and Diffraction of Light. Cambridge University Press, Cambridge, UK, sixth corrected edition, 1997. ISBN 0-521-63921-2 (paperback). xxviii + 808 pp.

LCCN QC355.2 .B67 1997. URL http://www.loc.gov/catdir/
description/cam029/98115588.html;   http://www.loc.gov/
catdir/toc/cam021/98115588.html. With contributions by A. B.

Bhatia and others.

Born:1999:POE

[BW99] Max Born and Emil Wolf. Principles of Optics: Electromagnetic Theory of Propagation, Interference and Diffraction of Light. Cambridge University Press, Cambridge, UK, seventh expanded edition, 1999. ISBN 0-521-64222-1, 0-521-63921-2 (paperback). xxxiii + 952 pp. LCCN QC355.2 .B67 1999. URL http://www.loc.gov/

catdir/description/cam029/98049429.html;   http://www.
loc.gov/catdir/samples/cam041/98049429.html;   http://
www.loc.gov/catdir/toc/cam021/98049429.html;   https://
www.cambridge.org/core/books/principles-of-optics/. With

contributions by A. B. Bhatia, P. C. Clemmow, D. Gabor, A. R. Stokes, A. M. Taylor, P. A. Wayman, and W. L. Wilcock.

Born:2019:POE

[BWB19] Max Born, Emil Wolf, and A. B. (Avadh Behari) Bhatia. Principles of Optics: Electromagnetic Theory of Propagation, Interference, and Diffraction of Light. Cambridge University Press, Cambridge, UK, seventh (expanded) anniversary edition, 60th anniversary edition, 2019. ISBN 1-108-47743-7 (hardcover), 1-108-76991-8 (e-pub). 990 (est.) pp. LCCN QC355.2 .B67

2019. URL https://assets.cambridge.org/97811084/77437/
toc/9781108477437_toc.pdf;   https://www.cambridge.org/
core/books/principles-of-optics/.

Born:1950:NSS

[BY50] Max Born and L. M. Yang. Nuclear shell structure and nuclear density. Nature, 166(4218):399, September 2, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v166/n4218/pdf/166399a0.pdf.

Cartwright:1989:BED

[Car89] Nancy Cartwright. The Born–Einstein debate: Where application and explanation separate. Synthese, 81(3):271–282, December 1989. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/article/10.
1007/BF00869317.

Cassidy:2005:BRB

[Cas05] David C. Cassidy. Book review: Born unto trouble: The End of the Certain World: The Life and Science of Max Born: The Nobel Physicist Who Ignited the Quantum Revolution, by Nancy Thorndike Greenspan. American Scientist, 93(4):372–374, July/ August 2005. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-

2786 (electronic). URL http://www.jstor.org/stable/pdfplus/
27858620.pdf.

Castro:2012:BRG

[Cas12] Carlos Castro. Born’s reciprocal gravity in curved phase-spaces and the cosmological constant. Foundations of Physics, 42(8): 1031–1055, August 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.com/

article/10.1007/s10701-012-9645-9.

Charles:2005:MMR

[Cha05] Daniel Charles. Master mind: the rise and fall of Fritz Haber, the Nobel laureate who launched the age of chemical warfare. Ecco, New York, NY, USA, 2005. ISBN 0-06-056272-2. xvii + 313 pp. LCCN QD22.H15 C48 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0910/2004057532-b.
html;   http://www.loc.gov/catdir/enhancements/fy0910/
2004057532-d.html.

Cheng:1970:GBD

[Che70] Chien-Chung Cheng. General Born diagram and Legendre transformation. American Journal of Physics, 38(8):956–??, August 1970. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v38/
i8/p956_s1.

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Condon:1962:YQP

[Con62] Edward U. Condon. 60 years of quantum physics. Physics Today, 15(10):37–49, October 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/PHTOAD/v15/i10. Delayed 1951 Presidential ad-

dress at the 1500th regular meeting of the American Philosophical Society of Washington, 2 December 1962, at the Natural History Museum Auditorium of the Smithsonian Institution, on the 60th anniversary of Planck’s constant, h. Reprinted in [WP85, pages 310–318].

Cook:2002:MB

[Coo02] Alan Cook. Max Born. Notes and Records of the Royal Society of London, 56(2):131–132, May 2002. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.

org/stable/pdfplus/3557662.pdf.

Copson:1961:BRB

[Cop61] E. T. Copson. Book review: Principles of Optics, by Max Born and Emil Wolf. Mathematical Gazette, 45(353):274, October 1961. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/stable/pdfplus/3612833.pdf.

Coughlan:2013:SWE

[Cou13] Sean Coughlan. The scientists who escaped the Nazis. BBC News Web story., July 17, 2013. URL http://www.bbc.com/news/ business-23261289. Interview with 92-year-old Gustav Born, son of Max and Hedi Born.

Cohen:1979:SPL

[CS79a] R. S. (Robert Sonné) Cohen and John J. Stachel, editors. Selected Papers of Léon Rosenfeld, volume 21 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. xxxiv + 929 pp. LCCN Q174 .B67 vol. 21 QC7. URL https:

//link.springer.com/book/10.1007/978-94-009-9349-5.

Cohen:1979:ECB

[CS79b] Robert S. Cohen and John J. Stachel. The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday). In Cohen and Stachel [CS79a], chapter II.10, pages 517–521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21

QC7. URL http://link.springer.com/chapter/10.1007/978-
94-009-9349-5_35.

Colles:1983:MBC

[CS83] M. John Colles and D. W. (David W.) Swift, editors. The Max Born Centenary Conference: OPTICS 82, ECOSA 82: September 7– 10, 1982, Edinburgh, Scotland, volume 369 of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 1983. ISBN 0-89252-404-9 (paperback). LCCN QC350 .M39 1982. URL http://link.

spie.org/PSISDG/0369/.

Condon:1937:BRR

[CZPS37] E. U. Condon, V. K. Zworykin, Leigh Page, and W. F. G. Swann. Book reviews: Recent books on physics: An Elementary Survey of Modern Physics, by G. F. Hull, The Physics of Electron Tubes, by L. R. Koller, L’Unité de la Force et L’Unité de la Matière dans la Conception Physique Uniforme du Monde, by Jan Bašta, Atomic Physics, by Max Born. Science (New Series), 86(2224):157–158, August 13, 1937. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/pdfplus/
1663721.pdf.

daCAndrade:1933:BRB

[da 33] E. N. da C.Andrade. Book review: Moderne Physik. Sieben Vorträge über Materie und Stahlung, by Max Born. Journal of Physical Chemistry, 37(6):830–831, January 1933. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic). See [Bor33a].

deBroglie:1953:LMO

[dB53] Louis de Broglie. L’interprétation de la mécanique ondulatoire à l’aide d’ondes à régions singulières. (French) [The interpretation of wave mechanics using waves with singular regions], 1953.

DeZela:2016:GTT

[De 16] F. De Zela. Gleason-type theorem for projective measurements, including qubits: The Born rule beyond quantum physics. Foundations of Physics, 46(10):1293–1306, October 2016. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-016-0020-0.

Demetrian:2003:NDR

[Dem03] M. Demetrian. A note on derivation of Rutherford formula within Born approximation. ArXiv Physics e-prints, February 2003. URL

http://adsabs.harvard.edu/abs/2003physics...2102D.

Dingler:1923:BRB

[Din23] Hugo Dingler. Book review: Die Relativitätstheorie Einsteins, by Max Born. Annalen der Philosophie, 3(4):631–632, ???? 1923. CODEN ???? ISSN 1866-6167. URL http://www.jstor.org/

stable/40601725;   http://www.jstor.org/stable/pdfplus/
40601725.pdf.

Dingle:1950:BRB

[Din50] Herbert Dingle. Book review: Natural Philosophy of Cause and Chance, by Max Born. The British Journal for the Philosophy of Science, 1(3):245–248, November 1950. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.

jstor.org/stable/pdfplus/685370.pdf.

Delbourgo:2008:BRP

[DL08] R. Delbourgo and D. Lashmar. Born reciprocity and the 1/r potential. Foundations of Physics, 38(11):995–1010, November 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-008-9247-8.

Deshpande:1969:BOT

[DM69] V. K. Deshpande and J. Mahanty. Born–Oppenheimer treatment of the hydrogen atom. American Journal of Physics, 37(8):823– ??, August 1969. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v37/i8/p823_s1.

Drezet:2017:HJB

[Dre17] A. Drezet. How to justify Born’s rule using the pilot wave theory of de Broglie? Annales de la Fondation Louis de Broglie, 42(1):103–

132, 2017. URL http://aflb.ensmp.fr/AFLB-421/aflb421m5.
htm.

Dyson:1963:FS

[Dys63] Freeman Dyson. Forum: On speculation. Bulletin of the Atomic Scientists, 19(7):32, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BB63a].

Einstein:1982:BGC

[EB82] Albert Einstein and Max Born. Briefwechsel: 1916–1955. (German) [Correspondence: 1916–1955]. Edition Erbrich, Frankfurt am Main, Germany, 1982. ISBN 3-88682-005-X. 330 pp. LCCN QC16.E5 A25 1982. With contributions by Hedwig Born, with an introduction by Bertrand Russell, and with a foreword by Werner Heisenberg. Reprint of the 1969 edition.

Einstein:1991:LMB

[Ein91] Albert Einstein. From Letters to Max Born. In Ferris and Fadiman [FF91], pages 808–809. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Enz:2002:BAG

[Enz02a] Charles P. (Charles Paul) Enz. Born’s assistant in Göttingen. In No time to be brief: a scientific biography of Wolfgang Pauli [Enz02b], pages 75–83. ISBN 0-19-856479-1. LCCN QC16.P37

E59 2002. URL http://www.loc.gov/catdir/enhancements/
fy0614/2002726902-d.html;   http://www.loc.gov/catdir/
enhancements/fy0614/2002726902-t.html.

Enz:2002:NTB

[Enz02b] Charles P. (Charles Paul) Enz. No time to be brief: a scientific biography of Wolfgang Pauli. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-856479-1. viii + 573 pp. LCCN QC16.P37 E59 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0614/2002726902-d.
html;   http://www.loc.gov/catdir/enhancements/fy0614/
2002726902-t.html.

Farago:1979:BIR

[Far79] P. S. Farago. Born’s intimate reminiscences: Book review: My Life: Recollections of a Nobel Laureate, by Max Born. Nature, 282 (5734):113, November 1, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v282/n5734/pdf/282113a0.pdf.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Fischer:2010:HQE

[Fis10a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2010:MB

[Fis10b] Ernst Peter Fischer. Max Born (1882–1970). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis10a], pages 94–105. ISBN 3-7766-2643-7. LCCN ????

Fischer:2012:HQE

[Fis12a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fischer:2012:MB

[Fis12b] Ernst Peter Fischer. Max Born (1882–1970). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis12a], pages 94–105. ISBN 3-596-19406-7. LCCN ????

Freire:2010:DMM

[FL10] Olival Freire and Christoph Lehner. ‘Dialectical materialism and modern physics’, an unpublished text by Max Born. Notes and Records of the Royal Society of London, 64(2):155–162, June 20, 2010. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Forman:1968:BRB

[For68] Paul Forman. Book review: My Life and My Views, by Max Born. Isis, 59(2):239–240, Summer 1968. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228305;   http://www.jstor.org/stable/
pdfplus/228305.pdf.

Forman:1970:BRBe

[For70] Paul Forman. Book review: Briefwechsel, 1916–1955, by Albert Einstein, Max Born, and Hedwig Born. Isis, 61(4):553–555, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229480;   http:
//www.jstor.org/stable/pdfplus/229480.pdf.

Fowler:1927:BRB

[Fow27] R. H. Fowler. Book review: The Mechanics of the Atom, by Max Born, J. W. Fisher, and D. R. Hartree. Mathematical Gazette, 13 (190):428–429, October 1927. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.jstor.

org/stable/pdfplus/3602769.pdf.

Fedak:2009:BJP

[FP09] William A. Fedak and Jeffrey J. Prentis. The 1925 Born and Jordan paper “On quantum mechanics”. American Journal of Physics, 77(2):128–??, February 2009. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.

org/resource/1/ajpias/v77/i2/p128_s1.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Frenkel:1934:BTE

[Fre34] J. Frenkel. On Born’s theory of the electron. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 146(859):930–934, October 15, 1934. CODEN PRLAAZ.

ISSN 0080-4630. URL http://rspa.royalsocietypublishing.
org/content/146/859/930. See [BI34c].

Freire:2001:SPP

[Fre01] Olival Freire, Jr. Science, philosophy and politics in the fifties: on the Max Born’s unpublished paper entitled “Dialectical materialism and modern physics”. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 10(3):248–254, March 2001. CODEN HISCDU. ISSN 0285-4821.

Fries:2008:CW

[FW08] Amos A. Fries and Clarence J. West. Chemical Warfare (1921). Kessinger Publishing, LLC, ????, 2008. ISBN 0-548-99545-1. 460 (est.) pp. LCCN ????

Galles:2001:BCT

[Gal01] Carlos D. Galles. Born coined the term [quantum mechanics]. Physics Today, 54(5):94, April 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v54/i4/p94_s1.

Gangopadhyay:2004:BOA

[GDR04] Gautam Gangopadhyay and Binayak Dutta-Roy. The Born– Oppenheimer approximation: a toy version. American Journal of Physics, 72(3):389–??, March 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v72/i3/p389_s1.

Greenspan:2008:MBB

[GE08] Nancy Thorndike Greenspan and Anita Ehlers. Max Born — Baumeister der Quantenwelt: eine Biographie. (German) [Max Born — builder of the quantum world: a biography]. Spektrum Taschenbuch. Spektrum Akademischer Verlag, Heidelberg, Germany, 2008. ISBN 3-8274-2080-6 (paperback). xvii + 380

pp. LCCN ???? URL http://deposit.d-nb.de/cgi-bin/
dokserv?id=3123772\%26prov=M\%26dok\_var=1\%26dok\_ext=
htm;   http://www.gbv.de/dms/weimar/toc/569653967_toc.

pdf. Translation to German by Anita Ehlers of [Gre05a].

Galison:2002:QMS

[GGK02] Peter Galison, Michael Gordin, and David Kaiser, editors. Quantum Mechanics: Science and Society. Routledge, London, UK, 2002. ISBN 1-136-70972-X. 433 (est.) pp. LCCN ????

Giulini:2013:MBB

[Giu13] Domenico Giulini. Max Born’s Vorlesungen über Atommechanik, Erster Band. In Research and Pedagogy: a History of Quantum Physics Through Its Textbooks [BN13], page ?? ISBN 3-8442-5871-

X. LCCN QC173.98. URL http://www.edition-open-access.
de/studies/2/.

GoeppertMayer:1988:MGM

[Goe88] Maria Goeppert Mayer. Maria Goeppert Mayer papers, 1925– 1973, MSS 20. Collection at the University of California,

San Diego., 1988. URL http://libraries.ucsd.edu/speccoll/
findingaids/mss0020.html.

Gottstein:1983:CGE

[Got83] Klaus Gottstein. Commentary: The Göttingen Eighteen. Bulletin of the Atomic Scientists, 39(8):62–63, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gowing:1979:ERM

[Gow79] Margaret Gowing. Essay review: My life: Recollections of a Nobel laureate, by Max Born. Contemporary Physics, 20(4):485–486, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Greenspan:2005:ECW

[Gre05a] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Greenspan:2005:MBP

[Gre05b] Nancy Thorndike Greenspan. Max Born and the peace movement. Physics World, 18(4):35–38, April 3, 2005. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/pwa-
pdf/18/4/phwv18i4a37.pdf.

Greenspan:2006:MBB

[Gre06] Nancy Thorndike Greenspan. Max Born — Baumeister der Quantenwelt. (German) [Max Born — builder of the quantum world]. Elsevier, Spektrum Akademischer Verlag, Munich, Germany, 2006. ISBN 3-8274-1640-X. xvii + 380 pp. Translation to German by Anita Ehlers of [Gre05a].

Gross:1947:BTS

[GS47] E. Gross and A. Stehanov. Born’s theory and the spectrum of light scattering of crystals. Nature, 159(4040):474–475, April 5, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v159/
n4040/pdf/159474a0.pdf.

Gimeno:2021:OCL

[GXB21] Gonzalo Gimeno, Mercedes Xipell, and Marià Baig. Operator calculus: the lost formulation of quantum mechanics. Archive for History of Exact Sciences, 75(3):283–322, May 2021. CODEN AH-ESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Haber:1919:BTW

[Hab19] F. (Fritz) Haber. Betrachtungen zur Theorie der Wärmetönung. (German) [Reflections on the theory of heat tint]. Verhandlungen der Deutschen Physikalischen Gesellschaft, 21(21–22):750–768, December 5, 1919. CODEN VDPEAZ. ISSN 0372-5448. This paper, with Born’s independent work [Bor19c], is the origin of the famous “Born–Haber cycle” in thermodynamics.

Haber:1922:LEC

[Hab22] Fritz Haber. Letter to the Editor: Chemical warfare. Nature, 104(2724):40, January 12, 1922. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v109/n2724/pdf/109040a0.pdf. See sharp

rebuttal [Tho22].

Haber:1986:PCC

[Hab86] L. F. (Ludwig Fritz) Haber. The poisonous cloud: chemical warfare in the First World War. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-858142-4. xiv + 415 pp. LCCN UG447 .H255 1986. US$33.55. URL http://www.loc.

gov/catdir/enhancements/fy0604/85010638-d.html;   http:
//www.nobel.se/chemistry/laureates/1918/haber-bio.html.

Hager:2008:AAJ

[Hag08] Thomas Hager. The alchemy of air: a Jewish genius, a doomed tycoon, and the scientific discovery that fed the world but fueled the rise of Hitler. Harmony Books, New York, NY, USA, 2008. ISBN 0-307-35178-5. xvii + 316 pp. LCCN QD21 .H26 2008. URL

http://www.loc.gov/catdir/toc/ecip089/2008003192.html.

Haidar:2012:MB

[Hai12] Riad Haidar. Max Born. Photoniques, 61(??):20–21, September/October 2012. CODEN ???? ISSN 1629-4475 (print), 2269-

8418 (electronic). URL http://www.photoniques.com/articles/
photon/abs/2012/05/photon201261p20/photon201261p20.html.

Hanle:1980:BRB

[Han80] Paul A. Hanle. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. Isis, 71(2):354–355, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230243;   http:
//www.jstor.org/stable/pdfplus/230243.pdf.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Heisenberg:1961:MPb

[HBSA61] Werner Heisenberg, Max Born, Erwin Schrödinger, and Pierre Auger, editors. On modern physics. Clarkson N. Potter. Inc., New York, NY, USA, 1961. 108 pp. LCCN QC71 .O5.

Hecht:2000:BRB

[Hec00] Eugene Hecht. Book review: Principles of Optics: Electromagnetic Theory of Propagation, Interference and Diffraction of Light, by Max Born and Emil Wolf. Physics Today, 53(10):77, October 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.org/resource/1/
phtoad/v53/i10/p77_s1.

Heisenberg:1925:QUK

[Hei25] W. Heisenberg. Über quantentheoretische Umdeutung kinematischer und mechanischer Beziehungen. (German) [On quantum theoretical reinterpretation of kinematic and mechanical relations]. Zeitschrift für Physik, 33(1):879–893, December 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01328377;
http://www.springerlink.com/content/w22847j178u03029/.

Heisenberg:1962:MBA

[Hei62] W. Heisenberg. Max Born zum achtzigsten Geburtstag. (German) [Max Born on his eightieth birthday]. Naturwissenschaften, 49(23):529, January 1962. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/p138q381372tp36r.

Heilbron:1972:BRP

[Hei72] John L. Heilbron. Book review: Physics in My Generation by Max Born; Bahnbrecher des Atomzeitalters. Grosse Naturforscher von Maxwell bis Heisenberg by Friedrich Herneck; Moseley and the Numbering of the Elements by Bernard Jaffe; The Big Machine by Robert Jungk. Isis, 63(1):111–112, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229207;   http:
//www.jstor.org/stable/pdfplus/229207.pdf.

Heilbron:1979:BRB

[Hei79a] J. L. Heilbron. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. Science (New Series), 204(4394):740–741, May 18, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/pdfplus/
1748293.pdf.

Heilbron:1979:MBB

[Hei79b] J. L. Heilbron. Max Born: Book review: My Life: Recollections of a Nobel Laureate. Science, 204(4394):740–741, May 18, 1979. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/204/4394/
740.full.pdf.

Haber:1970:MLF

[HH70] Charlotte Haber and Fritz Haber. Mein Leben mit Fritz Haber: Spiegelungen der Vergangenheit. (German) [My life with Fritz Haber: reflections of the past]. Econ, Düsseldorf, West Germany, 1970. 293 pp. LCCN ????

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hermann:1932:BGR

[HKP^{+}32a] C. Hermann, O. Kratky, F. Paneth, Max Born, Arnold Sommerfeld, Walther Gerlach, and Max Volmer. Besprechungen: Wyckoff, R. W. G., The structure of crystals, second edition. P. P. Ewald and C. Hermann, Strukturbericht 1913–1928. (German) [Book reviews]. Naturwissenschaften, 20(29):544–550, July 15, 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01503765.

Hermann:1932:BWR

[HKP^{+}32b] C. Hermann, O. Kratky, F. Paneth, Max Born, Arnold Sommerfeld, Walther Gerlach, and Max Volmer. Besprechungen: Wyckoff, R. W. G., The structure of crystals, second edition. P. P. Ewald and C. Hermann, Strukturbericht 1913–1928. (German) [Book reviews]. Naturwissenschaften, 20(29):544–550, July 15, 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01503765.

Hund:1988:PGF

[HMLM88] F. Hund, H. Maier-Leibnitz, and E. Mollwo. Physics in Gottingen — with Franck, Born and Pohl. European Journal of Physics, 9(3): 188–194, July 1988. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

Holley:1988:BRB

[Hol88] I. B. Holley, Jr. Book review: The poisonous cloud: chemical warfare in the First World War, by L. F. Haber. Technology and Culture, 29(1):158–159, January 1988. ISSN 0040-165X (print), 1097-

3729 (electronic). URL http://www.jstor.org/stable/info/

3105250. See [Hab86].

Holstein:2007:SBA

[Hol07] Barry R. Holstein. Second Born approximation and Coulomb scattering. American Journal of Physics, 75(6):537–??, June 2007. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v75/
i6/p537_s1.

Hovakimian:2018:VCB

[Hov18] Levon B. Hovakimian. On a validity criterion for the Born approximation. Armenian Journal of Physics, 11(4):273–277, 2018. CODEN AJPRCN. ISSN 1829-1171. URL http://ajp.asj-oa.

am/1005/.

Harris:1982:HFK

[HP82] Robert Harris and Jeremy Paxman. A higher form of killing: the secret story of gas and germ warfare. Chatto and Windus, London, UK, 1982. ISBN 0-7011-2585-3. xii + 274 + 16 pp. LCCN UG447.8 .H37 1982.

Harris:2002:HFK

[HP02] Robert Harris and Jeremy Paxman. A higher form of killing: the secret history of chemical and biological warfare. Random House Trade Paperbacks, New York, NY, USA, 2002. ISBN 0-8129-6653-8 (paperback). 301 pp. LCCN UG447 .H367 2002. URL http://www.

loc.gov/catdir/bios/random053/2002017948.html;   http:
//www.loc.gov/catdir/description/random042/2002017948.
html;   http://www.loc.gov/catdir/samples/random041/2002017948.
html.

Heisenberg:1960:DSF

[HSBA60] Werner Heisenberg, Erwin Schrödinger, Max Born, and Pierre Auger. Discussione sulla fisica moderna. (Italian) [Discussions on modern physics], volume 1952 of Quattro conferenze organizzate dalle “Rencontres internationales de Genève”. Biblioteca di cultura scientifica, 59. Paolo Boringhieri, Torino, 1960. v + 131 + 1 pp.

Ilyin:2016:BRT

[Ily16] Aleksey V. Ilyin. The Born rule and time-reversal symmetry of quantum equations of motion. Foundations of Physics, 46(7):845– 851, July 2016. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-016-0006-y.

Infeld:1955:GRG

[Inf55] Leopold Infeld. Die Geschichte der Relativitätstheorie. (German) [The history of the Theory of Relativity]. Naturwissenschaften, 42 (15):431–436, August 1, 1955. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Jammer:1974:ESI

[Jam74a] Max Jammer. Early semiclassical interpretations. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74b], chapter 2, pages 20–54. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:PQM

[Jam74b] Max Jammer. The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective. J. Wiley and Sons, New York, NY, USA, 1974. ISBN 0-471-43958-4. xi + 536 pp. LCCN QC173.98 .J35.

Johnson:2013:CMG

[Joh13] Jeffrey Allan Johnson. The case of the missing German quantum chemists: On molecular models, mobilization, and the paradoxes of modernizing chemistry in Nazi Germany. Historical Studies in the Natural Sciences, 43(4):391–452, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/hsns.2013.43.4.391.

Jones:2005:BA

[Jon05] Clifford Jones. Born, again. Physics World, 18(5):18, May 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (elec-

tronic). URL http://physicsworldarchive.iop.org/full/pwa-
pdf/18/5/phwv18i5a29.pdf.

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Kamp:1983:JFM

[K^{+}83] Norbert Kamp et al. James Franck und Max Born in Göttingen, volume 69 of Góttinger Universitätsreden. Vandenhoeck and Ruprecht, Göttingen, Germany, 1983. ISBN 3-525-82621-4. 37 pp. LCCN QC16.F67 J36 1983.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kockel:1960:BHW

[KED^{+}60] B. Kockel, W. H. Estphal, F. Dessauer, K. Strubecker, H. Siedentopf, F. Borgnis, H. Falkenhagen, W. Finkelnburg, and R. Vieweg. Bücher: Heisenberg: Wandlungen in den Grundlagen der Naturwissenschaft/Heisenberg: Physik und Philosophie/Frisch: Beiträge zur Physik und Chemie des 20. Jahrhunderts/Bellman: Introduction to Matrix Analysis/Mehlin: Astronomy/Born und Wolf: Principles of Optics/Mandl: Introduction to Quantum Field Theory/ Marley, Morgan: Health Physics/Churchman: Measurement Definitions and Theories. Physikalische Blätter, 16(6):344–347, June 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19600160610/abstract.

Khrennikov:2009:DMB

[Khr09] Andrei Khrennikov. Detection model based on representation of quantum particles by classical random fields: Born’s rule and beyond. Foundations of Physics, 39(9):997–1022, September 2009. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-009-9312-y.

Klein:1970:BRBb

[Kle70a] Martin J. Klein. Book review: Briefwechsel, 1916–1955. Albert Einstein and Hedwig and Max Born. Commentary by Max Born. Nymphenburger, Munich, 1969. 332 pp. DM 24.80. Science, 169 (3943):360–361, July 24, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/169/3943/360.3.full.pdf.

Klein:1970:BRBa

[Kle70b] Martin J. Klein. Book review: My Life and My Views. Max Born. Scribner, New York, 1968. vi, 216 pp. $4.95. Science, 169 (3943):360–361, July 24, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/169/3943/360.2.full.pdf.

Klein:1970:BRM

[Kle70c] Martin J. Klein. Book review: Max Born on his vocation: Physics in My Generation, by Max Born. Second edition. Springer-Verlag, New York, 1969. viii, 172 pp., illus. Paper, $3.80. Heidelberg Science Library. Science (New Series), 169(3943):360–361, July 24, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/pdfplus/1729567.pdf;
http://www.sciencemag.org/content/169/3943/360.1.full.
pdf.

Konno:1978:HRB

[Kon78] Hiroyuki Konno. The historical roots of Born’s probabilistic interpretation. Japanese Studies in the History of Science, 17(??): 129–145, ???? 1978. CODEN JSHIAE. ISSN 0090-0176.

Korner:1959:BRB

[Kör59] S. Körner. Book review: Experiment and Theory in Physics, by Max Born. The British Journal for the Philosophy of Science, 9(36):337–339, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.

org/stable/pdfplus/685199.pdf.

Kozack:1991:BAD

[Koz91] Richard E. Kozack. Born approximation and differential cross sections in nuclear physics. American Journal of Physics, 59(1): 74–??, January 1991. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v59/i1/p74_s1.

Kroner:1998:RMG

[Krö98] Hans-Peter Kröner. Rezensionen: Mitchell G. Ash/Alfons Söllner (Hrsgg.): Forced Emigration and Scientific Change. Emigré German-Speaking Scientists and Scholars after 1933. Cambridge: Cambridge University Press 1995. Berichte zur Wissenschaftsgeschichte, 21(2–3):196–197, ???? 1998. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Kemmer:1971:MB

[KS71] N. Kemmer and R. Schlapp. Max Born. 1882–1970. Biographical Memoirs of Fellows of the Royal Society, 17:17–52, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769700;   http:
//www.jstor.org/stable/pdfplus/769700.pdf.

Landsberg:1983:BCR

[Lan83] P. T. Landsberg. The Born Centenary: Remarks about classical thermodynamics. American Journal of Physics, 51(9):842– ??, September 1983. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v51/i9/p842_s1.

Lapidus:1969:BSS

[Lap69] I. Richard Lapidus. Born series for scattering by a one-dimensional delta-function potential. American Journal of Physics, 37(10): 1064–??, October 1969. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v37/i10/p1064_s2.

Lax:1955:BRB

[Lax55] M. Lax. Book review: Dynamical Theory of Crystal Lattices, by Max Born and Kun Huang. American Journal of Physics, 23(7): 474, October 1955. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v23/i7/p474_s1.

Lande:1968:DD

[LBB68] Alfred Landé, Max Born, and Walter Biem. Dialog on dualism. Physics Today, 21(8):55–56, August 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/link.aip.org/link/?PTO/21/55/1;   http://physicstoday.
org/resource/1/phtoad/v21/i8/p55_s1. Replies by Alfred

Landé to points raised in the preceding article and further comments by Max Born and Walter Biem. See [BB68].

Lande:1969:AQG

[LBBH69] Alfred Landé, Max Born, Walter Biem, and W. Heisenberg. Auffassungen über die Quantentheorie. (German) [Conceptions about quantum theory]. Physik Journal, 25(3):105–113, March 1969. CO-DEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Lemmerich:1982:MBJ

[Lem82] Jost Lemmerich. Max Born, James Franck, Physiker in ihrer Zeit: der Luxus d. Gewissens. (German) [Max Born, James Franck, physicists in their time: The Luxury of Conscience], volume 17 of Staatsbibliothek Preußischer Kulturbesitz. Ausstellungskataloge. Staatsbibliothek Preußischer Kulturbesitz, Berlin, Germany, 1982. ISBN 3-88226-148-X. xi + 1884 pp. LCCN QC7 .M33 1982. With contributions from Friedrich Hund, H. Maier-Leibnitz and Victor F. Weisskopf.

Lemmerich:1983:WPN

[Lem83] Jost Lemmerich. Der wissenschaftliche und publizistische Nachlaß Max Borns: Katalog. (German) [The Scientific and Journalistic Estate of Max Born: Catalog]. Harrassowitz, Wiesbaden, West Germany, 1983. ISBN 3-447-02306-6. 480 (est.) pp. LCCN ???? With the collaboration of Anita Kerkmann.

Lindsay:1958:BRB

[Lin58] R. Bruce Lindsay. Book review: Physik im Wandel meiner Zeit, by Max Born. Physics Today, 11(2):28, February 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/11/28/1;
http://physicstoday.org/resource/1/phtoad/v11/i2/p28_s1.

Lindner:1998:BRB

[Lin98] Erik Lindner. Book review: Produktion und Distribution wissenschaftlicher Literatur. Der Physiker Max Born und sein Verleger Ferdinand Springer 1913–1970, (Sonderdruck aus dem Archiv fur Geschichte des Buchwesens Bd. 45), by Frank Holl. Zeitschrift für Unternehmensgeschichte / Journal of Business History, 43(1): 106–107, ???? 1998. CODEN ???? ISSN 0342-2852. URL http:/

/www.jstor.org/stable/40695837;   http://www.jstor.org/
stable/pdfplus/40695837.pdf.

Lorentz:1910:ANF

[Lor10] H. A. (Hendrik Antoon) Lorentz. Alte und neue Fragen der Physik. 1. Vortrag. Über die Entwicklung unserer Vorstellungen vom Äther. (German) [Old and new questions of physics. Lecture 1: On the development of our ideas about the ether]. Physikalische Zeitschrift, 11(26):1234–1257, December 15, 1910. CODEN PHZ-

TAO. ISSN 0369-982X. URL http://babel.hathitrust.org/
cgi/pt?id=njp.32101054770845%3Bseq=794%3Bview=1up. A di-

gest of six lectures by H. A. Lorentz.

Mandolesi:2018:AWP

[Man18] André L. G. Mandolesi. Analysis of Wallace’s proof of the Born rule in Everettian quantum mechanics: Formal aspects. Foundations of Physics, 48(7):751–782, July 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Mandolesi:2019:AWP

[Man19] André L. G. Mandolesi. Analysis of Wallace’s proof of the Born rule in Everettian quantum mechanics II: Concepts and axioms. Foundations of Physics, 49(1):24–52, January 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Margenau:1950:BRB

[Mar50] Henry Margenau. Book review: Natural Philosophy of Cause and Chance, by Max Born. The Review of Metaphysics, 4(1):129– 130, September 1950. CODEN ???? ISSN 0034-6632 (print), 1549-4853 (electronic), 2154-1302. URL http://www.jstor.org/

stable/20123200;   http://www.jstor.org/stable/pdfplus/
20123200.pdf.

Margolin:1973:BRB

[Mar73] Jean-Claude Margolin. Book review: Correspondance 1916-1955 by Albert Einstein, Max Born, Bertrand Russell, Werner Heisenberg, and Pierre Leccia. Les Études philosophiques, 1(??):75, January/ March 1973. CODEN ???? ISSN 0014-2166 (print), 2101-0056

(electronic). URL http://www.jstor.org/stable/20846392;
http://www.jstor.org/stable/pdfplus/20846392.pdf.

Matthew:1978:MB

[Mat78] J. A. D. Matthew. Max Born 1882–1970. Physics Education, 13(4): 251–254, 1978. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-

6552 (electronic). URL http://stacks.iop.org/0031-9120/13/
i=4/a=010.

McCrea:1949:BRB

[McC49] W. H. McCrea. Book review: Natural Philosophy of Cause and Chance, by Max Born. Mathematical Gazette, 33(304):131–132, May 1949. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL http://www.jstor.org/stable/pdfplus/
3610830.pdf.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Milne:1949:BRB

[Mil49] E. A. Milne. Book review: Natural Philosophy of Cause and Chance, by Max Born. Philosophy, 24(91):370–372, October 1949. CODEN ???? ISSN 0031-8191 (print), 1469-817X (elec-

tronic). URL http://www.jstor.org/stable/3747181;   http:
//www.jstor.org/stable/pdfplus/3747181.pdf.

Milburn:1960:BRB

[Mil60] Richard H. Milburn. Book review: Principles of Optics, by Max Born and Emil Wolf. American Scientist, 48(2):166A, June 1960. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-

2786 (electronic). URL http://www.jstor.org/stable/pdfplus/
27827547.pdf.

Misra:1940:SCL

[Mis40] Rama Dhar Misra. On the stability of crystal lattices. II. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36(2):173–182, April 1940. CODEN PCPSA4. ISSN 0008-1981.

Morrison:1954:BRB

[Mor54] P. Morrison. Book review: Scientific Papers Presented to Max Born. Sir Edward Appleton et al. Hafner, New York, 1953. 94 pp.

Illus. $2.50. Science, 119(3095):550, April 23, 1954. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/119/3095/550.1.full.pdf.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Morrison:1954:BRP

[MPK^{+}54] P. Morrison, W. D. Parkinson, James S. Koehler, Martin D. Kamen, and Alan B. Macnee. Book review: Physics: Scientific Papers Presented to Max Born by Edward Appleton, book review: Thunderstorm Electricity, by Horace R. Byers, book review: Dislocations and Plastic Flow in Crystals, by A. H. Cottrell, book review: Dislocations in Crystals, by W. T. Read, book review: Radioactive Isotopes, by W. J. Whitehouse and J. L. Putnam, book review: Principles of Transistor Circuits, by Richard F. Shea. Science (New Series), 119(3095):550–553, April 23, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/pdfplus/1683166.pdf.

Morris:1969:BFH

[MS69] D. F. C. Morris and E. L. Short. The Born–Fajans–Haber correlation. Nature, 224(5223):950–952, December 6, 1969. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v224/
n5223/pdf/224950a0.pdf.

Miller:2010:FH

[MVM10] Frederic P. Miller, Agnes F. Vandome, and John McBrewster. Fritz Haber. VDM Publishing House Ltd., ????, 2010. ISBN 613-0-70202-7. 202 pp. LCCN ????

Nikolic:2008:WBB

[Nik08] Hrvoje Nikolić. Would Bohr be born if Bohm were born before Born? American Journal of Physics, 76(2):143–??, February 2008. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v76/
i2/p143_s1.

Nordheim:1971:BRB

[Nor71] Lothar W. Nordheim. Book review: Physics in My Generation, by Max Born. Physics Today, 24(6):53–??, June 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/24/53/2;
 http://physicstoday.org/resource/1/phtoad/v24/i6/p53_
s2.

Oesper:1968:BRB

[Oes68] Ralph E. Oesper. Book review: My Life and My Views, by Max Born. Journal of Chemical Education, 45(12):A988, December 1968. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Pachucki:2010:BOP

[Pac10] K. Pachucki. Born–Oppenheimer potential for H_{2}. Physical Review A (Atomic, Molecular, and Optical Physics), 82(3):032509– ??, September 2010. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://adsabs.harvard.

edu/abs/2010PhRvA..82c2509P.

Page:1963:BRB

[Pag63] Thornton Page. Book review: Physics and Politics, by Max Born. American Journal of Physics, 31(8):633–634, August 1963. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v31/i8/p633_
s2;   http://link.aip.org/link/?AJP/31/633/2.

Pais:1982:MBS

[Pai82] A. Pais. Max Born’s statistical interpretation of quantum mechanics. Science, 218(4578):1193–1198, December 17, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/1688979.
pdf;   http://www.sciencemag.org/content/218/4578/1193.
full.pdf.

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/toc/fy02/99046603.html.

Palazzo:2000:SVW

[Pal00] Albert Palazzo. Seeking victory on the western front: the British army and chemical warfare in World War I. University of Nebraska Press, Lincoln, NE, USA, 2000. ISBN 0-8032-3725-1. xii + 239 pp.

LCCN D639.C39 P35 2000. URL http://www.loc.gov/catdir/
enhancements/fy0709/99042637-b.html;   http://www.loc.
gov/catdir/enhancements/fy0709/99042637-d.html.

Pawlak:2008:HBM

[Paw08] Alexander Pawlak. Happy birthday, Max Born. Physik Journal, 7 (1):9–10, ???? 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Peierls:1971:SIB

[Pei71] Rudolf Peierls. Strong interactions [Born–Einstein correspondence]. New Scientist, ??(??):339, May 6, 1971. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Peierls:1997:AH

[Pei97] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Pryce:1960:BRB

[Pry60] M. H. L. Pryce. Book review: Physics in My Generation, by Max Born. The British Journal for the Philosophy of Science, 11(42):157–159, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.

org/stable/pdfplus/685593.pdf.

Panati:2007:TDB

[PST07] Gianlucaa Panati, Herberta Spohn, and Stefana Teufel. The timedependent Born–Oppenheimer approximation. Mathematical modelling and numerical analysis = Modelisation mathématique et analyse numérique: M^{2}AN, 41(??):297–314, March 2007. CODEN RMMAEV. ISSN 0764-583X (print), 1290-3841 (electronic).

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Romain:1965:BRB

[Rom65] Jacques E. Romain. Book review: Die Relativitätstheorie Einsteins, by Max Born. Physics Today, 18(6):54, June 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/18/54/1;
 http://physicstoday.org/resource/1/phtoad/v18/i6/p54_
s1.

Rosenfeld:1979:ECB

[Ros79] Léon Rosenfeld. The epistemological conflict between Einstein and Bohr (dedicated to Max Born on his 80th birthday). In Cohen and Stachel [CS79a], chapter II.10, pages 517–521. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL http://link.

springer.com/chapter/10.1007/978-94-009-9349-5_35.

Rushbrooke:1951:BGT

[RS51] G. S. Rushbrooke and H. I. Scoins. Born and Green’s theory of imperfect gases. Nature, 167(4244):366–367, March 3, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v167/
n4244/pdf/167366a0.pdf.

Renn:2007:GGRb

[RSS^{+}07] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 4. Gravitation in the Twilight of Classical Physics: the Promise of Mathematics, volume 250(4) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. 621–1152 pp. LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/985765933/34;   http:
//www.loc.gov/catdir/enhancements/fy0824/2007440325-d.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2007440325-t.html;   https://link.springer.com/book/10.
1007/978-1-4020-4000-9.

Renneberg:1994:STN

[RW94] Monika Renneberg and Mark Walker, editors. Science, Technology, and National Socialism. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-40374-X (hardcover), 0-521-52860-7 (paperback). xix + 422 pp. LCCN Q127.G3 S36 1994. URL http:/

/www.loc.gov/catdir/description/cam025/92041633.html;
http://www.loc.gov/catdir/samples/cam031/92041633.html;
 http://www.loc.gov/catdir/toc/cam028/92041633.html.

Salvia:2019:ECP

[Sal19] Stefano Salvia. Embattled cooperation(s): Peaceful atoms, pacifist physicists, and partisans of peace in the Early Cold War (1947– 1957). Physics in Perspective (PIP), 21(1):43–62, March 2019. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Schilpp:1949:AEPa

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schlegel:1979:BRB

[Sch79] Richard Schlegel. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. American Scientist, 67(5):623, September 1979. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-

2786 (electronic). URL http://www.jstor.org/stable/pdfplus/
27849521.pdf.

Schirrmacher:2005:DMA

[Sch05] Arne Schirrmacher, editor. Dreier Männer Arbeit in der frühen Bundesrepublik: Max Born, Werner Heisenberg und Pascual Jordan als politische Grenzgänger. (German) [Three men’s work in the early Federal Republic: Max Born, Werner Heisenberg and Pascual Jordan as political border crossers], volume 296 of Preprint / Max-Planck-Institut für Wissenschaftsgeschichte. Max-Planck-Inst. für Wissenschaftsgeschichte, Berlin, Germany, 2005. 50 + 6 pp. LCCN ????

Schirrmacher:2007:PPF

[Sch07] Arne Schirrmacher. Physik und Politik in der frühen Bundesrepublik Deutschland. Max Born, Werner Heisenberg und Pascual Jordan als politische Grenzgänger. (German) [Physics and politics in the early Federal Republic Germany. Max Born, Werner Heisenberg and Jordan as a political frontier]. Berichte zur Wissenschaftsgeschichte, 30(1):13–31, March 2007. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Segre:1964:BRB

[Seg64] Emilio Segrè. Book review: Ausgewählte Abhandlungen, by Max Born. Physics Today, 17(1):84–86, January 1964. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/17/84/2;
http://physicstoday.org/resource/1/phtoad/v17/i1/p84_s2.

Schlosshauer:2005:ZDB

[SF05] Maximilian Schlosshauer and Arthur Fine. On Zurek’s derivation of the Born rule. Foundations of Physics, 35(2):197–213, February 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/10.
1007/s10701-004-1941-6.

Schomaker:1952:BAE

[SG52] Verner Schomaker and Roy Glauber. The Born approximation in electron diffraction. Nature, 170(4320):290–291, August 16, 1952. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v170/
n4320/pdf/170290b0.pdf.

Stein:1982:LBA

[SG82] David E. Stein and Alex E. S. Green. Levinson-modified Born approximation as applied to the spherical square well and to Mie theory. American Journal of Physics, 50(12):1120–??, December 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v50/
i12/p1120_s1.

Sollner:2012:TBR

[SGM^{+}12] Immo Söllner, Benjamin Gschösser, Patrick Mai, Benedikt Pressl, and Gregor Weihs Zoltán Vörös. Testing Born’s rule in quantum mechanics for three mutually exclusive events. Foundations of Physics, 42(6):742–751, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/s10701-011-9597-5.

Shorter:2020:BRP

[Sho20] R. S. Shorter. Book review: Principles of optics, 7th edition — 60th anniversary edition, by M. Born and E. Wolf, Cambridge, Cambridge University Press, 2019, 992 pp., £54.99 (hardback), ISBN: 978-1-108-47743-7. Contemporary Physics, 61(3):217–218, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sigurdsson:1996:PLC

[Sig96] Skúli Sigurdsson. Physics, life, and contingency: Born, Schrödinger, and Weyl in exile. In Ash and Söllner [AS96], pages 48–70. ISBN 0-521-49741-8 (hardcover). LCCN E184.G3 F73 1996. URL http:/

/www.loc.gov/catdir/description/cam027/95024894.html;
http://www.loc.gov/catdir/samples/cam031/95024894.html;
 http://www.loc.gov/catdir/toc/cam026/95024894.html.

Szollosi-Janze:1998:FHB

[SJ98] Margit Szöllösi-Janze. Fritz Haber, 1868–1934: eine Biographie. C. H. Beck, München, Germany, 1998. ISBN 3-406-43548-3. 928 pp. LCCN QD22.H15 S96 1998. DM 98.00. URL http:/

/www.gbv.de/dms/faz-rez/FR11999081090553.pdf;   http://
www.gbv.de/dms/ilmenau/toc/239809467.PDF.

Saxena:1966:BCT

[SK66] S. C. Saxena and C. M. Kachhava. Born’s continuum theory of solids. American Journal of Physics, 34(8):704–??, August 1966. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v34/
i8/p704_s1.

Sommerfeld:1915:BMD

[Som15] Arnold Sommerfeld. Born, M.: Dynamik der Kristallgitter. Besprechung. (German) [Born, M.: Dynamics of crystal lattices. Review]. Naturwissenschaften, 3(50):669–670, December 10, 1915. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Sommerfeld:1933:RBS

[Som33] Arnold Sommerfeld. Rezension: Born–Sauter, Moderne Physik. (German) [Review: Born–Sauter, Modern Physics]. Metallkunde, 13(??):24–??, ???? 1933.

Sommerfeld:1934:BMS

[Som34] Arnold Sommerfeld. Born, M. und Sauter, F.: Sieben Vorträge über Materie und Strahlung. — Besprechung. (German) [Born, M. and Sauter, F.: Seven Lectures on Matter and Radiation. — Review]. Metallwirtschaft, Metallwissenschaft, Metalltechnik, 13(??):26–??, ???? 1934. CODEN METWAH. ISSN 0368-9581.

Sommerfeld:1950:BMN

[Som50] Arnold Sommerfeld. Born, M.: Natural Philosophy of cause and chance. — Besprechung. Zeitschrift für Naturforschung, 5(??):571– 572, ???? 1950. CODEN ZNTFA2. ISSN 0372-9516.

Sommer:2005:DBH

[Som05] Klaus P. Sommer. In das Deutschland ,,von Hilbert und Einstein”. Briefe von Einstein, Planck, Nernst, Debye, Born, Sommerfeld, Courant, Ehrenfest, Weyl und Althoff an David Hilbert, gefunden auf einem Göttinger Dachboden. (German) [In the Germany of “Hilbert and Einstein”. Letters of Einstein, Planck, Nernst, Debye, Born, Sommerfeld, Courant, Ehrenfest, Weyl and Althoff to David Hilbert, found in a Goettingen loft]. Berichte zur Wissenschaftsgeschichte, 28(4):283–303, December 2005. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Stoltzenberg:1998:FHC

[Sto98] Dietrich Stoltzenberg. Fritz Haber: Chemiker, Nobelpreisträger, Deutscher, Jude: eine Biographie. (German) [Fritz Haber: Chemist, Nobel Laureate, German, Jew: a biography]. VCH, Weinheim, Germany, second edition, 1998. ISBN 3-527-29573-9 (paperback). xiv + 669 pp. LCCN ????

Strong:1960:BRB

[Str60] John Strong. Book review: Principles of Optics. Electromagnetic theory of propagation, interference and diffraction of light, by Max Born and Emil Wolf. Science (New Series), 131(3399):495, February 19, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/pdfplus/1706118.pdf;   http://www.
sciencemag.org/content/131/3399/495.1.full.pdf.

Stroke:1964:BRE

[Str64] George W. Stroke. Book review: Electro-optical sciences: Principles of Optics. Electromagnetic theory of propagation, interference, and diffraction of light, by Max Born and Emil Wolf. Science (New Series), 146(3648):1154, November 27, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/pdfplus/1713930.pdf;   http://www.
sciencemag.org/content/146/3648/1154.1.full.pdf.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Sur:1999:AAC

[Sur99] Abha Sur. Aesthetics, authority, and control in an Indian laboratory: The Raman–Born controversy on lattice dynamics. Isis, 90(1):25–49, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237473.

Susskind:1957:BRB

[Süs57] Charles Süsskind. Book review: Physics in my Generation, by Max Born. Physics Today, 10(6):48–50, June 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/link.aip.org/link/?PTO/10/48/1;   http://physicstoday.
org/resource/1/phtoad/v10/i6/p48_s1.

Thorpe:1922:LEH

[Tho22] T. E. Thorpe. Letter to the Editor: [Haber on chemical warfare]. Nature, 104(2724):40–41, January 12, 1922. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:/

/www.nature.com/nature/journal/v109/n2724/pdf/109040b0.

pdf. See [Hab22].

Urey:1967:CAS

[Ure67] Harold C. Urey. Comments: Affording the space program. Bulletin of the Atomic Scientists, 23(2):24–25, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ano67, Bor68b].

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland, Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

vanderWaerden:2007:SQM

[vdW07] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics. Dover Publications, Inc., New York, NY, USA, 2007. ISBN 0-486-45892-X (paperback). xi + 430 pp.

LCCN QC174.12 .W34 2007. URL http://www.loc.gov/catdir/
enhancements/fy0702/2006050791-d.html.

vonLaue:1952:JFM

[vLP52] Max von Laue and R. W. Pohl. James Franck [und] Max Born. Zeitschrift für Physik, 133(1–2):i–iii, September 15, 1952. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01948675. Two photographs captioned: Zum 70. Geburtstag am 26. August 1952 (James Franck) and Zum 70. Geburtstag am 11. Dezember 1952 (Max Born), followed by birthday greetings.

Soldner:1921:BAL

[vS21] J. (Johann Georg von) Soldner. Über die Ablenkung eines Lichtstrahls von seiner geradlinigen Bewegung durch die Attraktion eines Weltkörpers, an welchem er nähe vorbeigeht; von J. Soldner, 1801. (German) [On the deviation of a light ray from its straight-line motion due to the attraction of a heavenly body that it passes by closely]. Annalen der Physik (1900) (series 4), 65(15):593–604, 1921. ISSN 0003-3804 (print),

1521-3889 (electronic). URL http://en.wikipedia.org/wiki/
Johann_Georg_von_Soldner;   http://en.wikisource.org/?
curid=755966;   http://onlinelibrary.wiley.com/doi/10.

1002/andp.19213701503. With a foreword, and considerable editing and abridging, by Philipp Lenard. An English translation of the complete original 1801 article is available in Wikisource.

vonWeizsacker:1957:SGA

[vW57] Carl Friedrich von Weizsäcker. Should Germany have atomic arms. Bulletin of the Atomic Scientists, 13(8):283–286, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Followup on [BBF^{+}57a].

Weiner:1972:MNP

[Wei72] Charles Weiner. 1932 — moving into the new physics. Physics Today, 25(5):40–49, May 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.org/

content/aip/magazine/physicstoday/article/25/5/10.1063/

1.3070853. Reprinted in [Wei85].

Weisskopf:1982:MGS

[Wei82] Victor Weisskopf. Meine Göttinger Studienjahre mit Born und Franck. (German) [My Göttingen study years with Born and Franck]. In Max Born, James Franck, Physiker in ihrer Zeit: der Luxus d. Gewissens. (German) [Max Born, James Franck, physicists in their time: The Luxury of Conscience] [Lem82], pages 80– 083. ISBN 3-88226-148-X. LCCN QC7 .M33 1982. With contributions from Friedrich Hund, H. Maier-Leibnitz and Victor F. Weisskopf.

Weiner:1985:MNP

[Wei85] Charles Weiner. 1932 — moving into the new physics. In Weart and Phillips [WP85], pages 332–339. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Wei72].

Weinstein:2012:MBA

[Wei12] Galina Weinstein. Max Born, Albert Einstein and Hermann Minkowski’s space–time formalism of special relativity. arxiv.org, ??(??):??, October 25, 2012. URL http://arxiv.org/abs/1210.

6929.

Williams:1967:CMB

[Wil67] Lloyd Williams. Comments: On Max Born’s “reflections”. Bulletin of the Atomic Scientists, 23(2):27–28, February 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bor65h].

Wolbarsht:1966:BRB

[Wol66] Myron L. Wolbarsht. Book review: Principles of Optics. Electomagnetic Theory of Propagation, Interference and Diffraction of Light, by Max Born and Emil Wolf. Quarterly review of biology, 41(4):440, December 1966. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL http://www.jstor.

org/stable/pdfplus/2818393.pdf.

Wolf:1983:RMB

[Wol83] Emil Wolf. Recollections of Max Born. Optics News, 9(6):10–16, November/December 1983. CODEN ONEWDU. ISSN 0098-907X.

Wolf:1995:RMB

[Wol95] Emil Wolf. Recollections of Max Born. Astrophysics and Space Science, 227(1–2):277–297, May 1995. CODEN APSSBE. ISSN 0004-640X (print), 1572-946X (electronic). URL http://link.

springer.com/article/10.1007/BF00678085.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.